### v1

# Predicate Invention from Pixels via Pretrained Vision-Language Models

Abstract
Our aim is to learn to solve long-horizon decision-making problems in highly-variable, combinatorially-complex robotics domains given raw sensor input in the form of images. Previous work has shown that one way to achieve this aim is to learn a structured abstract transition model in the form of symbolic predicates and operators, and then plan within this model to solve novel tasks at test time. However, these learned models do not ground directly into pixels from just a handful of demonstrations. In this work, we propose to invent predicates that operate directly over input images by leveraging the capabilities of pretrained vision-language models (VLMs). Our key idea is that, given a set of demonstrations, a VLM can be used to propose a set of predicates that are potentially relevant for decision-making and then to determine the truth values of these predicates in both the given demonstrations and new image inputs. We build upon an existing framework for predicate invention, which generates feature-based predicates operating on object-centric states, to also generate visual predicates that operate on images. Experimentally, we show that our approach — pix2pred — is able to invent semantically meaningful predicates that enable generalization to novel, complex, and long-horizon tasks across two simulated robotic environments.
1 Introduction
A core research objective in robotics is to develop a robot that can make long-term decisions from low-level sensory inputs to accomplish a very broad range of tasks. Recent work has shown that model-free imitation learning can solve complex tasks directly from pixels (Chi et al. 2023; Zhao et al. 2023). However, sequential and compositional generalization to novel tasks beyond the training distribution remain open challenges, especially when task horizons are long and demonstration datasets are small. In this work, we take an aggressively model-based approach to address these challenges. In particular, we use sparse demonstration data to learn abstract world models (Silver et al. 2023; Liang et al. 2024) and plan in those world models to achieve strong compositional and sequential generalization.
Figure 1: Example train vs. test task in the “Burger” environment. Training tasks involve creating a single burger with a single ingredient or stacking two ingredients. Tasks at test time involve more objects, different object arrangements, and novel compositions of goals as compared to those at training time. For example, one of the tasks at test time is to create three burgers – two with multiple ingredients – and where the robot starts out holding something (no training task starts out with the robot holding something).
Figure 2: Pix2Pred overview. pix2pred are given an initial set of predicates, a set of controllers (temporally extended actions like pick, cook, etc.) and demonstrations of solving tasks in the form of goal-directed observation-action sequences. Our method learns more predicates and learns operators in terms of the learned and given predicates. Given a novel observation and goal at test time, we use a planner to generate a plan with our learned operators that achieves the goal. Goals in the train and test tasks are defined in terms of the given predicates; goals are written as text in the figure to simplify the figure.
As a simple, pedagogical example, consider the “Burger” domain depicted in Figure 1. The agent is given demonstrations showing how to cook a patty, how to cut lettuce, how to stack ingredients, and how to make a sandwich with a single ingredient. It is then tasked with making multiple burgers, each with a cooked patty and chopped lettuce, from novel initial images where some patties may be cooked already, some lettuce chopped already, and where the robot might start out holding an object. This test task is challenging for model-free imitation learning because it is much longer-horizon than the demonstrated tasks and involves novel numbers and configurations of objects. However, if the agent can learn to recognize a small number of abstract state features, such as when a particular patty is Cooked, when a particular head of lettuce is Chopped, and whether or not the robot is currently holding an object, – as well as abstract transition rules involving these concepts: cooking a patty causes it to go from Raw to Cooked, chopping lettuce causes it to go from Whole to Chopped, a robot can only pick up an object if it is not Holding something else – it can combine these concepts in many ways when planning and generalize substantially to new tasks.
We are specifically interested in learning PDDL models (Fox and Long 2003; McDermott et al. 1998) in the form of symbolic predicates (e.g., Cooked) and operators (e.g., cooking a patty causes it to go from Raw to Cooked) that can be leveraged by powerful classical planning algorithms (Helmert 2006). Such models feature several biases that are conducive to efficient planning and learning. Firstly, they are relational: world states and goals are characterized in terms of properties and relations between objects. Secondly, they are object-oriented: the number and properties of objects may vary across different problem instances, allowing the same model to support a wide range of specific tasks. Finally, they are sparse: each modeled action depends on and affects only a relatively small number of other objects, abstracting out irrelevant details and predicting core cause-effect interactions. Several recent works have demonstrated that such models can be learned efficiently to support generalization to tasks featuring new states, new goals, more objects, and longer time horizons than those demonstrated at training time (Silver et al. 2021, 2022, 2023; Asai and Fukunaga 2018; Ahmetoglu et al. 2022). However, missing from previous work is a method for learning predicates that can be computed from raw visual observations from very few demonstrations.
To address this gap, our first key idea is to use Vision-Language Models (VLMs) to propose and ground predicates. Given demonstrations, we leverage the commonsense knowledge of the VLM propose predicates with natural language names. Given an invented predicate and new states, we can again use the VLM to ground and evaluate the predicate directly on images. However, we find experimentally that the VLM’s proposed set of predicates do not support efficient planning or generalization across training tasks. Given this, our second key idea is to generate a large pool of candidates and subselect predicates by explicitly optimizing for efficient and effective planning. We extend a recent method for program-synthesis-based predicate invention (Silver et al. 2023) that searches over sets of predicates via hill-climbing to minimize a fast-planning objective. After running this procedure to subselect predicates, we obtain a compact set of predicates that supports efficient and effective downstream decision-making.
In experiments, we evaluate the extent to which our method — pix2pred — is able to learn symbolic predicates and operators that generalize to novel problems. Across four tasks in two simulated environments, we compare to several related baselines and find that pix2pred consistently achieves the highest success rate on novel problem instances involving more objects, more complex goals, and longer horizons than those demonstrated during training
2 Related Work
Decision-Making with Foundation Models. Our work is strongly inspired by impressive recently-demonstrated capabilities of Large (Vision) Language Models (LLMs, VLMs) across a variety of challenging text and image tasks (OpenAI 2023; Team et al. 2023; Dubey et al. 2024). We build on a large body of work that leverages foundation models for decision-making in robotics tasks (Ichter et al. 2023; Singh et al. 2023; Liang et al. 2022; Huang et al. 2022; Curtis et al. 2024; Huang et al. 2023, 2024; Hu et al. 2023a; Duan et al. 2024; Yang et al. 2024) (see Hu et al. (2023b) for a recent survey). However, many of these approaches operate under a number of different restrictive assumptions. Some are confined to a particular task distribution (e.g., only pick-and-place) or seek to synthesize new relatively short-horizon skills (e.g., pouring a liquid into a cup) (Huang et al. 2023, 2024; Duan et al. 2024). Others assume skills are provided, and attempt to use foundation models to compose them to solve longer-horizon tasks (Ichter et al. 2023; Singh et al. 2023; Liang et al. 2022; Huang et al. 2022; Curtis et al. 2024; Hu et al. 2023a; Quartey et al. 2024; Kumar et al. 2024a). These approaches all use a foundation model directly for planning, which has been shown to perform significantly worse than dedicated planning algorithms that use an appropriate abstraction (Valmeekam et al. 2022; Kambhampati et al. 2024). By contrast, our approach leverages foundation models to construct abstractions that are fed into a planning system at inference time. We treat the work of Hu et al. (2023a) as representative of previous approaches that use foundation models to plan, and compare directly to variants of this approach in our experiments (Section 5).
Learning Abstractions for Planning. We build on a long line of work that learns abstractions for efficient planning (Bertsekas, Castanon et al. 1988; Jong and Stone 2005; Abel, Hershkowitz, and Littman 2016; Ugur and Piater 2015; Asai and Fukunaga 2018; Konidaris, Kaelbling, and Lozano-Pérez 2018; Silver et al. 2021; James, Rosman, and Konidaris 2022; Mao et al. 2022; Chitnis et al. 2022; Kumar et al. 2023; Silver et al. 2023; Han et al. 2024; Liu et al. 2024). Several of these works assume that predicates or operators are already given (Chitnis et al. 2021; Kumar et al. 2023; Mao et al. 2022; Siegel et al. 2024) or focus on the problem of learning predicates and other abstractions from online interaction (James, Rosman, and Konidaris 2022; Liang et al. 2024), or assume access to dense natural language descriptions with each demonstration (Liu et al. 2024; Han et al. 2024). By contrast, we aim to invent predicates and operators in an offline fashion from image-based demonstrations with no supervision on the kinds or number of predicates to learn. Like us, “skills to symbols” (Konidaris, Kaelbling, and Lozano-Pérez 2018) performs predicate invention given skills, but operates on point cloud data and uses a bisimulation objective to select predicates that is different from the fast-planning objective that we use. Our work builds directly on the method from  Silver et al. (2023), which learns predicates and operators that operate over a handcrafted, low-dimensional feature space. We extend this approach to learn predicates and operators that operate directly on images, and we demonstrate experimentally (Section 5) that this is necessary to solve tasks where attributes of objects important for decision-making cannot be easily extracted and encoded into a low-dimensional feature space.
3 Problem Setting and Background
We consider the problem of learning from demonstrations in deterministic, fully-observed environments. pix2pred takes as input a set of core object-types with meaningful natural-language names, an initial set of predicates, a set of parameterized controllers, and demonstrations of solving tasks in the form of goal-directed observation-action sequences, where goals are expressed in terms of the initial predicates and the sequences are segmented by controller. Each observation contains raw sensor data that includes one or more images characterizing the world state at that time, as well as data from sensors on the robot. We assume access to a perception function that can identify basic attributes of objects (e.g., position, dimensions, etc.) from the raw sensor data and can annotate objects of the given object-types on the images with unique names. Each controller execution in the demonstration is labeled with its name, the names of the objects it operates on in terms of names given by the perception function, and the continuous parameters it was executed with. From this, pix2pred learns predicates and operators (specifically in the form of a PDDL domain file (McDermott et al. 1998)), as well as samplers that produce the continuous parameters for the controllers. Below, we describe the precise structure of our demonstrations more formally. We also provide background on symbolic predicates, operators, samplers, and an existing framework for predicate invention that we build on.
3.1 Demonstrations
Our demonstrations are provided within a set of training tasks drawn from some distribution $\mathcal{T}$. We are ultimately interested in learning to solve any task $T$ drawn from this distribution. All tasks occur within a common environment.
Environments and Tasks
We model an environment as a tuple $\langle I,A,f,\Lambda,P\rangle$, where $I$ is the observation space, $A$ is the action space, and $f$ is an unknown transition function, denoted $f:I\times A\rightarrow I$. An observation $\text{obs}\in I$ consists of a sequence of one or more images ($\text{obs}^{\text{img}}$) representing camera images from cameras operated by the robot, as well as a vector of continuous real-valued numbers ($\text{obs}^{\text{feat.}}$) representing raw sensor data associated with the robot, such as GPS and encoder data. The agent receives some observation $\text{obs}_{t}$ at some timestep $t$, and can execute an action $a\in A$. The environment will then return a new observation $\text{obs}_{t+1}\leftarrow f(\text{obs}_{t},a)$. We assume $f$ is deterministic. $\Lambda$ is a finite set of object types, where each type $\lambda\in\Lambda$ has a name (e.g., robot, patty, etc.) and a tuple of real-valued features representing basic attributes of the object (e.g., position for most objects, position and joint angles for the robot).
Within an environment, a task is a tuple $\langle\mathcal{O},i_{0},g\rangle$. $\mathcal{O}$ is a set of objects, each with a type from $\Lambda$. The environment’s perception function $P:I\times\mathcal{O}\rightarrow\mathcal{X}$ processes each observation into a structured state used for downstream decision-making. These states have two parts: an image-based state, ($x^{\text{img}}$), and a feature-based state, ($x^{\text{feat}}$), concatenated together to form a state $x\in\mathcal{X}$. The perception function gives a unique name to each object in the observations and annotates the images with their names, as shown in Figure 2; this forms the image-based state. In practice this can be achieved by combining open-vocabulary object-detection and segmentation models (Ren et al. (2024)). The perception function also identifies the basic attributes of each object according to its type; this forms the feature-based state. In practice, this can be implemented with custom detectors for each property of each object, and is simple when the attributes are basic properties like pose and dimensions. We assume that objects are fully-observable to the perception function (i.e., there is no observation in which certain objects or types of objects are unobservable). $i_{0}$ is the initial observation of the task, and $x_{0}$ represents the initial state of a task produced by the perception function, where $x_{o}\leftarrow P(i_{0},\mathcal{O})$. We will henceforth refer to a task as involving this initial state: $\langle\mathcal{O},x_{0},g\rangle$. We describe the goal $g$ in a the next section.
Initial Predicate Set and Goals
In order to specify the goal $g$ of a task, as well as provide a basis for decision-making and learning, we provide an initial set of predicates $\Psi_{\text{init}}$ as part of the demonstrations. Each predicate $\psi$ can be thought of as a template for a binary classifier over states. Specifically, a predicate is characterized by a name, an ordered list of types $(\lambda_{1},\dots,\lambda_{m})$ and a classifier function $c_{\psi}:\mathcal{X}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$ parameterized by typed object arguments. A ground atom $\underline{\psi}$ consists of a predicate $\psi$ and objects $(o_{1},\dots,o_{m})$. Each ground atom induces a particular binary state classifier $c_{\underline{\psi}}:\mathcal{X}\to\{\text{true},\text{false}\}$, where $c_{\underline{\psi}}(x)\triangleq c_{\psi}(x,(o_{1},\dots,o_{m}))$. For example, the ground atom Cooked(patty0) would have the value “true” if the object patty0 is cooked. [^1] It would have this value if its classifier was accurate. We discuss what happens if this classifier is inaccurate in Section 4.3.
We assume the initial predicates $\Psi_{\text{init}}$ are sufficient for representing task goals, and may additionally contain concepts useful across all tasks in the environment, but are insufficient on their own in that additional predicates will be necessary to capture all concepts critical for solving tasks. Specifically, the goal $g$ of a task is a set of ground atoms involving predicates in $\Psi_{\text{init}}$ and objects in $\mathcal{O}$. A goal $g$ is said to hold in a state $x$ if for all ground atoms $\underline{\psi}\in g$, the classifier $c_{\underline{\psi}}(x)$ returns true. We use the term “goal predicates” for predicates that are used in defining the goal. We provide a discussion about the necessity of predicate invention despite having these predicates in Appendix A.4.
Actions as Hybrid Controllers
The environment’s action space $A$ is defined by a finite set of controllers $\mathcal{C}$. Each controller $C((\lambda_{1},\dots,\lambda_{v}),\Theta)\in\mathcal{C}$ has a semantically-meaningful name, as well as optional discrete typed parameters $(\lambda_{1},\ldots,\lambda_{v})$ and continuous parameters $\Theta$. For instance, a controller Pick for picking up an object might have one discrete parameter of type item and a $\Theta$ that is a placeholder for a specific grasp. An action $a\in\mathcal{A}$ in a task is a controller $C\in\mathcal{C}$ with discrete and continuous arguments fully specified: $a=C((o_{1},\ldots o_{v}),\theta)$.
Final Demonstration Structure
We provide the agent with a demonstration set $\mathcal{D}$. Each demonstration consists of: (1) a training task $T\sim\mathcal{T}$, (2) a near-optimal plan that solves the task, and (3) the state sequence of that plan with values for all initial predicate ground atoms.
Recall that a task is a tuple $\langle\mathcal{O},x_{0},g\rangle$. A plan $\pi=(a_{1},\ldots,a_{m})$ is a sequence of actions $a\in\mathcal{A}$ such that successive execution of the actions in the environment, starting from $x_{0}$, results in a final state $x_{m}$ where $g$ holds. A state sequence of a plan is the set of states $(x_{1},\ldots,x_{m+1})$ encountered as that plan is executed in an environment. An example of this input is shown in Figure 2.
Given a handful (less than $15$) of these demonstrations, the agent’s objective is to invent additional predicates and a set of operator definitions in terms of the initial and learned predicates that enables the agent to plan and efficiently solve a set of held-out tesk tasks drawn from $\mathcal{T}$. Test tasks involve more objects, different object configurations, novel compositions of goals, and longer-horizons than the training tasks solved in the demonstrations.
3.2 Operators and Planning
Predicates $\Psi$ induce an abstract state space $S_{\Psi}$ of the task’s underlying continuous state space $\mathcal{X}$. Operators help specify an abstract transition model over $S_{\Psi}$.
More specifically, given a particular state $x_{t}$, as well as predicates $\Psi$, we can compute the corresponding abstract state (denoted by $s_{t}$) by selecting all the ground atoms that are “true” in $x_{t}$. Formally, $s_{t}=\textsc{Abstract}(x_{t},\Psi)\triangleq\{\underline{\psi}:c_{\underline{\psi}}(x_{t})=\text{true},\forall\psi\in\Psi\}.$ Then, each operator $\omega$ is a tuple $\omega=\langle\overline{v},P,E^{+},E^{-},\mathcal{C}\rangle$. Here, $\overline{v}$ are typed variables representing the operator’s arguments, $P$, $E^{+}$ and $E^{-}$ are sets of predicates representing operator preconditions, add effects and delete effects respectively, and $\mathcal{C}$ is a controller associated with the operator. Note importantly that the discrete parameters of $\mathcal{C}$ must be a subset of $\overline{v}$. Specifying objects of the correct type as arguments induces a ground operator $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$. Ground operators define a (partial) abstract transition function: given $\underline{\omega}=\langle\underline{P},\underline{E^{+}},\underline{E^{-}},\underline{\mathcal{C}}\rangle$, if $\underline{P}\subseteq s$, then the successor abstract state $s^{\prime}$ is $(s\setminus\underline{E^{-}})\cup\underline{E^{+}}$ where $\setminus$ is set difference.
A ground operator specifies a corresponding ground controller $\underline{\mathcal{C}}$. To extract an action $a$ from this ground controller, we must specify values for the parameters $\Theta$ associated with controller $\mathcal{C}$. Following previous work (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022), we leverage continuous parameter samplers to do this.
Given these components, we can perform planning to solve test tasks. We assume access to a planner that takes a set of predicates, operators, and samplers, as well as a task $T=(\mathcal{O},x_{0},g)$ associated with an environment, and outputs a plan that achieves the goal at the symbolic level, if such a plan exists. We implement such a planner via off-the-shelf classical planning algorithms (Helmert 2006), with details in Appendix A.1, though we note that our predicate invention framework is compatible with other planning strategies.
3.3 Predicate Invention Problem
Given demonstrations $\mathcal{D}$ structured as described in Section 3.1, our objective is to learn additional predicates ($\Psi$), operators ($\Omega$), and samplers ($\Sigma$) that enable aggressive generalization to novel test tasks from the task distribution $\mathcal{T}$ by calling a planner with these learned components. We directly leverage an existing approach to learn neural network samplers from our demonstration data (Appendix A.3). We are mainly interested in learning predicates and operators.
Predicate Invention via Hill-Climbing
We build on the framework introduced by Silver et al. (2023) for predicate invention. Their method first generates an initial pool of candidate predicates $\Psi_{\text{pool}}$ from a grammar and then selects the subset of predicates from this pool that maximizes an objective function designed to measure planning efficiency over the demonstrations, where the selection is done via a hill-climbing optimization. Predicates subselected from the pool $\Psi_{\text{pool}}$ are added to the initial set $\Psi_{\text{init}}$ to form a final set of predicates $\Psi$. After, operators are learned in terms of the predicates $\Psi$. Intuitively, optimizing their particular objective fuction yields the predicate set (and corresponding operator set) that produces plans that most closely align with the demonstrations on the training tasks.
Importantly, this previous work makes two strong assumptions that do not hold in our setting: (1) the state space $\mathcal{X}$ is composed entirely of continuous object features (i.e., $\mathcal{X}$ only consists of $x^{\text{feat}}$), and (2) learned predicate classifiers are not noisy. We relax these assumptions and enable predicates that operate over images.
4 Inventing Predicates from Pixels
Our main interest is to invent predicates from pixels (i.e., predicates that operate directly over one or more images in our image-based state $x^{\text{img}}$). We wish to do this from a handful of structured demonstrations described in Section 3.1. These demonstrations do not possess any direct supervision for additional predicates (i.e., the demonstrations provide no indication of the number or structure of any predicates we should invent). The hill-climbing framework for predicate invention (Section 3.3) is able to learn predicates in such a setting, but requires a pool of initial candidate predicates to subselect from. To define such an initial pool, we must specify two things: what the predicates should be (i.e., what visual concepts are relevant for decision-making), and how should they be implemented (i.e., for every predicate $\psi$ in the pool, how can we define a corresponding classifier function $c_{\psi}$).
We propose to use a VLM to both generate an initial pool of predicates that might be relevant to decision-making, and implement the classifier function for each predicate in the pool. Recent work has demonstrated that VLMs are able to answer a variety of common-sense natural language questions given one or more images (Antol et al. 2015; Kim et al. 2024; Majumdar et al. 2024). The core intuition motivating our approach is that predicate generation and implementation correspond to two different types of questions. Predicate generation can be accomplished by giving the VLM access to our demonstrations, and asking “what visual concepts are changing between false and true after each action execution?”. Predicate classifier implementation can be accomplished by giving the VLM access to a particular image and asking it “is this particular ground predicate true or false in the current state?”. In the remainder of this section, we explain how exactly we implement this procedure and how we modify and leverage the hill-climbing predicate invention framework described in Section 3.3 to yield a compact subset of predicates optimized for planning. [^2] See Figure 8 for a detailed overview of our entire pipeline.
4.1 Implementing Visual Predicates with a VLM
Recall that each state $x\in\mathcal{X}$ is composed of two parts: a feature-based state $x^{\text{feat}}$ and an image-based state $x^{\text{img}}$. Visual predicates operate exclusively over $x^{\text{img}}$. Specifically, a visual predicate is a tuple of a name, a sequence of $m$ types, and a classifier function: $\psi_{\text{vis}}=\langle\text{name},(\lambda_{1},\dots,\lambda_{m}),c_{\psi}\rangle$. Here, the classifier function returns a boolean given an image-based state and object arguments $c_{\psi}:\mathcal{X}^{\text{img}}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$.
We propose to implement the classifier function $c_{\psi}$ by querying a VLM. Specifically, we create a prompt string $\text{txt}_{\psi}$ with the predicate’s name and arguments to be filled in by the predicate’s typed object arguments $(\lambda_{1},\dots,\lambda_{m})$ (e.g. Cooked(?obj1)). [^3] We create the prompt string $\text{txt}_{\psi}$ with the predicate’s name but other choices are possible. We could ask the VLM, “Is patty1 cooked?” instead of “What is the value of Cooked(patty1)?") When the predicate is ground, we substitute the specific objects used to ground it into the prompt (e.g. Cooked(patty1)). We implement the classifier function by simply passing the prompt with this ground atom string into a VLM with some additional instruction, asking it to output either “true”, “false”, or “unknown” (which we take to be equivalent to “false”).
While this idea is straightforward, we found that several additional modifications to the text prompt $\text{txt}_{\psi}$ and images $x^{\text{img}}$ were critical to obtain accurate labels for ground atoms from the VLM. We discuss these in detail and provide examples of our prompts in Appendix A.5.
4.2 Proposing an Initial Pool of Visual Predicates
Figure 3: VLM predicate proposal example. We prompt (Appendix A.5) a VLM to propose a large pool of predicates that describe important concepts in the given demonstrations. Technically, the VLM outputs ground atoms as strings, and we parse the predicate (visualized in orange) from each ground atom. Shown here are proposals generated from demonstrations in the task distribution “Combo Burger” within the “Burger” environment (Section 5). The VLM proposes 563 predicates and we filter these into 136 “valid” predicates; the figure shows 56 of these.
Figure 4: VLM predicate labeling example. We prompt (Appendix A.5) a VLM to determine the truth values of ground atoms (each represented as a string) in the given image, labeling each as either true, false, or unknown. Shown here are a subset of the labels generated for a particular state from a task in the task distribution“Combo Burger” within the “Burger” environment (Section 5).
We generate a pool candidate of visual predicates by prompting a VLM on each demonstration $d\in\mathcal{D}$. For a demonstration $d\in\mathcal{D}$ of length $k=len(d)$, we extract the image-based state at each timestep $x^{\text{img}}_{t},t\in[0,k-1]$ and add a text heading to each image corresponding to which timestep $t$ in the demonstration it belongs to. We then pass all these images in sequence, along with the actions executed in between them ($a_{0},a_{1},\ldots,a_{k_{d}-1}$) directly to a VLM. We prompt the VLM to output a set of proposals for ground atoms based on these inputs.
Given a set of ground atoms proposed for each demonstration, we then parse this set to discover atoms that are syntactically incorrect. In particular, we remove atoms that include object names that are not part of the demonstration $d$ (for instance, the VLM might propose an atom ontop(patty1, pot1), even though there is no object named pot1). We then lift each of these atoms into a visual predicate by creating typed variables for each of the objects identified in each proposed ground atom (for instance, an atom cooked(patty1) would be lifted to cooked(?p: patty)). Finally, we remove any duplicate lifted atoms and add only unique predicates to our pool. We obtain a final pool of predicates $\Psi_{\text{pool}}^{\text{vis}}$ by running this procedure on each demonstration.
4.3 Predicate Invention via Subset Selection From Noisy Data
Given the procedure described in Section 4.2, we can generate an initial pool of visual predicates $\Psi_{\text{pool}}^{\text{vis}}$. We combine these with feature-based predicates (generated via a programmatic grammar as described in Silver et al. (2023)) to create an overall pool $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$. We wish to run hill-climbing optimization as described in Section 3.3 to subselect a small set of relevant predicates that are optimized for efficient and effective planning over our training tasks.
Figure 5: Labeling of ground atoms with visual predicates involving the patty object type in Burger. This figure depicts the visual state before and after the cook skill in demonstration #0 for the task distribution “Combo Burger”. It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the patty object in these states. The VLM proposes a wide variety of predicates involving the patty: clear, cooked, cool, grilling, raw, and uncook. Of these, cooked(?p:patty) is most relevant to achieving the goal while also being accurately labeled. Our approach automatically selects this predicate from the pool. See Figure 9 in the appendix for an analogous situation with the chop skill.
Unfortunately, we cannot directly apply the particular hill-climbing optimization from previous work because our visual predicates are noisy. Visual predicates might not be consistent across states in the demonstrations due to occasional hallucination or mislabeling on the part of the VLM. This noise creates outliers in the abstract state space transition data and causes the operator learning algorithm to overfit and create a large number of unnecessary operators to model these outlier transitions. These extra operators make planning inefficient by increasing the branching factor of the planner’s search algorithm. They also cause the hill-climbing predicate subselection – which interally performs operator learning on every subset of predicates it scores – to overfit and select a large number of unnecessary predicates. These overfit predicates can appear in the definitions of all operators – not just the operators created to model the outlier transition data – because operators are learned to be chainable (add effects of an earlier operator satisfy the preconditions of a later operator). The overfit predicates and overfit operators fail to generalize outside of the training distribution. We combat this problem by modifying the operator learning approach to be more tolerant to noise  [^4] Modifications are described in detail in Appendix A.2, and by regularizing the hill-climbing optimization itself via early stopping to discourage overfitting.
5 Experiments
Our experiments are designed to answer the following questions.
  •       Q1.
    How well does pix2pred generalize to novel, more complex goals when compared to an imitation approach that doesn’t use a planner?
  •       Q2.
    How critical is it to perform explicit score-based optimization for subselection of predicates?
  •       Q3.
    How critical is it to invent both visual and feature-based predicates?
Figure 6: Environments. Top row: train task examples. Bottom row: test task examples. See Section 5 for descriptions of each environment.
Environments. We now describe our experimental environments and tasks, with details in Appendix A.6. We implement 4 task distributions across 2 environments. Each environment was introduced by previous work and lightly adapted for our setting.
  • $\bullet$
    Kitchen: A simulated robotic environment featuring a robotic arm positioned in front of a kitchen that includes a stove, a microwave, and controllable lights (Gupta et al. 2019). The task is to boil water in a kettle by turning on a specific stove burner and pushing the kettle onto it. The task has a short horizon, requiring only two high-level skills in sequence to achieve the goal. Demonstrations show the robot solving the task using a particular burner from various kettle initial positions, and at test time the robot must use a different burner on the same stove.
  • $\bullet$
    Burger: A simulated 2D grid world environment (based on the Robotouille environment introduced by (Wang et al. 2023)) featuring a robot that must prepare and assemble burgers from various components. Preparation includes chopping lettuce on a cutting board and cooking patties on a grill. We experiment with three task distributions within this environment, with each distribution featuring its own set of training and testing tasks: (1) Bigger Burger involves producing a burger with extra patties after being shown how to create a burger with a single patty, (2) More Burger Stacks involves producing multiple burgers after being shown how to create a single burger, and (3) Combo Burger involves producing burgers with multiple ingredients after being shown how to create burgers with a single ingredient. The tasks have a long horizon. For example, test tasks in Bigger Burger, More Burger Stacks, and Combo Burger can require a sequence of 16, 29, and 30 controllers to achieve the goal, respectively.
Figure 7: pix2pred versus baselines. Percent of evaluation tasks solved per task across both simulated environments. All results are averaged over 5 seeds. Black bars denote standard deviations.
Approaches. We evaluate our method against several ablations and baselines from the literature, including those that attempt to solve test tasks directly without learning any abstractions.
  • $\bullet$
    Ours: Inventing visual and feature-based predicates with pix2pred, learning operators in terms of these predicates, and planning with these operators at test time.
  • $\bullet$
    VLM subselect: An ablation of pix2pred where we have the VLM directly subselect a compact set of visual predicates instead of delegating subselection to the hill-climbing optimization. We still learn feature-based predicates via hill-climbing in this case.
  • $\bullet$
    No subselect: An ablation of pix2pred where we perform no hill-climbing, but rather select all predicates in $\Psi_{\text{pool}}$.
  • $\bullet$
    No invent: An ablation of pix2pred where we do not propose or subselect any additional predicates beyond the provided ones in $\Psi_{\text{init}}$.
  • $\bullet$
    No visual: An ablation of pix2pred where we only invent feature-based predicates.
  • $\bullet$
    No feature: An ablation of pix2pred where we only invent visual predicates.
  • $\bullet$
    VLM feat. pred: A baseline inspired by recent work from Han et al. (2024). Here, a VLM writes predicates as functions over the feature-based state instead of being able to invent its own novel visual concepts.
  • $\bullet$
    ViLa: A baseline that does not learn abstractions but uses a VLM to plan on test tasks directly (Hu et al. 2023a). We test two variants of this: ViLa-pure which is the original approach as implemented by Hu et al. (2023a) adapted to our setting, and ViLa-fewshot, which additionally provides the demonstrations as part of the few-shot prompt to give the VLM some context of what kinds of plans it should return for testing tasks.
Experimental Setup. Each seed varies the number and locations of objects in the train and test tasks, and we average results over $5$ random seeds for each task distribution. For each seed, we sample 10 test tasks. We use GPT-4o (OpenAI 2023) as the VLM for all approaches. For our tasks Burger, as mentioned in section 4.3, we early stop the predicate subselection optimization when the hill-climbing objective value goes below $2000$ (a value we found by running pix2pred on a validation set). We evaluate each approach by providing the test task and allowing the approach to make and execute exactly one single plan.
Results and Analysis. Figure 7 shows our success rate plots. pix2pred outperforms previous approaches on 3 out of 4 of the task distributions by a wide margin. In particular, pix2pred seems to generalize to more complex tasks much better than using few-shot learning with a VLM (i.e., ViLa) (Q1). In the short-horizon kitchen task that requires 2 actions, we find that ViLa with demonstrations provided performs comparably to pix2pred. However, the performance of this baseline drops-off significantly in the longer-horizon burger tasks that require more than 15 actions. In particular, we found that ViLa seems to solve testing tasks by pattern-matching aspects of the training demonstrations and is unable to solve test tasks that are unlike any of the demonstrations. For instance, the agent never starts out holding an object in any of the training demonstrations, but it does so in about 50% of the testing task instances in the Burger environment. We observed that ViLa does not correctly put down the held object before attempting to pick up other objects, whereas pix2pred learns a predicate corresponding to the agent’s hands being empty and correctly selects this as a precondition for picking up any other object. We find that pix2pred only fails due to noise in VLM labeling: the VLM infrequently labels the initial state incorrectly such that a critical precondition is false, which prevents our planner from making a valid plan.
We also find that explicit optimization via hill-climbing leads to much better performance on testing tasks when compared to simply selecting predicates using a VLM (Q2). This can be seen from the fact that the VLM subselect baseline performs signifcantly worse than our main approach across all 4 task distributions. Qualitatively, we notice that this baseline tends to select far more predicates than necessary: in the burger domain, it selects predicates Cooked(?p1: patty) and Chopped(?l1: lettuce) (necessary for decision-making), but also extraneous predicates like Raw(?p1: patty), Available(?p1: patty), and Whole(?l1: lettuce) that the VLM labels quite inconsistently. This results in a complex set of operators that overfit strongly to the training tasks and do not generalize well. Additionally, we see that the “no subselect” baseline fails completely at all tasks, indicating that some subselection is critical; learning operators in terms of the more than 500 predicates in the proposal pool leads to extremely overfit operators.
We find that both feature-based and visual predicates are important in most of our tasks (Q3). The “no invent“ baseline fails completely in all the tasks because the initial predicates are insufficient. The “no visual.” baseline performs somewhat well in the Kitchen task, but fails in the Burger tasks where visual predicates, specifically predicates corresponding to Cooked and Chopped, are necessary, since the feature-based state does not contain sufficient information to classify these concepts. Interestingly, the predicate corresponding to whether the robot’s hand is empty can be classified more accurately in Burger tasks if features of the robot’s hand are used vs. attempting to recognize that concept visually. This leads to our approach doing better than the “no feat.” baseline in the Burger tasks, since our approach is not only limited to purely visual predicates and can invent whichever predicate has higher classification accuracy.
6 Limitations and Conclusion
We proposed pix2pred, a method for inventing symbolic predicates that operate over raw images. We found that our approach, from just a handful of demonstrations (3-12), is able to invent semantically meaningful predicates that afford efficient planning and generalization to novel tasks across a range of domains and problems.
There are several noteworthy limitations of our present work. Firstly, VLM hallucinations adversely impact predicate proposal and atom labelling – especially with small amounts of demonstration data – which can inhibit our approach’s ability to learn useful predicates and operators. Secondly, the hill-climbing framework we use for predicate invention can be extremely slow, especially as the number of demonstrations and size of the initial predicate pool grows. Finally, our approach assumes input demonstrations are segmented in terms of a provided set of parameterized controllers with names that correspond to their function.
One interesting future direction is to expand the structure of predicates a VLM is allowed to produce, perhaps by allowing it to write arbitrarily-long text descriptions or even code to define new predicates (Liang et al. 2024). Another is to improve the underlying hill-climbing optimization procedure to be significantly more efficient and tolerant to noise, enabling pix2pred to scale to much larger domains that require potentially many tens or hundreds of predicates. Eventually we’d like to extend pix2pred to automatically segment demonstrations and learn controllers so that we can learn to plan directly from the simplest low-level demonstration data.
Appendix A Appendix
A.1 Planner Implementation Details
1 Input: Task $\langle\mathcal{O},x_{0},g\rangle$, predicates $\Psi$, operators $\Omega$, samplers $\Sigma$.
2 $s_{0}\leftarrow\textsc{Abstract}(x_{0},\Psi)$
3 Call classical planner to generate an abstract plan: $\text{plan}^{\text{abs}}=\underline{\omega_{0}},\underline{\omega_{1}},...,\underline{\omega_{m}}\leftarrow\text{Planner}(s_{0},\mathcal{O},\Psi,\Omega,g)$
4 Extract controller sequence with discrete args filled in: $\text{skeleton}=\underline{\mathcal{C}_{0}},\underline{\mathcal{C}_{1}},...,\underline{\mathcal{C}_{m}}$
5 For $i=0,\dots,m$:
6  Sample $\theta\sim\sigma_{\omega_{i}}$, where $\theta$ represents the continuous parameters of $\mathcal{C}_{i}$ and $\sigma_{\omega_{i}}$ is the sampler associated with $\omega_{i}$
87  Use $\theta$ to fully ground controller $\underline{\mathcal{C}_{i}}$ into an action $a_{i}$. execute action $a_{i}$ and obtain the following state $x_{i+1}\leftarrow f(x_{i},a_{i})$
Algorithm 1 Planning and Execution pseudocode
Algorithm 1 shows the pseudocode for the planning and execution strategy we implement following recent work (Kumar et al. 2024b). Given an initial state $x_{0}$ for a task, we simply evaluate the classifiers of the given predicates $\Psi$ to convert the state into an abstract state $s_{0}$. We then call a classical planner with the task object set $\mathcal{O}$ and operators $\Omega$ to compute an abstract plan that achieves the goal (if one exists). We extract the ground controller sequence for this abstract plan, and then simply greedily execute each controller sequentially by calling the sampler associated with each operator.
We note that more sophisticated planning strategies are possible. In particular, provided a photorealistic simulation environment for the transition function $f$, we could leverage many task and motion planning (Garrett et al. 2021), such as bilevel planning (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022).
A.2 Operator Learning
We adapt the “cluster and intersect” operator learning strategy from previous work (Chitnis et al. 2022) to handle noise in predicate values inherent to our setting. Specifically, we learn a set of operators $\Omega$ from our demonstrations $\mathcal{D}$ and predicates $\Psi$ in four steps. Of these, the first three steps are largely taken directly from previous work: we introduce a modification to the third step, as well as the final step to handle noise.
  1. 1.
    Partitioning: Each demonstration can be expressed as a sequence of transitions $\{(x,a,x^{\prime})\}$, with $x,x^{\prime}\in\mathcal{X}$ and $a\in\mathcal{A}$. Recall that each action $a$ is a controller with particular discrete and continuous arguments specified; let $\underline{\mathcal{C}}$ denote the corresponding controller with the same discrete object argument values, but continuous parameter values left unspecified. First, we use $\Psi$ to Abstract all states $x,x^{\prime}$ in the demonstrations $\mathcal{D}$, creating a dataset of transitions $\{(s,a,s^{\prime})\}$ with $s,s^{\prime}\in\mathcal{S}_{\Psi}$. Next, we partition these transitions via the following equivalence relation: $(s_{1},a_{1},s_{1}^{\prime})\equiv(s_{2},a_{2},s_{2}^{\prime})$ if the effects and controllers unify, that is, if there exists a mapping between the objects such that $\underline{\mathcal{C}_{1}}$, $(s_{1}-s_{1}^{\prime})$, and $(s_{1}^{\prime}-s_{1})$ are equivalent to $\underline{\mathcal{C}_{2}}$, $(s_{2}-s_{2}^{\prime})$, and $(s_{2}^{\prime}-s_{2})$ respectively. After this step, we have effectively ‘clustered’ all transitions in $\mathcal{D}$ together: we can associate each transition $\{(s,a,s^{\prime})\}$ with a particular equivalence class.
  2. 2.
    Arguments and Effects induction: For each equivalence class created in the previous step, we create $\overline{v}$ by selecting an arbitrary transition $(s,a,s^{\prime})$ and replacing each object that appears in the controller $\underline{\mathcal{C}}$ or effects with a variable of the same type. This further induces a substitution $\delta:\overline{v}\to\mathcal{O}$ for the objects $\mathcal{O}$ in this transition. Given this, the $E^{+}$, and $E^{-}$ can then be created by applying $\delta$ to $(s^{\prime}-s)$, and $(s-s^{\prime})$ respectively. By construction, for all other transitions $\tau$ in the same equivalence class, there exists an injective substitution $\delta_{\tau}$ under which the controller arguments and effects are equivalent to the newly created $E^{+}$, and $E^{-}$.
  3. 3.
    Precondition learning: The only remaining component required to turn each equivalence class into an operator is the operator preconditions. For this, we perform an intersection over all abstract states in each equivalence class (Bonet and Geffner 2019; Curtis et al. 2021). Recall that an abstract state is simply the collection of ground atoms that are ‘true’, thus taking an intersection amounts to finding the set of atoms that are always true across every initial state of every transition $(s,a,s^{\prime})$ in the equivalence class. However, since some of our predicate classifiers might be noisy they might not always hold. Thus, we take a ‘soft’ intersection: we take any atom as a precondition that is true across more than a specific percentage (set by a hyperparameter $h_{\text{pre\_frac}}$) of transitions in the equivalence class. More specifically, $P\leftarrow\bigcap_{\tau=(s,\cdot,\cdot)}\delta^{-1}_{\tau}(s)$ if $\frac{|\delta^{-1}_{\tau}(s)|}{|\mathcal{D}|}\geq h_{\text{pre\_frac}}$, where $\delta^{-1}_{\tau}(s)$ substitutes all occurrences of the objects in $s$ with the parameters in $\overline{v}$ following an inversion of $\delta_{\tau}$, and discards any atoms involving objects that are not in the image of $\delta_{\tau}$. $|\delta^{-1}_{\tau}(s)|$ denotes the number of transitions in $\mathcal{D}$ in which the lifted atom $|\delta^{-1}_{\tau}(s)|$ holds, and $|\mathcal{D}|$ denotes the total number of transitions in $\mathcal{D}$. In our experiments, we set $h_{\text{pre\_frac}}$ to 0.8.
  4. 4.
    Pruning low-data operators: Noise in the atom values often causes there to be equivalence classes with very few data points, since transitions that have been affected by noise do not unify with other transitions in our dataset. These lead to operators that are overfit to those particular noisy transitions, which are undesirable. We combat this via simply discarding learned operators that have data below a certain fraction (denoted by hyperparameter $h_{\text{data\_frac}}$) of the total transitions associated with a particular controller $\mathcal{C}$. In particular, let $|\tau_{\mathcal{C}}|$ denote the number of transitions in $\mathcal{D}$ where the action $a$ involves a particular controller $\mathcal{C}$. For any operator $\omega$, let $|\tau_{\omega}|$ denote the number of datapoints associated with the equivalence class used to construct that operator. We only keep an operator $\omega$ if $\frac{|\tau_{\omega}|}{|\tau_{\mathcal{C}}|}\geq h_{\text{data\_frac}}$. In our experiments, we set $h_{\text{data\_frac}}=0.05$.
A.3 Sampler Learning
To enable execution, we must also learn samplers for proposing continuous controller parameters (Algorithm 1). Adapting prior work (Silver et al. 2023; Kumar et al. 2023; Chitnis et al. 2022), we train one sampler per operator, defined as:
$\sigma(x,o_{1},\dots,o_{k})=s_{\sigma}(x[o_{1}]\oplus\cdots\oplus x[o_{k}]),$
where $x[o]$ is the feature vector for $o$, $\oplus$ denotes concatenation, and $s_{\sigma}$ is a learned model. Treating this as supervised learning, we use operator-specific datasets $\tau_{\omega}$, where each of these datasets is composed of all transitions from the equivalence class used to induce this operator (Appendix A.2). Each transition $(s_{i},a_{i+1},s_{i+1})\in\tau_{\omega}$ maps operator arguments $\overline{v}$ to objects via $\delta:\overline{v}\to\mathcal{O}_{\tau}$. Using this mapping, the input for training is $x[\delta(v_{1})]\oplus\cdots\oplus x[\delta(v_{k})]$, where $(v_{1},\dots,v_{k})=\overline{v}$, and the output is the continuous parameter vector $\theta$, which are the continuous parameters used for $a_{i+1}$.
Each sampler is parameterized by two neural networks. The first predicts a Gaussian distribution over $\theta$, regressing its mean and covariance. The second network classifies $(x[o_{1}]\oplus\cdots\oplus x[o_{k}],\theta)$ as valid or invalid, enabling rejection sampling from the Gaussian. Negative examples are transitions outside $\tau_{\omega}$ but using the same controller.
A.4 More on Goals
The agent may sometimes invent visual predicates that can be interpreted as component of the task goal. For example, the goal of a task may be to make a burger with a cooked patty inside it. Here, the goal predicate would be an indicator for if this is achieved or not, perhaps called BurgerComplete(bottom_bun, patty, top_bun), in which case Cooked(patty) is an implicit component of that goal, in that BurgerComplete’s classifier implicitly checks that the patty is cooked. This is in contrast with a situation where the goal of a task is to fill a cup of water, and where the cup starts out upside down so that you can’t pour into it. Here, Upright(cup) may not be a component of HasWater(cup), in that the classifier for HasWater may not be explicitly checking if the cup is upright. In the former case, one could argue that there is no usefulness to inventing Cooked if we already have the goal predicate BurgerComplete, as we should be able to extract this concept from inside the goal predicate. But this may not be possible depending on how the classifier for the goal predicate was implemented – we may not have access to its internals. And even if we did, it may not be realistically possible to generate a dataset of burgers with and without cooked patties and label them with BurgerComplete’s classifier to obtain a dataset to train a classifier for Cooked. Furthermore, in our experimental setup, we would not necessarily need the classifier for BurgerComplete at all: at training time, any ground atoms associated with this predicate could be manually labeled in the demonstrations (easy to do when there are few goals per task), and at test time, because we execute one single plan without execution monitoring and replanning, we only need to check if the agent has solved the task once when the agent has completed its plan, which we could also do manually (analogously, the head chef at a restaurant does a manual check on the final burgers produced by the cooking staff).
A.5 Additional Prompting Details
Here, we provide additional details how we prompt a VLM for both labeling and proposal of atoms.
Atom Labeling.
Recall from Section 4.1 that the goal of atom labeling is to obtain the truth value of a particular ground atom $\underline{\psi}(o_{0},\ldots,o_{l})$ given an image-based state at timestep $t$ in a trajectory $x_{t}^{\text{img}}$. We do this by prompting a VLM with a text prompt, as well as a string representation of the atom (e.g. Cooked(patty1)) and asking it to output the truth value of the atom. Since we usually want to query for the values of many atoms at once, we provide all atoms in one single query and ask a VLM to label their truth values simultaneously.
We use a different prompt depending on whether timestep $t=0$ or $t>0$. For $t=0$, we simply provide the initial image-based state $x_{0}^{\text{img}}$ and a prompt (shown below) that asks the model to label all the values of atoms listed in the prompt with either “True’, “False” or “Unknown”.
You are a vision system for a robot. Your job is to output the values of the following predicates based on the provided visual scene. For each predicate, output True, False, or Unknown if the relevant objects are not in the scene or the value of the predicate simply cannot be determined. Output each predicate value as a bulleted list with each predicate and value on a different line. For each output value, provide an explanation as to why you labelled this predicate as having this particular value. Use the format: $<$predicate$>$: $<$truth_value$>$. $<$explanation$>$.
Predicates:
We then specify the ground atom strings whose values we’d like to query. For instance, in the “Bigger Burger” task, these might be:
on_table(patty1)
on_table(patty2)
on_table(top_bun1)
on_table(top_bun2)
prepared(patty1)
prepared(patty2)
raw(patty1)
raw(patty2)
uncooked(patty1)
uncooked(patty2)
We found two aspects of this were critical to labeling accuracy. Firstly, chain-of-thought prompting (Wei et al. 2022) was extremely useful: our prompt asks the VLM to explicitly provide reasoning for its choices in addition to labeling the truth value. Secondly, recall that the image-based state $x^{\text{img}}$ includes segmentation of object names rendered onto the image itself (as depicted in Figure 2). This is a form of set-of-marks prompting (Yang et al. 2023) and is critical to the VLM being able to identify objects being referred to in the text prompt.
For all states $x_{t}$ in a trajectory that are not the initial state (i.e., $t>0$), we prompt the VLM with the following information: (1) the image-based state from the current and previous timesteps (i.e. $x_{t}^{\text{img}}$ and $x_{t-1}^{\text{img}}$), (2) the action executed to get to the current state (i.e. $a_{t-1})$, and (3) the response of the VLM from the previous timestep.
We modify the prompt to the VLM to be as follows (here we show some of the predicates queried and a particular action for a certain timestep in a trajectory for a Burger task):
You are a vision system for a robot. You are provided with two images corresponding to the states before and after a particular skill is executed. You are given a list of predicates below, and you are given the values of these predicates in the image before the skill is executed. Your job is to output the values of the following predicates in the image after the skill is executed. Pay careful attention to the visual changes between the two images to figure out which predicates change and which predicates do not change. For the predicates that change, list these separately at the end of your response. Note that in some scenes, there might be no changes. First, output a description of what changes you expect to happen based on the skill that was just run, explicitly noting the skill that was run. Second, output a description of what visual changes you see happen between the before and after images, looking specifically at the objects involved in the skill’s arguments, noting what objects these are. Next, output each predicate value in the after image as a bulleted list with each predicate and value on a different line. For each predicate value, provide an explanation as to why you labeled this predicate as having this particular value. Use the format: $<$predicate$>$: $<$truth_value$>$. $<$explanation$>$.
Your response should have three sections. Here is an outline of what your response should look like:
[START OUTLINE]
# Expected changes based on the executed skill
[insert your analysis on the expected changes you will see based on the skill that was executed]
# Visual changes observed between the images
[insert your analysis on the visual changes observed between the images]
# Predicate values in the after image
[insert your bulleted list of ‘* ¡predicate¿: ¡truth value¿. ¡explanation¿‘]
[END OUTLINE]
Predicates: available(robot)
busy(robot)
cooked(patty1)
empty(bottom_bun1)
empty(grill)
empty(robot)
empty_grill(grill)
free(robot)
Skill executed between states: Place[robot:robot, top_bun1:top_bun, patty1:patty]
We also provide the response of the VLM from a previous timestep, an example of which is shown below.
Predicate values in the first scene, before the skill was executed: # Expected changes based on the executed skill
The skill executed is ”Pick[robot:robot, top_bun1:top_bun]”. We expect the robot to pick up the top bun. The top bun should no longer be on the ground or table and should be in the robot’s possession.
# Visual changes observed between the images
In the before image, the top bun is on the ground. In the after image, the robot is holding the top bun, indicating that the robot has successfully picked it up.
# Predicate values in the after image
* available(robot): Unknown. The previous value was unknown, and the pick action does not provide information about availability.
* busy(robot): Unknown. The previous value was unknown, and the pick action does not provide information about busyness.
* cooked(patty1): True. The patty remains cooked, as it was cooked before the action. Previously, the value was true.
* empty(bottom_bun1): False. The bottom bun remains unchanged. Previously, the value was false.
* empty(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
* empty(robot): False. The robot is now holding the top bun. Previously, the value was true.
* empty_grill(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
* free(robot): Unknown. The previous value was unknown, and the pick action does not provide information about freedom. …
We found that prompting the VLM to describe differences between two scenes in a chain-of-thought fashion led to much more accurate output than simply labeling atoms based on a single scene.
There are two additional prompting techniques we found helpful and important for labeling accuracy: (1) asking the VLM to “double-check” its label output based on the explanations it provided, and (2) augmenting the image-based state to include a close-up crop of the image for the objects the robot is currently interacting with. For this double-checking, we provide the VLM with the following prompt:
Sometimes your reasoning about the value of a predicate at the current timestep uses an incorrect value of that predicate in the previous timestep. Below, I give you give you the values of the predicates at the previous timestep once again. Please check your reasoning and provide a corrected version of your previous answer, if it needs correcting. Regardless of whether or not it needs correctly, your reply should be formatted exactly the same as the previous answer.
We use double-checking only at training time to maximize labeling accuracy before hill-climbing. We augment images with additional object crops only for tasks in the Burger domain.
Figure 9: Labeling of ground atoms with visual predicates involving the lettuce object type in Burger. This figure depicts the visual state before and after the chop skill in demonstration #1 for the task distribution “Combo Burger”. It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the lettuce object in these states. The VLM proposes a wide variety of predicates involving the patty: assemble, chopped, clear, combine, cut, intact, sliced, uncut, and whole. Of these, chopped(?l:lettuce) and cut(?l:lettuce) are the most relevant to achieving the goal while also being accurately labeled. Our approach automatically selects the former predicate from the pool.
Atom Proposal
Recall from Section 4.1 that the objective of atom proposal is to generate an initial pool of visual predicates given a set of demonstrations $\mathcal{D}$. We prompt a VLM to propose ground atoms on each demonstration $d\in D$, and then aggregate these and “lift” them into predicates. Importantly, it is crucial to have diversity in the initial predicate pool: our approach will subselect and remove irrelevant predicates, but will fail if the initial pool does not have a predicate that is important for decision-making across tasks from the task distribution $\mathcal{T}$. To promote diversity in the pool, we explicitly prompt the VLM to list all ground atoms that might be relevant to decision-making, and also to output at least one synonym and antonym of every ground atom it proposes (e.g. if it proposes ‘cooked(patty1)’, it might also propose ‘grilled(patty1)’ and ‘uncooked(patty1)’).
Below, we provide our full prompt for atom proposal:
You are a robotic vision system whose job is to output a structured set of predicates useful for describing important concepts in the following demonstration of a task. You will be provided with a list of actions used during the task, as well as images of states before and after every action execution. Please provide predicates in terms of the following objects: {objs}. For each predicate, output it in the following format: predicate_name(obj1, obj2, obj3…). Start by generating predicates that change before and after each action. After this, generate any other predicates that perhaps do not change but are still important to describing the demonstration shown. For each predicate you generate, also generate some predicates that are synonyms and antonyms so that any predicate that is even tangentially relevant to the demonstrations is generated.
Once again, we found that chain-of-thought prompting was critical to good proposals. We also found that ensuring all the image-based states have segmented objects with names rendered on them was important to get a good set of proposals.
We then provide a text representation of the action sequence executed in the demonstration trajectory. An example in a Burger task is:
Skills executed in trajectory:
Pick[robot:robot, patty1:patty]
Place[robot:robot, patty1:patty, grill:grill]
Cook[robot:robot, patty1:patty, grill:grill]
Pick[robot:robot, patty1:patty]
Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun]
Pick[robot:robot, top_bun1:top_bun]
Place[robot:robot, top_bun1:top_bun, patty1:patty]
An example response from the VLM in a Burger task is shown below.
### Predicates for Each Action
1. **Pick[robot:robot, patty1:patty]**
- Before: on(patty1, ground)
- After: holding(robot, patty1)
- Synonyms: grasp(robot, patty1)
- Antonyms: release(robot, patty1)
2. **Place[robot:robot, patty1:patty, grill:grill]**
- Before: holding(robot, patty1)
- After: on(patty1, grill)
- Synonyms: position(patty1, grill)
- Antonyms: off(patty1, grill)
3. **Cook[robot:robot, patty1:patty, grill:grill]**
- Before: raw(patty1)
- After: cooked(patty1)
- Synonyms: grill(patty1)
- Antonyms: uncooked(patty1)
4. **Pick[robot:robot, patty1:patty]**
- Before: on(patty1, grill)
- After: holding(robot, patty1)
- Synonyms: grasp(robot, patty1)
- Antonyms: release(robot, patty1)
5. **Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun]**
- Before: holding(robot, patty1)
- After: on(patty1, bottom_bun1)
- Synonyms: position(patty1, bottom_bun1)
- Antonyms: off(patty1, bottom_bun1)
6. **Pick[robot:robot, top_bun1:top_bun]**
- Before: on(top_bun1, ground)
- After: holding(robot, top_bun1)
- Synonyms: grasp(robot, top_bun1)
- Antonyms: release(robot, top_bun1)
7. **Place[robot:robot, top_bun1:top_bun, patty1:patty]**
- Before: holding(robot, top_bun1)
- After: on(top_bun1, patty1)
- Synonyms: position(top_bun1, patty1)
- Antonyms: off(top_bun1, patty1)
### Other Important Predicates
- **Initial State:**
- on(patty1, ground)
- on(bottom_bun1, ground)
- on(top_bun1, ground)
- empty(robot)
- **Final State:**
- on(top_bun1, patty1)
- on(patty1, bottom_bun1)
- empty(robot)
- **Synonyms:**
- empty(robot), free(robot)
- on(obj1, obj2). above(obj1, obj2)
- **Antonyms:** - empty(robot), holding(robot, obj)
- on(obj1, obj2), off(obj1, obj2)
A.6 Additional Experimental Details
Additional Environment Details.
Here, we describe in detail the initial predicates, training demonstrations, and test tasks for each task in each of our experimental environments. Note that the controllers listed below have discrete object parameters (indicated by the ‘?’) as well as continuous parameters where applicable $\theta$ (shown within []).
  • $\bullet$
    Kitchen:
    • $\bullet$
      Given Predicates:
      • $\bullet$
        KettleBoiling(?k: kettle, ?kn: knob, ?b burner): is true only if ?kn is turned on, and the kettle ?k is ontop of ?b, and if ?kn and ?b are ’linked’ (i.e., ?kn is the knob that causes burner ?b to glow red with heat).
    • $\bullet$
      Skills:
      • $\bullet$
        TurnOnKnob(?g: gripper, ?kn: knob, [push_dir]): Moves the gripper ?g to a fixed location near the knob, and pushes with angle “push_dir” with respect to the horizontal axis of the gripper for a fixed number of timesteps in an effort to flick the corresponding knob on.
      • $\bullet$
        PushKettleOntoBurner(?g: gripper, ?k: kettle, ?b: burner, [push_x. push_y, push_z]): Moves the gripper behind the current location of kettle ?k and then pushes along the 3D vector [push_x. push_y, push_z] for a fixed number of timesteps.
    • $\bullet$
      Training demonstrations: we provide 3 demonstrations that execute TurnOnKnob and then PushKettleOntoBurner in sequence to achieve KettleBoiling(kettle1, knob2, burner2), where burner2 and knob2 are in the back left on the stove. The kettle starts out on the front left of the stove.
    • $\bullet$
      Test tasks: Given the kettle starts out on the front right burner, we task the agent with moving it to the back right burner. In this case, plans are only two steps, and actually simply replaying the demonstration plans will work.
  • $\bullet$
    Burger: The three different task distributions we implement in this domain share common skills. We list these before listing the task-distribution-specific initial predicates, demonstrations, and goals.
    General info:
    • $\bullet$
      A “burger” consists of a top bun above a bottom bun, with one or more items in between.
    • $\bullet$
      Every type has a row/column/z attribute. The z attribute changes when an object is picked up or placed onto other objects. The grill and cutting board are of type object; the patty, lettuce, cheese, bottom bun, and top bun are of type item, which is a subtype of type object; the robot is of type robot, and has additional attributes “fingers” that indicates how open its gripper is, and “dir” that indicates the direction it is facing.
    Skills:
    • $\bullet$
      Pick(?r: robot, ?i: item, []). Moves the robot ?r to a cell adjacent to item ?i and picks it up if the robot isn’t currently holding anything.
    • $\bullet$
      Place(?r: robot, ?i: item, ?o: object, []). Moves the robot ?r to a cell adjacent to object ?o and places item ?i atop object ?o if it is holding item ?i.
    • $\bullet$
      Cook(?r: robot, ?p: patty, ?g: grill). Given the robot is adjacent to grill ?g and patty ?p is atop the grill, cooks the patty so it appears grilled.
    • $\bullet$
      Chop(?r: robot, ?l: lettuce, ?c: cutting_board). Given the robot is adjacent to cutting board ?c and lettuce ?l is atop the cutting board, chops the lettuce with a knife that’s on the cutting board so it appears chopped.
    • $\bullet$
      Bigger Burger
      • $\bullet$
        Given predicates:
        • $\bullet$
          On(?o1: object, ?o2: object): turns true when ?o1 is atop ?o2.
        • $\bullet$
          OnGround(?o1: object): turns true when ?o1 is atop a cell ?o2 that forms the ground.
        • $\bullet$
          Clear(?o: object): is true only when there is no object atop the object ?o.
        • $\bullet$
          Holding(?r: robot, ?i: item): turns true only when the robot ?r is holding the item ?i.
        • $\bullet$
          SomewhereAboveAndPrepped(?p: patty, ?b: bottom_bun): turns true when ?p is somewhere above ?b and ?p is cooked.
        • $\bullet$
          RightAboveAndPrepped(?p: patty, ?c: cutting_board): turns true when ?p is right above ?c and ?p is cooked.
        • $\bullet$
          RightAboveAndPrepped(?p: patty, ?g: grill): turns true when ?p is right above ?g and ?p is cooked.
        • $\bullet$
          RightAboveAndPrepped(?p1: patty, ?p2: patty): turns true when ?p1 is right above ?p2 and ?p1 is cooked.
      • $\bullet$
        Training demonstrations: We provide 4 demonstrations that make one burger with a single cooked patty, 4 demonstrations that cook two patties and stack them on the cutting board, and 4 demonstrations that cook two patties and stack them on the grill.
      • $\bullet$
        Test tasks: There are 10 test tasks. In 5 of these test tasks, we ask the agent to make a burger with 2 cooked patties in it. In the other 5 test tasks, we ask the agent to make a burger with 2 cooked patties in it in addition to making a single open-face burger with a cooked patty, and the agent starts out holding a raw patty.
    • $\bullet$
      More Burger Stacks
      • $\bullet$
        Given predicates:
        • $\bullet$
          On(?o1: object, ?o2: object): turns true when ?o1 is atop ?o2.
        • $\bullet$
          OnGround(?o1: object): turns true when ?o1 is atop a cell ?o2 that forms the ground.
        • $\bullet$
          Clear(?o: object): is true only when there is no object atop the object ?o.
        • $\bullet$
          Holding(?r: robot, ?i: item): turns true only when the robot ?r is holding the item ?i.
        • $\bullet$
          SomewhereAboveAndPrepped(?p: patty, ?b: bottom_bun): turns true when ?p is somewhere above ?b and ?p is cooked.
      • $\bullet$
        Training demonstrations: We provide 1 demonstration that makes two burgers, each with a single cooked patty, and 11 demonstrations that make one burger with a single cooked patty.
      • $\bullet$
        Test tasks: There are 10 test tasks. In 5 of these test tasks, we ask the agent to make 5 “open-face” burgers – burgers that consist of a cooked patty on a bottom bun. In the other 5 test tasks, we ask the agent to make 6 open-face burgers, and the agent starts out holding a raw patty.
    • $\bullet$
      Combo Burger
      • $\bullet$
        Given predicates:
        • $\bullet$
          On(?o1: object, ?o2: object): turns true when ?o1 is atop ?o2.
        • $\bullet$
          OnGround(?o1: object): turns true when ?o1 is atop a cell ?o2 that forms the ground.
        • $\bullet$
          Clear(?o: object): is true only when there is no object atop the object ?o.
        • $\bullet$
          Holding(?r: robot, ?i: item): turns true only when the robot ?r is holding the item ?i.
        • $\bullet$
          SomewhereAboveAndPrepped(?p: patty, ?b: bottom_bun): turns true when ?p is somewhere above ?b and ?p is cooked.
        • $\bullet$
          SomewhereAboveAndPrepped(?l: lettuce, ?b: bottom_bun): turns true when ?l is somewhere above ?b and ?l is choppeed.
        • $\bullet$
          SomewhereAboveAndPrepped(?l: lettuce, ?p: patty): turns true when ?l is somewhere above ?p and ?l is chopped.
      • $\bullet$
        Training demonstrations: We provide 3 demonstrations that make one burger with a single cooked patty, 3 demonstrations that make one burger with a single chopped lettuce, 3 demonstrations that place a raw patty on the cutting board, chop lettuce, and place the chopped lettuce on the patty, and 3 demonstrations that place a raw patty on the grill, chop lettuce, and place the chopped lettuce on the patty.
      • $\bullet$
        Test tasks: There are 10 test tasks. In 5 of these test tasks, we ask the agent to make two burgers, each with chopped lettuce on a cooked patty. In the other 5 test tasks, we ask the agent to make two burgers, each with chopped lettuce on a cooked patty, in addition to making another burger with a single cooked patty, and the agent starts out holding a raw patty.
A.7 Learned Predicate and Operator Examples
Figures 10, 11, 12, and show our learned predicates and operators in our “Kitchen”, “Bigger Burger”, “Burger More Stacks” and “Combo Burger” tasks respectively.
Figure 10: Learned predicates and operators for the “Kitchen” task.
Figure 11: Learned predicates and operators for the “Bigger Burger” task distribution.
Figure 12: Learned predicates and operators for the “More Burger Stacks” task distribution.
Figure 13: Learned predicates and operators for the “Combo Burger” task distribution.

### v4

@@
-Predicate Invention from Pixels via Pretrained Vision-Language Models
+From Pixels to Predicates: Learning Symbolic World Models via Pretrained Vision-Language Models
 Abstract
-Our aim is to learn to solve long-horizon decision-making problems in highly-variable, combinatorially-complex robotics domains given raw sensor input in the form of images.
-Previous work has shown that one way to achieve this aim is to learn a structured abstract transition model in the form of symbolic predicates and operators, and then plan within this model to solve novel tasks at test time.
-However, these learned models do not ground directly into pixels from just a handful of demonstrations.
-In this work, we propose to invent predicates that operate directly over input images by leveraging the capabilities of pretrained vision-language models (VLMs).
-Our key idea is that, given a set of demonstrations, a VLM can be used to propose a set of predicates that are potentially relevant for decision-making and then to determine the truth values of these predicates in both the given demonstrations and new image inputs.
-We build upon an existing framework for predicate invention, which generates feature-based predicates operating on object-centric states, to also generate visual predicates that operate on images.
-Experimentally, we show that our approach — pix2pred — is able to invent semantically meaningful predicates that enable generalization to novel, complex, and long-horizon tasks across two simulated robotic environments.
+Our aim is to learn to solve long-horizon decision-making problems in complex robotics domains given low-level skills and a handful of short-horizon demonstrations containing sequences of images.
+To this end, we focus on learning abstract symbolic world models that facilitate zero-shot generalization to novel goals via planning.
+A critical component of such models is the set of symbolic predicates that define properties of and relationships between objects.
+In this work, we leverage pretrained vision-language models (VLMs) to propose a large set of visual predicates potentially relevant for decision-making, and to evaluate those predicates directly from camera images.
+At training time, we pass the proposed predicates and demonstrations into an optimization-based model-learning algorithm to obtain an abstract symbolic world model that is defined in terms of a compact subset of the proposed predicates.
+At test time, given a novel goal in a novel setting, we use the VLM to construct a symbolic description of the current world state, and then use a search-based planning algorithm to find a sequence of low-level skills that achieves the goal.
+We demonstrate empirically across experiments in both simulation and the real world that our method can generalize aggressively, applying its learned world model to solve problems with a wide variety of object types, arrangements, numbers of objects, and visual backgrounds, as well as novel goals and much longer horizons than those seen at training time.
+Keywords: Planning, Learning from Demonstration, Vision-Language Models
 1 Introduction
-A core research objective in robotics is to develop a robot that can make long-term decisions from low-level sensory inputs to accomplish a very broad range of tasks.
-Recent work has shown that model-free imitation learning can solve complex tasks directly from pixels (Chi et al.
-2023; Zhao et al.
-2023).
-However, sequential and compositional generalization to novel tasks beyond the training distribution remain open challenges, especially when task horizons are long and demonstration datasets are small.
-In this work, we take an aggressively model-based approach to address these challenges.
-In particular, we use sparse demonstration data to learn abstract world models (Silver et al.
-2023; Liang et al.
-2024) and plan in those world models to achieve strong compositional and sequential generalization.
-Figure 1: Example train vs.
-test task in the “Burger” environment.
-Training tasks involve creating a single burger with a single ingredient or stacking two ingredients.
-Tasks at test time involve more objects, different object arrangements, and novel compositions of goals as compared to those at training time.
-For example, one of the tasks at test time is to create three burgers – two with multiple ingredients – and where the robot starts out holding something (no training task starts out with the robot holding something).
-Figure 2: Pix2Pred overview.
-pix2pred are given an initial set of predicates, a set of controllers (temporally extended actions like pick, cook, etc.) and demonstrations of solving tasks in the form of goal-directed observation-action sequences.
-Our method learns more predicates and learns operators in terms of the learned and given predicates.
-Given a novel observation and goal at test time, we use a planner to generate a plan with our learned operators that achieves the goal.
-Goals in the train and test tasks are defined in terms of the given predicates; goals are written as text in the figure to simplify the figure.
-As a simple, pedagogical example, consider the “Burger” domain depicted in Figure 1.
-The agent is given demonstrations showing how to cook a patty, how to cut lettuce, how to stack ingredients, and how to make a sandwich with a single ingredient.
-It is then tasked with making multiple burgers, each with a cooked patty and chopped lettuce, from novel initial images where some patties may be cooked already, some lettuce chopped already, and where the robot might start out holding an object.
-This test task is challenging for model-free imitation learning because it is much longer-horizon than the demonstrated tasks and involves novel numbers and configurations of objects.
-However, if the agent can learn to recognize a small number of abstract state features, such as when a particular patty is Cooked, when a particular head of lettuce is Chopped, and whether or not the robot is currently holding an object, – as well as abstract transition rules involving these concepts: cooking a patty causes it to go from Raw to Cooked, chopping lettuce causes it to go from Whole to Chopped, a robot can only pick up an object if it is not Holding something else – it can combine these concepts in many ways when planning and generalize substantially to new tasks.
-We are specifically interested in learning PDDL models (Fox and Long 2003; McDermott et al.
-1998) in the form of symbolic predicates (e.g., Cooked) and operators (e.g., cooking a patty causes it to go from Raw to Cooked) that can be leveraged by powerful classical planning algorithms (Helmert 2006).
-Such models feature several biases that are conducive to efficient planning and learning.
-Firstly, they are relational: world states and goals are characterized in terms of properties and relations between objects.
-Secondly, they are object-oriented: the number and properties of objects may vary across different problem instances, allowing the same model to support a wide range of specific tasks.
-Finally, they are sparse: each modeled action depends on and affects only a relatively small number of other objects, abstracting out irrelevant details and predicting core cause-effect interactions.
-Several recent works have demonstrated that such models can be learned efficiently to support generalization to tasks featuring new states, new goals, more objects, and longer time horizons than those demonstrated at training time (Silver et al.
-2021, 2022, 2023; Asai and Fukunaga 2018; Ahmetoglu et al.
-2022).
-However, missing from previous work is a method for learning predicates that can be computed from raw visual observations from very few demonstrations.
-To address this gap, our first key idea is to use Vision-Language Models (VLMs) to propose and ground predicates.
-Given demonstrations, we leverage the commonsense knowledge of the VLM propose predicates with natural language names.
-Given an invented predicate and new states, we can again use the VLM to ground and evaluate the predicate directly on images.
-However, we find experimentally that the VLM’s proposed set of predicates do not support efficient planning or generalization across training tasks.
-Given this, our second key idea is to generate a large pool of candidates and subselect predicates by explicitly optimizing for efficient and effective planning.
-We extend a recent method for program-synthesis-based predicate invention (Silver et al.
-2023) that searches over sets of predicates via hill-climbing to minimize a fast-planning objective.
-After running this procedure to subselect predicates, we obtain a compact set of predicates that supports efficient and effective downstream decision-making.
-In experiments, we evaluate the extent to which our method — pix2pred — is able to learn symbolic predicates and operators that generalize to novel problems.
-Across four tasks in two simulated environments, we compare to several related baselines and find that pix2pred consistently achieves the highest success rate on novel problem instances involving more objects, more complex goals, and longer horizons than those demonstrated during training
+A core research objective in robotics is to make long-term decisions from low-level sensory inputs to accomplish a very broad range of tasks [36].
+For instance, a useful household robot should be able to satisfy a variety of requests, such as “wipe and clear my dining table”, “make me a cup of healthy juice”, or even “cook me a protein-rich meal”, across a wide range of novel household environments.
+Even when provided with a library of chainable short-horizon parameterized skills (such as moveto, pick, place, wipe, turn-knob), solving such tasks remains challenging.
+The robot must not only select and sequence the right skills, but also ground each one with the correct parameters—identifying which objects to act upon, where to move, how to grasp, where to place, or how far to turn a knob.
+Critically, the space of possible variations in objects, configurations, visual backgrounds, and goal specifications is extremely large: even seemingly simple instructions can require substantially different skill sequences and parameters in different situations.
+In this work, we choose to tackle these challenges in a learning-from-demonstration setting with an aggressively model-based approach.
+Given a handful of demonstrations that use known parameterized skills, we learn symbolic world models [42, 30] composed of abstract, interpretable properties and object relations (i.e., predicates) that ground directly into low-level sensory inputs (i.e., pixels).
+Because these models capture task-agnostic world dynamics, they can generalize across objects, embodiments (e.g., from human to robot), and tasks (e.g., from simple settings at training time to more complex, novel settings at test time).
+Additionally, the structure of the models affords efficient search-based task planning [33, 10] to solve novel tasks at inference time [14].
+Figure 1: Overview of pix2pred in the Cleanup domain: Given 66 human demonstrations showcasing the effects of distinct skills (e.g., wiping, dumping) and a small initial predicate set, pix2pred invents new predicates (e.g., NoObjectsOnTop(?table)) and learns symbolic operators.
+At test time, it uses a search-based planner that operates over the learned model to solve a novel multi-step task in a visually distinct environment—e.g., retrieving an eraser from a bin, clearing an obstacle, wiping the table, and returning the eraser.
+As an illustrative setting, consider the Cleanup domain shown in Figure 1.
+At training time, the robot receives a few egocentric visual demonstrations of a human sequencing available skills — such as grasping, dumping, and wiping — to accomplish two specific goals: wiping a table, and throwing away a particular object.
+At inference time, it must achieve a novel goal of wiping, and putting away the eraser afterward in an entirely different room where the eraser is initially in a bin (and thus the robot must first dump the bin before using the eraser and then putting it back in the bin).
+This sequencing is challenging for model-free imitation learning as it requires producing an entirely different action sequence in a scenario with novel objects, and backgrounds.
+However, if the agent can learn and reason over a small set of conceptual distinctions in the world state, including whether an object is an eraser or a surface is clear, then it can compose these via planning to solve unseen tasks.
+Learning predicates for symbolic world models [25, 42, 13, 30] in this setting is challenging for two reasons.
+First, we are provided no initial information or explicit constraints on the predicates to be learned: we must decide the number of predicates, the types and numbers of objects they should operate over, as well as their implementation (i.e, how they process input visual information).
+Second, we must invent predicates that lead to a symbolic world model that generalizes well.
+There are an infinite number of possible world models that could describe the dynamics illustrated by a set of demonstrations: we wish to select one that enables efficient planning across the entire distribution of tasks the demonstration tasks are drawn from.
+To address these challenges, we introduce a new predicate-invention method — pix2pred — that exploits the perceptual grounding and common-sense capabilities of Vision-Language Models (VLMs) within a program synthesis framework.
+Our method assumes as input demonstrations annotated by intuitively-named parameterized skills and objects and a small set of initial predicates primarily used to describe the demonstrations’ goals.
+From these, it generates a diverse pool of candidate predicates — both programmatic, feature-based predicates (as in [42]) and VLM-derived, pixel-grounded predicates that capture high-level, semantically meaningful concepts (e.g., NoObjectsOnTop(?table)).
+Given this pool, we run a subselection procedure that learns symbolic operators and selects the predicate and operator set that optimizes an efficient planning objective [42].
+This generate-then-subselect procedure allows pix2pred to filter redundant or unreliable predicates and retain only those that accurately ground in pixels and are useful for downstream decision-making.
+In experiments, we find that pix2pred is able to learn symbolic world models that operate over pixels from just a handful (<15<15) of egocentric human demonstrations (Figure 1).
+These models enable a real-world Boston Dynamics Spot robot to solve two complex, long-horizon tasks with different layouts, objects, and visual appearances.
+We leverage three simulated domains developed in different previous works to rigorously benchmark our approach’s performance against several ablations and baselines from the literature.
+From this, we find that pix2pred achieves the highest success rates on novel problem instances involving more objects, more complex goals, and longer horizons.
 2 Related Work
 Decision-Making with Foundation Models.
-Our work is strongly inspired by impressive recently-demonstrated capabilities of Large (Vision) Language Models (LLMs, VLMs) across a variety of challenging text and image tasks (OpenAI 2023; Team et al.
-2023; Dubey et al.
-2024).
-We build on a large body of work that leverages foundation models for decision-making in robotics tasks (Ichter et al.
-2023; Singh et al.
-2023; Liang et al.
-2022; Huang et al.
-2022; Curtis et al.
-2024; Huang et al.
-2023, 2024; Hu et al.
-2023a; Duan et al.
-2024; Yang et al.
-2024) (see Hu et al.
-(2023b) for a recent survey).
-However, many of these approaches operate under a number of different restrictive assumptions.
-Some are confined to a particular task distribution (e.g., only pick-and-place) or seek to synthesize new relatively short-horizon skills (e.g., pouring a liquid into a cup) (Huang et al.
-2023, 2024; Duan et al.
-2024).
-Others assume skills are provided, and attempt to use foundation models to compose them to solve longer-horizon tasks (Ichter et al.
-2023; Singh et al.
-2023; Liang et al.
-2022; Huang et al.
-2022; Curtis et al.
-2024; Hu et al.
-2023a; Quartey et al.
-2024; Kumar et al.
-2024a).
-These approaches all use a foundation model directly for planning, which has been shown to perform significantly worse than dedicated planning algorithms that use an appropriate abstraction (Valmeekam et al.
-2022; Kambhampati et al.
-2024).
-By contrast, our approach leverages foundation models to construct abstractions that are fed into a planning system at inference time.
+Our work is inspired by recently-demonstrated capabilities of large (vision) language models (LLMs, VLMs) in a variety of challenging text and image tasks [35, 46].
+We build on a large body of work that leverages foundation models for decision-making in robotics [20, 44, 29, 19, 7, 18, 17, 16, 9, 51, 39] (see Hu et al.
+[15] for a recent survey).
+However, these approaches operate under a number of different restrictive assumptions.
+Some are confined to a particular task distribution (e.g., only pick-and-place) or seek to synthesize new relatively short-horizon skills (e.g., pouring liquid into a cup) [18, 17, 9].
+Others assume skills are provided, and attempt to use foundation models to compose them to solve longer-horizon tasks [20, 44, 29, 19, 7, 16, 37, 27].
+We similarly assume skills are provided, but leverage foundation models to construct abstractions that are used by a search-based planning system at inference time.
 We treat the work of Hu et al.
-(2023a) as representative of previous approaches that use foundation models to plan, and compare directly to variants of this approach in our experiments (Section 5).
+[16] as representative of previous foundation model planning approaches and compare directly to variants in our experiments (Section 5).
+Figure 2: Labeling and proposal (Left) A small subset of the proposed predicates for the Cleanup domain.
+(Right) Truth values for these predicates (green is true and red is false) in a particular state, as determined by the VLM.
+This data is ultimately used to subselect the correct predicates.
 Learning Abstractions for Planning.
-We build on a long line of work that learns abstractions for efficient planning (Bertsekas, Castanon et al.
-1988; Jong and Stone 2005; Abel, Hershkowitz, and Littman 2016; Ugur and Piater 2015; Asai and Fukunaga 2018; Konidaris, Kaelbling, and Lozano-Pérez 2018; Silver et al.
-2021; James, Rosman, and Konidaris 2022; Mao et al.
-2022; Chitnis et al.
-2022; Kumar et al.
-2023; Silver et al.
-2023; Han et al.
-2024; Liu et al.
-2024).
-Several of these works assume that predicates or operators are already given (Chitnis et al.
-2021; Kumar et al.
-2023; Mao et al.
-2022; Siegel et al.
-2024) or focus on the problem of learning predicates and other abstractions from online interaction (James, Rosman, and Konidaris 2022; Liang et al.
-2024), or assume access to dense natural language descriptions with each demonstration (Liu et al.
-2024; Han et al.
-2024).
+We build on a long line of work that learns abstractions for efficient planning [3, 22, 1, 47, 2, 25, 43, 41, 21, 34, 6, 26, 42, 13, 31, 38].
+Several of these methods assume that predicates or operators are already given [5, 34, 32, 40] or focus on the problem of learning predicates and other abstractions from online interaction [21, 30], or assume access to dense natural language descriptions with each demonstration [31, 13].
 By contrast, we aim to invent predicates and operators in an offline fashion from image-based demonstrations with no supervision on the kinds or number of predicates to learn.
-Like us, “skills to symbols” (Konidaris, Kaelbling, and Lozano-Pérez 2018) performs predicate invention given skills, but operates on point cloud data and uses a bisimulation objective to select predicates that is different from the fast-planning objective that we use.
-Our work builds directly on the method from  Silver et al.
-(2023), which learns predicates and operators that operate over a handcrafted, low-dimensional feature space.
-We extend this approach to learn predicates and operators that operate directly on images, and we demonstrate experimentally (Section 5) that this is necessary to solve tasks where attributes of objects important for decision-making cannot be easily extracted and encoded into a low-dimensional feature space.
-3 Problem Setting and Background
-We consider the problem of learning from demonstrations in deterministic, fully-observed environments.
-pix2pred takes as input a set of core object-types with meaningful natural-language names, an initial set of predicates, a set of parameterized controllers, and demonstrations of solving tasks in the form of goal-directed observation-action sequences, where goals are expressed in terms of the initial predicates and the sequences are segmented by controller.
-Each observation contains raw sensor data that includes one or more images characterizing the world state at that time, as well as data from sensors on the robot.
-We assume access to a perception function that can identify basic attributes of objects (e.g., position, dimensions, etc.) from the raw sensor data and can annotate objects of the given object-types on the images with unique names.
-Each controller execution in the demonstration is labeled with its name, the names of the objects it operates on in terms of names given by the perception function, and the continuous parameters it was executed with.
-From this, pix2pred learns predicates and operators (specifically in the form of a PDDL domain file (McDermott et al.
-1998)), as well as samplers that produce the continuous parameters for the controllers.
-Below, we describe the precise structure of our demonstrations more formally.
-We also provide background on symbolic predicates, operators, samplers, and an existing framework for predicate invention that we build on.
-3.1 Demonstrations
-Our demonstrations are provided within a set of training tasks drawn from some distribution $\mathcal{T}$.
-We are ultimately interested in learning to solve any task $T$ drawn from this distribution.
-All tasks occur within a common environment.
-Environments and Tasks
-We model an environment as a tuple $\langle I,A,f,\Lambda,P\rangle$, where $I$ is the observation space, $A$ is the action space, and $f$ is an unknown transition function, denoted $f:I\times A\rightarrow I$.
-An observation $\text{obs}\in I$ consists of a sequence of one or more images ($\text{obs}^{\text{img}}$) representing camera images from cameras operated by the robot, as well as a vector of continuous real-valued numbers ($\text{obs}^{\text{feat.}}$) representing raw sensor data associated with the robot, such as GPS and encoder data.
-The agent receives some observation $\text{obs}_{t}$ at some timestep $t$, and can execute an action $a\in A$.
-The environment will then return a new observation $\text{obs}_{t+1}\leftarrow f(\text{obs}_{t},a)$.
-We assume $f$ is deterministic.
-$\Lambda$ is a finite set of object types, where each type $\lambda\in\Lambda$ has a name (e.g., robot, patty, etc.) and a tuple of real-valued features representing basic attributes of the object (e.g., position for most objects, position and joint angles for the robot).
-Within an environment, a task is a tuple $\langle\mathcal{O},i_{0},g\rangle$.
-$\mathcal{O}$ is a set of objects, each with a type from $\Lambda$.
-The environment’s perception function $P:I\times\mathcal{O}\rightarrow\mathcal{X}$ processes each observation into a structured state used for downstream decision-making.
-These states have two parts: an image-based state, ($x^{\text{img}}$), and a feature-based state, ($x^{\text{feat}}$), concatenated together to form a state $x\in\mathcal{X}$.
-The perception function gives a unique name to each object in the observations and annotates the images with their names, as shown in Figure 2; this forms the image-based state.
-In practice this can be achieved by combining open-vocabulary object-detection and segmentation models (Ren et al.
-(2024)).
-The perception function also identifies the basic attributes of each object according to its type; this forms the feature-based state.
-In practice, this can be implemented with custom detectors for each property of each object, and is simple when the attributes are basic properties like pose and dimensions.
-We assume that objects are fully-observable to the perception function (i.e., there is no observation in which certain objects or types of objects are unobservable).
-$i_{0}$ is the initial observation of the task, and $x_{0}$ represents the initial state of a task produced by the perception function, where $x_{o}\leftarrow P(i_{0},\mathcal{O})$.
-We will henceforth refer to a task as involving this initial state: $\langle\mathcal{O},x_{0},g\rangle$.
-We describe the goal $g$ in a the next section.
-Initial Predicate Set and Goals
-In order to specify the goal $g$ of a task, as well as provide a basis for decision-making and learning, we provide an initial set of predicates $\Psi_{\text{init}}$ as part of the demonstrations.
-Each predicate $\psi$ can be thought of as a template for a binary classifier over states.
-Specifically, a predicate is characterized by a name, an ordered list of types $(\lambda_{1},\dots,\lambda_{m})$ and a classifier function $c_{\psi}:\mathcal{X}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$ parameterized by typed object arguments.
+Our approach takes strong inspiration from “skills to symbols” [25] and follow-up work [21], which also aim to learn symbolic world models from a set of skills.
+James et al.
+[21] in particular learns abstractions by estimating option initiation sets and effects from interaction data, partitioning options into subgoal types, and merging objects into effect-equivalence classes to form predicates.
+In contrast, our method differs in two key ways.
+First, we leverage a vision–language model to propose and ground predicates directly from raw demonstrations, rather than inducing them through state partitioning.
+This enables learning from just a few short, embodiment-agnostic demonstrations, while biasing the representation toward human-interpretable, “commonsense” predicates (e.g., IsEraser(), IsColorGreen).
+Second, instead of relying on effect-equivalence criteria for learning predicates, we explicitly optimize predicate selection with a planning-efficiency objective, ensuring that the learned world models are directly tailored for effective long-horizon planning.
+Our work builds directly on the bilevel planning framework of Silver et al.
+[42], which learns predicates, operators, and samplers compatible with Task and Motion Planning (TAMP) [11].
+However, unlike Silver et al., which assumes a handcrafted, noiseless object-centric state space, our approach introduces algorithmic modifications (e.g., soft precondition intersection) to handle the noise inherent in VLM-generated visual predicates.
+Closely related concurrent works, VisualPredicator [30] and SkillWrapper [52], also leverage VLMs for predicate invention but are unable to learn from offline embodiment-agnostic demonstrations.
+In particular, these approaches typically require the agent to execute exploratory skills and learn from contrastive signals between successful and failed skill or plan executions.
+Our method targets the learning-from-demonstration setting: it learns entirely from a small set of offline, embodiment-agnostic demonstrations (e.g., human videos) that contain only positive skill executions.
+This allows us to acquire models without requiring online interaction or environment resets, enabling us to validate our approach on real-world tasks with a Boston Dynamics Spot robot (Section 5) trained using human video demonstrations.
+Ultimately, we view our method as complementary to these online approaches [30, 52]: pix2pred could be used to initialize a world model from offline data, which is subsequently refined into a provably sound and complete model online by these methods.
+3 Background and Problem Setting
+In this section, we describe how we model states, actions, and objects, and then provide background on predicates and symbolic world models.
+We also describe the inputs and outputs of the training (learning) and inference (deployment) phases of our method.
+3.1 States, Actions, and Objects
+We wish to perform decision making given an image-based state $s_{t}$ at timestep $t$ given a set of user-specified objects $\mathcal{O}$.
+Each object $o\in\mathcal{O}$ consists of a name, an optional type (denoted $\lambda$), and a descriptor string.
+If a type is specified, it is associated with a continuous feature vector composed of object properties (e.g.
+pose, size).
+Following previous work [42, 28], we assume that the agent is equipped with feature detectors (e.g.
+pose estimation within a pre-specified map) corresponding to each of these properties.
+Descriptions are short user-provided natural-language phrases (such as “blue mop” or “small round coffee table”) that, in the context of a particular set of images, suffice to designate a specific object.
+Importantly, we assume throughout that these descriptors enable clear disambiguation and identification of the particular object within the image-based state, and the corresponding objects are present in the initial state.
+The state $s_{t}$ itself is comprised of a set of images: $s^{\text{img}}_{t}=\{i^{0}_{t},i^{1}_{t},\ldots,i^{j}_{t}\}$ (potentially captured from multiple cameras) and an “object-centric” state $s^{\text{obj}}_{t}$, which includes the set of continuous feature vectors (e.g.
+pose, size) for every unique object with corresponding type that is in the images.
+Overall, $s_{t}\leftarrow s^{\text{img}}_{t}\mathbin{\|}s^{\text{obj}}_{t}$, where $\mathbin{\|}$ denotes concatenation.
+Thus, the overall state space $S$ is composed of an image-based state space, and a feature-based state space.
+AA is the agent’s action space defined by a finite set of skills $\mathcal{C}$ that are temporally-extended low-level robot behaviors.
+Each skill $C((\lambda_{1},\dots,\lambda_{v}),\Theta)\in\mathcal{C}$ has a semantically-meaningful name, a policy function, optional discrete typed object parameters $(\lambda_{1},\ldots,\lambda_{v})$, and optional continuous parameters $\Theta$.
+For instance, a skill Grasp for picking up an object might have one discrete parameter of type object and a 6-dimensional $\Theta$ that is a placeholder for a specific grasp.
+An action $a\in\mathcal{A}$ is a skill $C\in\mathcal{C}$ with discrete and continuous arguments fully specified: $a=C((o_{1},\ldots o_{v}),\theta)$.
+An action executes an underlying policy from the policy function (e.g.
+a learned grasp policy) to modify the state.
+Figure 3: Domains. Top row: train task example illustrations. Bottom row: test task example illustrations.
+3.2 Predicates, Symbolic World Models, and Planning
+A predicate is characterized by a name, an ordered list of $m$ arguments $(\lambda_{1},\dots,\lambda_{m})$ (note that the types of the arguments are not necessarily unique).
+and a classifier function $c_{\psi}:S\times\Lambda^{m}\to\{\text{true},\text{false}\}$ parameterized by the arguments.
 A ground atom $\underline{\psi}$ consists of a predicate $\psi$ and objects $(o_{1},\dots,o_{m})$.
-Each ground atom induces a particular binary state classifier $c_{\underline{\psi}}:\mathcal{X}\to\{\text{true},\text{false}\}$, where $c_{\underline{\psi}}(x)\triangleq c_{\psi}(x,(o_{1},\dots,o_{m}))$.
-For example, the ground atom Cooked(patty0) would have the value “true” if the object patty0 is cooked.
-[^1] It would have this value if its classifier was accurate.
-We discuss what happens if this classifier is inaccurate in Section 4.3.
-We assume the initial predicates $\Psi_{\text{init}}$ are sufficient for representing task goals, and may additionally contain concepts useful across all tasks in the environment, but are insufficient on their own in that additional predicates will be necessary to capture all concepts critical for solving tasks.
-Specifically, the goal $g$ of a task is a set of ground atoms involving predicates in $\Psi_{\text{init}}$ and objects in $\mathcal{O}$.
-A goal $g$ is said to hold in a state $x$ if for all ground atoms $\underline{\psi}\in g$, the classifier $c_{\underline{\psi}}(x)$ returns true.
-We use the term “goal predicates” for predicates that are used in defining the goal.
-We provide a discussion about the necessity of predicate invention despite having these predicates in Appendix A.4.
-Actions as Hybrid Controllers
-The environment’s action space $A$ is defined by a finite set of controllers $\mathcal{C}$.
-Each controller $C((\lambda_{1},\dots,\lambda_{v}),\Theta)\in\mathcal{C}$ has a semantically-meaningful name, as well as optional discrete typed parameters $(\lambda_{1},\ldots,\lambda_{v})$ and continuous parameters $\Theta$.
-For instance, a controller Pick for picking up an object might have one discrete parameter of type item and a $\Theta$ that is a placeholder for a specific grasp.
-An action $a\in\mathcal{A}$ in a task is a controller $C\in\mathcal{C}$ with discrete and continuous arguments fully specified: $a=C((o_{1},\ldots o_{v}),\theta)$.
-Final Demonstration Structure
-We provide the agent with a demonstration set $\mathcal{D}$.
-Each demonstration consists of: (1) a training task $T\sim\mathcal{T}$, (2) a near-optimal plan that solves the task, and (3) the state sequence of that plan with values for all initial predicate ground atoms.
-Recall that a task is a tuple $\langle\mathcal{O},x_{0},g\rangle$.
-A plan $\pi=(a_{1},\ldots,a_{m})$ is a sequence of actions $a\in\mathcal{A}$ such that successive execution of the actions in the environment, starting from $x_{0}$, results in a final state $x_{m}$ where $g$ holds.
-A state sequence of a plan is the set of states $(x_{1},\ldots,x_{m+1})$ encountered as that plan is executed in an environment.
-An example of this input is shown in Figure 2.
-Given a handful (less than $15$) of these demonstrations, the agent’s objective is to invent additional predicates and a set of operator definitions in terms of the initial and learned predicates that enables the agent to plan and efficiently solve a set of held-out tesk tasks drawn from $\mathcal{T}$.
-Test tasks involve more objects, different object configurations, novel compositions of goals, and longer-horizons than the training tasks solved in the demonstrations.
-3.2 Operators and Planning
-Predicates $\Psi$ induce an abstract state space $S_{\Psi}$ of the task’s underlying continuous state space $\mathcal{X}$. Operators help specify an abstract transition model over $S_{\Psi}$.
-More specifically, given a particular state $x_{t}$, as well as predicates $\Psi$, we can compute the corresponding abstract state (denoted by $s_{t}$) by selecting all the ground atoms that are “true” in $x_{t}$.
-Formally, $s_{t}=\textsc{Abstract}(x_{t},\Psi)\triangleq\{\underline{\psi}:c_{\underline{\psi}}(x_{t})=\text{true},\forall\psi\in\Psi\}.$ Then, each operator $\omega$ is a tuple $\omega=\langle\overline{v},P,E^{+},E^{-},\mathcal{C}\rangle$.
+Each ground atom induces a particular binary state classifier $c_{\underline{\psi}}:S\to\{\text{true},\text{false}\}$, where $c_{\underline{\psi}}(x)\triangleq c_{\psi}(x,(o_{1},\dots,o_{m}))$.
+For example, the ground atom Holding(spot, apple) would have the value “true” if the apple is being held by the spot.
+In this work, we distinguish between feature-based predicates that operate exclusively over the object-centric state, and visual predicates that operate on the image-based state.
+A set of predicates $\Psi$ induce an abstract state space $S^{\Psi}$ of the underlying continuous state space $S$.
+In order to enable planning, we require an action model $\Omega$ that specifies an abstract transition model over $S^{\Psi}$.
+We learn such an action model in the form of symbolic PDDL-style [33, 10] operators that are each linked to a particular skill from the domain’s action space.
+In domains where skills have continuous parameters, each operator is associated with a generative sampler that proposes continuous parameters for the associated skill [6].
+Planning.
+Together, the predicates and operators specify a symbolic abstract world model that can be used in conjunction with samplers for efficient hierarchical planning such as Task and Motion Planning (TAMP) [45, 11].
+In this work, we choose to use a simpler planning procedure that does not require a precise simulator or transition model [28].
+At a high level, our planning procedure takes in the initial state $s_{0}$ and first evaluates all the provided predicates $\Psi$ to yield an abstract state.
+Then, it calls a PDDL planner to achieve the goal from this abstract state.
+We can simply execute each action of this plan directly and check whether the goal is achieved at the end (which we do in all our simulation domains), or execute actions one-by-one and replan from resulting states as necessary (which we demonstrate in some of our real-world deployment).
+The precise details of our planning procedure are not necessary to understand what follows; we defer them to Appendix A.2.
+3.3 Learning and Deployment
+Learning Phase.
+We learn from a set of $n$ demonstrations $\mathcal{D}$, starting with a set of initial predicates $\Psi^{\text{init}}$.
+Each demonstration $d$ consists of (1) a set of objects $\mathcal{O}^{d}$, (2) a k-step trajectory $\tau^{d}=[s_{0},a_{0},s_{1},a_{1},\ldots,s_{k}]$ of states and actions, and (3) a conjunctive goal expression $g^{d}$ that holds true only in the final state $s_{k}$.
+Here, $g^{d}$ is a conjunction of ground atoms with predicates from $\Psi^{\text{init}}$, and objects from the object set $\mathcal{O}^{d}$.
+The initial predicate set $\Psi^{\text{init}}$ consists of predicates that are necessary to specify the demonstration goals (we denote this subset $\Psi^{G}$), and optionally some additional predicates that might be known ahead of time.
+As a concrete example, for the Cleanup domain in Figure 1, the initial predicate set consists of {WipedOffTable(?obj), InsideContainer(?obj1, ?obj2), Holding(?robot, ?obj), HandEmpty(?robot)}, the first demonstration has the goal WipedOffTable(play_table), and the second has the goal InsideContainer(green_block, cardboard_box).
+Deployment Phase.
+Given (1) a new set of objects $\mathcal{O}^{\text{test}}$ with the same types as seen in the demonstrations, (2) a novel initial state $s_{0}$, and (3) a novel goal $g^{\text{test}}$ expressed using $\Psi^{\text{init}}$, we must output a sequence of $m$ actions (which we call a plan) that achieves some new state $s_{m}$ where $g^{\text{test}}$ holds.
+We refer to the tuple of objects, initial state, and goal as a task.
+4 Learning Symbolic World Models
+Given a demonstration set $\mathcal{D}$ (Section 3.3), we learn a PDDL-style [33, 10] symbolic world model $\mathcal{W}^{\mathcal{D}}$ by inventing additional predicates $\Psi^{\mathcal{D}}$, learning operators defined by both the newly-invented predicates and the initial predicates $\Psi^{\mathcal{D}}\cup\Psi^{\text{init}}$, and, where applicable, training generative samplers for continuous skill parameters [6].
+We adapt prior work for sampler learning (Appendix A.4) and focus on inventing visual predicates and operators from demonstrations.
+4.1 Proposing an Initial Pool of Visual Predicates
+The first step of our approach is to generate a pool of candidate visual predicate names by prompting a VLM on each demonstration $d\in\mathcal{D}$.
+For a demonstration $d\in\mathcal{D}$ of length $k=len(d)$, we extract the image-based state at each timestep $s^{\text{img}}_{t},t\in[0,k-1]$ (i.e., the states before and after each skill execution).
+We add a text heading to each image corresponding to which timestep $t$ in the demonstration it belongs to.
+We then pass all these images in sequence, along with description of the actions (name and specific arguments) executed in between them ($a_{0},a_{1},\ldots,a_{k-1}$), and the descriptors of the objects in each demonstration, directly to a VLM.
+We prompt the VLM to output a set of proposals (in the form of natural language strings) for ground atoms based on these inputs.
+Given a set of ground atoms proposed for each demonstration, we then automatically parse this set to discover atoms that are syntactically incorrect.
+In particular, we remove atoms that include object names that are not part of the demonstration $d$ (for instance, the VLM might propose an atom Inside(apple, room1), though there is no object named room1).
+We then deterministically lift each of these atoms into a visual predicate by creating typed variables for each object (for instance, an atom OnTop(apple, short_round_coffee_table) would be lifted to OnTop(?m: movable, ?t: table), where movable and table are input types specified as part of specifying the objects as defining the problem and environment).
+Finally, we remove any duplicate lifted atoms and add only unique predicates to our pool.
+We obtain a final pool of predicates $\Psi_{\text{pool}}^{\text{vis}}$ by running this procedure on each demonstration.
+4.2 Implementing Visual Predicates with a VLM
+We implement the classifier function $c_{\psi}$ for all visual predicates $\psi^{\text{vis}}$ as a query to a VLM.
+Specifically, we create a prompt string $\text{txt}_{\psi}$ with each predicate’s name and arguments (e.g.
+NoObjectsOnTopTable(?surface)).
+When the predicate is ground, we substitute the descriptors of the specific objects used to ground it into the prompt (e.g.
+NoObjectsOnTopTable (short_round_coffee_table)).
+To evaluate the predicates, we pass the prompt with these ground atom strings into a VLM with additional instructions, asking it to output either “true”, “false”, or “unknown” (which we take to be equivalent to “false”) for each predicate.
+If the current state is not the initial state (i.e., for $s_{t}:t>0$) of a task, we also provide the previous state’s image-based state, ground atom labeling, and executed action.
+We discuss our approach to prompting for atom labeling and proposal in detail and provide examples of our prompts in Appendix A.5.
+4.3 Learning Symbolic World Models via Optimization
+Algorithm 1 shows the pseudocode of our hill-climbing procedure for learning predicates and operators.
+Here, $\Psi^{\text{vis}}_{\text{pool}}$ is generated by the procedure described in Section 4.1.
+Where possible (such as in simulation domains where object features are available), we combine these with additional “feature-based” predicates that are generated from a programmatic grammar over the object-centric state space, as described in Silver et al.
+[42]), to create an overall pool $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$.
+The feature-based predicates are intended to capture relations that are challenging to capture visually (e.g.
+relations involving robot joint angles).
+$\Psi_{\text{pool}}$ is passed into the hill-climbing procedure from [42] that optimizes a planning-based objective ($J(\ldots)$).
+Internally, it performs operator learning (OperatorLearn) and outputs operators in terms of a small subset of predicates from $\Psi_{\text{pool}}$.
+Input: Demos $\mathcal{D}$, initial predicates $\Psi_{\text{init}}$
+$\Psi_{\text{pool}}\leftarrow\Psi_{\text{init}}\cup\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$;
+$\Psi_{\text{selected}}\leftarrow\emptyset$, $J_{\text{prev}},J_{\text{curr}}\leftarrow\infty$;
+while $J_{\text{prev}}-J_{\text{curr}}>J_{\text{thresh}}$ do
+ $J_{\text{prev}}\leftarrow J_{\text{curr}}$;
+ $\psi^{*}\leftarrow\arg\min_{\psi\in\Psi_{\text{pool}}}J(\mathcal{D},\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}}\cup\{\psi\}),\Psi_{\text{selected}}\cup\{\psi\})$;
+ $J_{\text{curr}}\leftarrow J(\mathcal{D},\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}}\cup\{\psi^{*}\}),\Psi_{\text{selected}}\cup\{\psi^{*}\})$;
+ $\Psi_{\text{selected}}\leftarrow\Psi_{\text{selected}}\cup\{\psi^{*}\}$, $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}\setminus\{\psi^{*}\}$;
+return predicates $\Psi_{\text{selected}}\cup\Psi^{G}$, $\text{OperatorLearn}(\mathcal{D},\Psi_{\text{selected}})$
+Algorithm 1 Model Learning
+We make a few core modifications to the version of Algorithm 1 from [42].
+We regularize the optimization by early-stopping the hill-climbing optimization via a hyperparameter ($J_{\text{thresh}}$) that we tune on a small validation set.
+Noise also tends to adversely impact operator learning: small differences in atom values before and after a particular skill is executed lead to learning many operators that are each overfit to specific cases, instead of one general operator that handles many transitions.
+We combat this by modifying the precondition induction procedure of operator learning to be more noise aware, as well as simply deleting operators that model less than a certain fraction of the data (controlled by a hyperparameter).
+See Appendix A.3 for more details.
+A key advantage of our approach is that the optimization will automatically select among several closely-related and even synonymous predicates (such as NextTo vs.
+Touching in Figure 2) to find the predicates that are most accurately labeled and also important for planning across the training demonstrations.
+5 Experiments
+We design experiments to answer the following questions.
+  •   Q1.
+    How well does pix2pred generalize to novel, more complex goals, especially when compared to an imitation approach that doesn’t use a planner?
+  •   Q2.
+    How critical is it to perform explicit score-based optimization for subselection of predicates?
+Domains.
+We now describe our experimental domains and tasks (illustrated in Figure 3), with details in Appendix A.6.
+We implement 2 real-world tasks on a Boston Dynamics Spot robot in different real-world buildings.
+For the purposes of extensive comparison against baselines, and to demonstrate the broad applicability of our method across sequential decision-making problems, we use 3 simulated domains that were introduced by previous work and lightly adapted for our setting.
+  • $\bullet$
+    Cleanup: A real-world Spot clears objects into a bin (sometimes dumping the bin) and wipes a tabletop. Testing tasks vary objects, backgrounds, and require more complex compositions of skills.
+  • $\bullet$
+    Juice: A real-world Spot operates a cold-press juicer by setting up cups, loading fruit, and running the machine.
+Testing tasks introduce new objects, backgrounds, and goals requiring demonstration composition.
+  • $\bullet$
+    Kitchen: Originally introduced by Gupta et al. [12]: a robot must boil water by placing a kettle on a stove burner and turning it on. Testing tasks involve using a different burner.
+  • $\bullet$
+    Burger: A simulated burger-making grid world [48].
+We test three scenarios: Bigger Burger tasks require stacking extra patties, More Burger Stacks tasks involve assembling multiple burgers instead of one, and Combo Burger tasks demand combining multiple different ingredients at test time without seeing combining at training time.
+Test tasks involve sequences of up to 30 skill executions
+  • $\bullet$
+    Coffee: A simulated task [41] where a robot dispenses coffee into a jug from a coffee machine and pours coffee into cups of varying sizes. Test tasks add more cups and novel object poses.
+Figure 4: pix2pred versus baselines in simulation. Percent of evaluation tasks solved across all our simulated domains. All results are averaged over 5 seeds. Black bars denote standard deviations.
+Approaches. We evaluate the following ablations and baselines in all simulated domains.
+  • $\bullet$
+    Ours: Our full main method: pix2pred.
+  • $\bullet$
+    VLM subselect: Ablation where the VLM directly selects a compact set of visual predicates ($\Psi^{\text{vis}}$); feature-based predicates ($\Psi^{\text{feat}}$) are still selected via hill-climbing.
+  • $\bullet$
+    No subselect: Ablation with no hill-climbing; selects all predicates in $\Psi_{\text{pool}}$.
+  • $\bullet$
+    No invent: Ablation with no new predicates beyond $\Psi_{\text{init}}$.
+  • $\bullet$
+    No visual [42]: Ablation inventing only feature-based predicates ($\Psi^{\text{feat}}$), equivalent to.
+  • $\bullet$
+    No feature: Ablation inventing only visual predicates.
+  • $\bullet$
+    VLM feat. pred [16]: Baseline where a VLM is used to directly output predicates expressed as arbitrary code functions over $S^{\text{obj}}$ [13].
+  • $\bullet$
+    ViLa: Baseline where a VLM plans directly without abstractions; we test ViLa-pure (original version) and ViLa-fewshot (novel variant that adds demonstrations to the few-shot prompt).
+Experimental Setup.
+In simulation, we vary object counts and locations across $5$ seeds, sample $10$ test tasks per seed, and allow one plan execution per task.
+In real-world domains, all demonstrations are collected from the point-of-view of a human, no object-centric state is available in the training data, and learned world models are evaluated on a robot.
+We compare our method to ViLA-fewshot using a fixed training set ($6$ demos for Cleanup, $10$ for Juice) and measure success as whether a correct plan is generated from a given initial state across $5$ generalization-focused test tasks.
+GPT-4o [35] is used as the VLM in simulation, and Gemini-2.0-flash [46] in the real world.
+All approaches use the same set of underlying skills in each domain.
+Specific details, including the precise prompts, skill implementations, and hyperparameters used can be found in Appendix A.6.
+Task Approach New New More Novel Novel
+obj. vis. obj. goal 1 goal 2
+Cleanup Ours 3/3 3/3 3/3 3/3 3/3
+ViLA-fewshot 3/3 3/3 2/3 0/3 3/3
+Juice Ours 2/3 2/3 3/3 2/3 1/3
+ViLA-fewshot 0/3 2/3 3/3 0/3 0/3
+Table 1: Real-world generalization results.
+We evaluate on 5 different tasks testing different aspects of generalization in our 2 real world domains after training on the exact same dataset.
+Success out of 3 random initial states is reported.
+Additional details on task construction and evaluation are in Appendix A.6.
+Results and Analysis.
+Figure 4 shows success rates in simulated domains.
+pix2pred outperforms prior approaches on 4 of 5 domains tested, often by a wide margin, and generalizes far better to complex tasks than ViLa few-shot (Q1).
+On the short-horizon kitchen task (2 actions), ViLa with demonstrations performs comparably to pix2pred, but its performance drops sharply on longer-horizon burger tasks (15+ actions).
+We find that ViLa tends to pattern-match training demonstrations and struggles with test tasks requiring true generalization—for example, handling initial states where the agent starts out holding an object (seen in $\sim$50% of Burger test tasks but never in training).
+ViLa fails to put down objects before picking up new ones, while pix2pred invents a HandEmpty predicate that enables the planner to handle such cases.
+Our method also outperforms ViLA in real-world domains (Table 1).
+ViLA handles new backgrounds and object instances and quantities to some extent but struggles with novel goals requiring composition across demonstrations, especially in the complex Juice domain.
+In contrast, pix2pred generalizes across object instances, visual backgrounds, object counts, and novel goal compositions with just $6$ and $10$ human-collected demonstrations.
+Qualitatively, pix2pred invents few but meaningful predicates (e.g., OnFloor and IsEraser in Cleanup; JuiceMachineOpen, InsideWasteValveRegion, and NearJuiceValve in Juice).
+See Appendix A.8 for additional details.
+We also find that explicit predicate subselection significantly improves test performance over direct VLM predicate selection (Q2).
+The VLM subselect baseline consistently underperforms pix2pred across all four task distributions.
+Qualitatively, VLM subselect picks many unnecessary predicates.
+In the Burger domain, for instance, it selects useful concepts like Cooked(?p1:patty) and Chopped(?l1:lettuce), but also inconsistent and redundant ones like Raw(?p1:patty), Available(?p1:patty), and Whole(?l1:lettuce).
+This bloats the operator set, causing overfitting and poor generalization.
+Finally, the “no subselect” baseline fails entirely, showing that subselection is critical.
+In most of our domains, the VLM proposes over $100$ distinct predicates; trying to learn operators over these directly leads to severe overfitting.
+There are two possible causes of failure for our approach: (i) pix2pred learns a correct model, but fails to make a correct plan in test tasks because the VLM incorrectly classifies one or more predicates at test-time, and (ii) pix2pred fails to learn a model that generalizes to planning on test tasks (despite predicate values being “correct”).
+Upon manual inspection we found that all failures in our real-world environments, and the majority of failures in simulation environments, are due to (i).
+6 Conclusion
+We proposed pix2pred, a method for inventing symbolic world models capable of enabling planning directly from input pixels.
+We found that our approach, from just a handful of demonstrations (3-12), is able to invent semantically meaningful predicates that afford efficient planning and aggressive generalization to novel object instances, numbers, task horizons, and goals across a range of problems.
+7 Limitations
+There are several important limitations of pix2pred in its current form.
+First, a key assumption of our approach is that, for each new task, all relevant objects in the scene are specified with unambiguous descriptors.
+This is currently done manually for the real-world environments and can be time-consuming.
+A second assumption is the full observability of objects in all states.
+These assumptions do not hold in all tasks of interest, especially when deploying in novel real-world environments.
+Future work could attempt to extract descriptors automatically from foundation models via a form of automatic prompt optimization [23], and extend the core ideas of pix2pred to partially observable settings.
+A further limitation lies in the hill-climbing optimization procedure (Section 4.3) at the core of our model learning approach: it can be extremely slow — taking many seconds to evaluate the objective value of a single candidate predicate set — especially as the number of demonstrations and size of the initial predicate pool grow.
+We also found it to be quite sensitive to the choice of hyperparameters.
+Future work could develop more efficient and noise-tolerant optimization algorithms for predicate selection, as well as improved objectives for them.
+Relatedly, operator learning algorithms could be improved to better handle large input predicate sets and noisy predicates.
+Finally, our approach assumes input demonstrations are segmented in terms of a provided set of parameterized skills with names that correspond to their function.
+These skills, as well as their underlying policies, are assumed to be provided ahead of time.
+Future work could integrate pix2pred with low-level skill-learning approaches, enabling learning from very low-level robot demonstrations or potentially even from large-scale internet video corpora.
+Overall, we hope our work inspires future progress toward learning structured and interpretable world models that enable efficient inference in and strong generalization across complex robotic manipulation problems.
+Appendix A Appendix
+A.1 Symbolic Operators
+Given a particular state $s_{t}$, as well as predicates $\Psi$, we can compute the corresponding abstract state (denoted by $s^{\Psi}_{t}$) by selecting all the ground atoms that are “true” in $s_{t}$.
+Formally, $s^{\Psi}_{t}=\textsc{Abstract}(s_{t},\Psi)\triangleq\{\underline{\psi}:c_{\underline{\psi}}(s_{t})=\text{true},\forall\psi\in\Psi\}.$ Then, each operator $\omega$ is a tuple $\omega=\langle\overline{v},P,E^{+},E^{-},\mathcal{C}\rangle$.
 Here, $\overline{v}$ are typed variables representing the operator’s arguments, $P$, $E^{+}$ and $E^{-}$ are sets of predicates representing operator preconditions, add effects and delete effects respectively, and $\mathcal{C}$ is a controller associated with the operator.
@@
 To extract an action $a$ from this ground controller, we must specify values for the parameters $\Theta$ associated with controller $\mathcal{C}$.
-Following previous work (Silver et al.
-2023; Kumar et al.
-2023; Chitnis et al.
-2022), we leverage continuous parameter samplers to do this.
-Given these components, we can perform planning to solve test tasks.
-We assume access to a planner that takes a set of predicates, operators, and samplers, as well as a task $T=(\mathcal{O},x_{0},g)$ associated with an environment, and outputs a plan that achieves the goal at the symbolic level, if such a plan exists.
-We implement such a planner via off-the-shelf classical planning algorithms (Helmert 2006), with details in Appendix A.1, though we note that our predicate invention framework is compatible with other planning strategies.
-3.3 Predicate Invention Problem
-Given demonstrations $\mathcal{D}$ structured as described in Section 3.1, our objective is to learn additional predicates ($\Psi$), operators ($\Omega$), and samplers ($\Sigma$) that enable aggressive generalization to novel test tasks from the task distribution $\mathcal{T}$ by calling a planner with these learned components.
-We directly leverage an existing approach to learn neural network samplers from our demonstration data (Appendix A.3).
-We are mainly interested in learning predicates and operators.
-Predicate Invention via Hill-Climbing
-We build on the framework introduced by Silver et al.
-(2023) for predicate invention.
-Their method first generates an initial pool of candidate predicates $\Psi_{\text{pool}}$ from a grammar and then selects the subset of predicates from this pool that maximizes an objective function designed to measure planning efficiency over the demonstrations, where the selection is done via a hill-climbing optimization.
-Predicates subselected from the pool $\Psi_{\text{pool}}$ are added to the initial set $\Psi_{\text{init}}$ to form a final set of predicates $\Psi$.
-After, operators are learned in terms of the predicates $\Psi$.
-Intuitively, optimizing their particular objective fuction yields the predicate set (and corresponding operator set) that produces plans that most closely align with the demonstrations on the training tasks.
-Importantly, this previous work makes two strong assumptions that do not hold in our setting: (1) the state space $\mathcal{X}$ is composed entirely of continuous object features (i.e., $\mathcal{X}$ only consists of $x^{\text{feat}}$), and (2) learned predicate classifiers are not noisy.
-We relax these assumptions and enable predicates that operate over images.
-4 Inventing Predicates from Pixels
-Our main interest is to invent predicates from pixels (i.e., predicates that operate directly over one or more images in our image-based state $x^{\text{img}}$).
-We wish to do this from a handful of structured demonstrations described in Section 3.1.
-These demonstrations do not possess any direct supervision for additional predicates (i.e., the demonstrations provide no indication of the number or structure of any predicates we should invent).
-The hill-climbing framework for predicate invention (Section 3.3) is able to learn predicates in such a setting, but requires a pool of initial candidate predicates to subselect from.
-To define such an initial pool, we must specify two things: what the predicates should be (i.e., what visual concepts are relevant for decision-making), and how should they be implemented (i.e., for every predicate $\psi$ in the pool, how can we define a corresponding classifier function $c_{\psi}$).
-We propose to use a VLM to both generate an initial pool of predicates that might be relevant to decision-making, and implement the classifier function for each predicate in the pool.
-Recent work has demonstrated that VLMs are able to answer a variety of common-sense natural language questions given one or more images (Antol et al.
-2015; Kim et al.
-2024; Majumdar et al.
-2024).
-The core intuition motivating our approach is that predicate generation and implementation correspond to two different types of questions.
-Predicate generation can be accomplished by giving the VLM access to our demonstrations, and asking “what visual concepts are changing between false and true after each action execution?”.
-Predicate classifier implementation can be accomplished by giving the VLM access to a particular image and asking it “is this particular ground predicate true or false in the current state?”.
-In the remainder of this section, we explain how exactly we implement this procedure and how we modify and leverage the hill-climbing predicate invention framework described in Section 3.3 to yield a compact subset of predicates optimized for planning.
-[^2] See Figure 8 for a detailed overview of our entire pipeline.
-4.1 Implementing Visual Predicates with a VLM
-Recall that each state $x\in\mathcal{X}$ is composed of two parts: a feature-based state $x^{\text{feat}}$ and an image-based state $x^{\text{img}}$.
-Visual predicates operate exclusively over $x^{\text{img}}$.
-Specifically, a visual predicate is a tuple of a name, a sequence of $m$ types, and a classifier function: $\psi_{\text{vis}}=\langle\text{name},(\lambda_{1},\dots,\lambda_{m}),c_{\psi}\rangle$.
-Here, the classifier function returns a boolean given an image-based state and object arguments $c_{\psi}:\mathcal{X}^{\text{img}}\times\mathcal{O}^{m}\to\{\text{true},\text{false}\}$.
-We propose to implement the classifier function $c_{\psi}$ by querying a VLM.
-Specifically, we create a prompt string $\text{txt}_{\psi}$ with the predicate’s name and arguments to be filled in by the predicate’s typed object arguments $(\lambda_{1},\dots,\lambda_{m})$ (e.g.
-Cooked(?obj1)).
-[^3] We create the prompt string $\text{txt}_{\psi}$ with the predicate’s name but other choices are possible.
-We could ask the VLM, “Is patty1 cooked?” instead of “What is the value of Cooked(patty1)?") When the predicate is ground, we substitute the specific objects used to ground it into the prompt (e.g.
-Cooked(patty1)).
-We implement the classifier function by simply passing the prompt with this ground atom string into a VLM with some additional instruction, asking it to output either “true”, “false”, or “unknown” (which we take to be equivalent to “false”).
-While this idea is straightforward, we found that several additional modifications to the text prompt $\text{txt}_{\psi}$ and images $x^{\text{img}}$ were critical to obtain accurate labels for ground atoms from the VLM.
-We discuss these in detail and provide examples of our prompts in Appendix A.5.
-4.2 Proposing an Initial Pool of Visual Predicates
-Figure 3: VLM predicate proposal example.
-We prompt (Appendix A.5) a VLM to propose a large pool of predicates that describe important concepts in the given demonstrations.
-Technically, the VLM outputs ground atoms as strings, and we parse the predicate (visualized in orange) from each ground atom.
-Shown here are proposals generated from demonstrations in the task distribution “Combo Burger” within the “Burger” environment (Section 5).
-The VLM proposes 563 predicates and we filter these into 136 “valid” predicates; the figure shows 56 of these.
-Figure 4: VLM predicate labeling example.
-We prompt (Appendix A.5) a VLM to determine the truth values of ground atoms (each represented as a string) in the given image, labeling each as either true, false, or unknown.
-Shown here are a subset of the labels generated for a particular state from a task in the task distribution“Combo Burger” within the “Burger” environment (Section 5).
-We generate a pool candidate of visual predicates by prompting a VLM on each demonstration $d\in\mathcal{D}$.
-For a demonstration $d\in\mathcal{D}$ of length $k=len(d)$, we extract the image-based state at each timestep $x^{\text{img}}_{t},t\in[0,k-1]$ and add a text heading to each image corresponding to which timestep $t$ in the demonstration it belongs to.
-We then pass all these images in sequence, along with the actions executed in between them ($a_{0},a_{1},\ldots,a_{k_{d}-1}$) directly to a VLM.
-We prompt the VLM to output a set of proposals for ground atoms based on these inputs.
-Given a set of ground atoms proposed for each demonstration, we then parse this set to discover atoms that are syntactically incorrect.
-In particular, we remove atoms that include object names that are not part of the demonstration $d$ (for instance, the VLM might propose an atom ontop(patty1, pot1), even though there is no object named pot1).
-We then lift each of these atoms into a visual predicate by creating typed variables for each of the objects identified in each proposed ground atom (for instance, an atom cooked(patty1) would be lifted to cooked(?p: patty)).
-Finally, we remove any duplicate lifted atoms and add only unique predicates to our pool.
-We obtain a final pool of predicates $\Psi_{\text{pool}}^{\text{vis}}$ by running this procedure on each demonstration.
-4.3 Predicate Invention via Subset Selection From Noisy Data
-Given the procedure described in Section 4.2, we can generate an initial pool of visual predicates $\Psi_{\text{pool}}^{\text{vis}}$.
-We combine these with feature-based predicates (generated via a programmatic grammar as described in Silver et al.
-(2023)) to create an overall pool $\Psi_{\text{pool}}\leftarrow\Psi_{\text{pool}}^{\text{vis}}\cup\Psi_{\text{pool}}^{\text{feat}}$.
-We wish to run hill-climbing optimization as described in Section 3.3 to subselect a small set of relevant predicates that are optimized for efficient and effective planning over our training tasks.
-Figure 5: Labeling of ground atoms with visual predicates involving the patty object type in Burger.
-This figure depicts the visual state before and after the cook skill in demonstration #0 for the task distribution “Combo Burger”.
-It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the patty object in these states.
-The VLM proposes a wide variety of predicates involving the patty: clear, cooked, cool, grilling, raw, and uncook.
-Of these, cooked(?p:patty) is most relevant to achieving the goal while also being accurately labeled.
-Our approach automatically selects this predicate from the pool.
-See Figure 9 in the appendix for an analogous situation with the chop skill.
-Unfortunately, we cannot directly apply the particular hill-climbing optimization from previous work because our visual predicates are noisy.
-Visual predicates might not be consistent across states in the demonstrations due to occasional hallucination or mislabeling on the part of the VLM.
-This noise creates outliers in the abstract state space transition data and causes the operator learning algorithm to overfit and create a large number of unnecessary operators to model these outlier transitions.
-These extra operators make planning inefficient by increasing the branching factor of the planner’s search algorithm.
-They also cause the hill-climbing predicate subselection – which interally performs operator learning on every subset of predicates it scores – to overfit and select a large number of unnecessary predicates.
-These overfit predicates can appear in the definitions of all operators – not just the operators created to model the outlier transition data – because operators are learned to be chainable (add effects of an earlier operator satisfy the preconditions of a later operator).
-The overfit predicates and overfit operators fail to generalize outside of the training distribution.
-We combat this problem by modifying the operator learning approach to be more tolerant to noise  [^4] Modifications are described in detail in Appendix A.2, and by regularizing the hill-climbing optimization itself via early stopping to discourage overfitting.
-5 Experiments
-Our experiments are designed to answer the following questions.
-  •       Q1.
-    How well does pix2pred generalize to novel, more complex goals when compared to an imitation approach that doesn’t use a planner?
-  •       Q2.
-    How critical is it to perform explicit score-based optimization for subselection of predicates?
-  •       Q3.
-    How critical is it to invent both visual and feature-based predicates?
-Figure 6: Environments. Top row: train task examples. Bottom row: test task examples. See Section 5 for descriptions of each environment.
-Environments.
-We now describe our experimental environments and tasks, with details in Appendix A.6.
-We implement 4 task distributions across 2 environments.
-Each environment was introduced by previous work and lightly adapted for our setting.
-  • $\bullet$
-    Kitchen: A simulated robotic environment featuring a robotic arm positioned in front of a kitchen that includes a stove, a microwave, and controllable lights (Gupta et al.
-2019).
-The task is to boil water in a kettle by turning on a specific stove burner and pushing the kettle onto it.
-The task has a short horizon, requiring only two high-level skills in sequence to achieve the goal.
-Demonstrations show the robot solving the task using a particular burner from various kettle initial positions, and at test time the robot must use a different burner on the same stove.
-  • $\bullet$
-    Burger: A simulated 2D grid world environment (based on the Robotouille environment introduced by (Wang et al.
-2023)) featuring a robot that must prepare and assemble burgers from various components.
-Preparation includes chopping lettuce on a cutting board and cooking patties on a grill.
-We experiment with three task distributions within this environment, with each distribution featuring its own set of training and testing tasks: (1) Bigger Burger involves producing a burger with extra patties after being shown how to create a burger with a single patty, (2) More Burger Stacks involves producing multiple burgers after being shown how to create a single burger, and (3) Combo Burger involves producing burgers with multiple ingredients after being shown how to create burgers with a single ingredient.
-The tasks have a long horizon.
-For example, test tasks in Bigger Burger, More Burger Stacks, and Combo Burger can require a sequence of 16, 29, and 30 controllers to achieve the goal, respectively.
-Figure 7: pix2pred versus baselines. Percent of evaluation tasks solved per task across both simulated environments. All results are averaged over 5 seeds. Black bars denote standard deviations.
-Approaches. We evaluate our method against several ablations and baselines from the literature, including those that attempt to solve test tasks directly without learning any abstractions.
-  • $\bullet$
-    Ours: Inventing visual and feature-based predicates with pix2pred, learning operators in terms of these predicates, and planning with these operators at test time.
-  • $\bullet$
-    VLM subselect: An ablation of pix2pred where we have the VLM directly subselect a compact set of visual predicates instead of delegating subselection to the hill-climbing optimization.
-We still learn feature-based predicates via hill-climbing in this case.
-  • $\bullet$
-    No subselect: An ablation of pix2pred where we perform no hill-climbing, but rather select all predicates in $\Psi_{\text{pool}}$.
-  • $\bullet$
-    No invent: An ablation of pix2pred where we do not propose or subselect any additional predicates beyond the provided ones in $\Psi_{\text{init}}$.
-  • $\bullet$
-    No visual: An ablation of pix2pred where we only invent feature-based predicates.
-  • $\bullet$
-    No feature: An ablation of pix2pred where we only invent visual predicates.
-  • $\bullet$
-    VLM feat.
-pred: A baseline inspired by recent work from Han et al.
-(2024).
-Here, a VLM writes predicates as functions over the feature-based state instead of being able to invent its own novel visual concepts.
-  • $\bullet$
-    ViLa: A baseline that does not learn abstractions but uses a VLM to plan on test tasks directly (Hu et al.
-2023a).
-We test two variants of this: ViLa-pure which is the original approach as implemented by Hu et al.
-(2023a) adapted to our setting, and ViLa-fewshot, which additionally provides the demonstrations as part of the few-shot prompt to give the VLM some context of what kinds of plans it should return for testing tasks.
-Experimental Setup.
-Each seed varies the number and locations of objects in the train and test tasks, and we average results over $5$ random seeds for each task distribution.
-For each seed, we sample 10 test tasks.
-We use GPT-4o (OpenAI 2023) as the VLM for all approaches.
-For our tasks Burger, as mentioned in section 4.3, we early stop the predicate subselection optimization when the hill-climbing objective value goes below $2000$ (a value we found by running pix2pred on a validation set).
-We evaluate each approach by providing the test task and allowing the approach to make and execute exactly one single plan.
-Results and Analysis.
-Figure 7 shows our success rate plots.
-pix2pred outperforms previous approaches on 3 out of 4 of the task distributions by a wide margin.
-In particular, pix2pred seems to generalize to more complex tasks much better than using few-shot learning with a VLM (i.e., ViLa) (Q1).
-In the short-horizon kitchen task that requires 2 actions, we find that ViLa with demonstrations provided performs comparably to pix2pred.
-However, the performance of this baseline drops-off significantly in the longer-horizon burger tasks that require more than 15 actions.
-In particular, we found that ViLa seems to solve testing tasks by pattern-matching aspects of the training demonstrations and is unable to solve test tasks that are unlike any of the demonstrations.
-For instance, the agent never starts out holding an object in any of the training demonstrations, but it does so in about 50% of the testing task instances in the Burger environment.
-We observed that ViLa does not correctly put down the held object before attempting to pick up other objects, whereas pix2pred learns a predicate corresponding to the agent’s hands being empty and correctly selects this as a precondition for picking up any other object.
-We find that pix2pred only fails due to noise in VLM labeling: the VLM infrequently labels the initial state incorrectly such that a critical precondition is false, which prevents our planner from making a valid plan.
-We also find that explicit optimization via hill-climbing leads to much better performance on testing tasks when compared to simply selecting predicates using a VLM (Q2).
-This can be seen from the fact that the VLM subselect baseline performs signifcantly worse than our main approach across all 4 task distributions.
-Qualitatively, we notice that this baseline tends to select far more predicates than necessary: in the burger domain, it selects predicates Cooked(?p1: patty) and Chopped(?l1: lettuce) (necessary for decision-making), but also extraneous predicates like Raw(?p1: patty), Available(?p1: patty), and Whole(?l1: lettuce) that the VLM labels quite inconsistently.
-This results in a complex set of operators that overfit strongly to the training tasks and do not generalize well.
-Additionally, we see that the “no subselect” baseline fails completely at all tasks, indicating that some subselection is critical; learning operators in terms of the more than 500 predicates in the proposal pool leads to extremely overfit operators.
-We find that both feature-based and visual predicates are important in most of our tasks (Q3).
-The “no invent“ baseline fails completely in all the tasks because the initial predicates are insufficient.
-The “no visual.” baseline performs somewhat well in the Kitchen task, but fails in the Burger tasks where visual predicates, specifically predicates corresponding to Cooked and Chopped, are necessary, since the feature-based state does not contain sufficient information to classify these concepts.
-Interestingly, the predicate corresponding to whether the robot’s hand is empty can be classified more accurately in Burger tasks if features of the robot’s hand are used vs.
-attempting to recognize that concept visually.
-This leads to our approach doing better than the “no feat.” baseline in the Burger tasks, since our approach is not only limited to purely visual predicates and can invent whichever predicate has higher classification accuracy.
-6 Limitations and Conclusion
-We proposed pix2pred, a method for inventing symbolic predicates that operate over raw images.
-We found that our approach, from just a handful of demonstrations (3-12), is able to invent semantically meaningful predicates that afford efficient planning and generalization to novel tasks across a range of domains and problems.
-There are several noteworthy limitations of our present work.
-Firstly, VLM hallucinations adversely impact predicate proposal and atom labelling – especially with small amounts of demonstration data – which can inhibit our approach’s ability to learn useful predicates and operators.
-Secondly, the hill-climbing framework we use for predicate invention can be extremely slow, especially as the number of demonstrations and size of the initial predicate pool grows.
-Finally, our approach assumes input demonstrations are segmented in terms of a provided set of parameterized controllers with names that correspond to their function.
-One interesting future direction is to expand the structure of predicates a VLM is allowed to produce, perhaps by allowing it to write arbitrarily-long text descriptions or even code to define new predicates (Liang et al.
-2024).
-Another is to improve the underlying hill-climbing optimization procedure to be significantly more efficient and tolerant to noise, enabling pix2pred to scale to much larger domains that require potentially many tens or hundreds of predicates.
-Eventually we’d like to extend pix2pred to automatically segment demonstrations and learn controllers so that we can learn to plan directly from the simplest low-level demonstration data.
-Appendix A Appendix
-A.1 Planner Implementation Details
-1 Input: Task $\langle\mathcal{O},x_{0},g\rangle$, predicates $\Psi$, operators $\Omega$, samplers $\Sigma$.
-2 $s_{0}\leftarrow\textsc{Abstract}(x_{0},\Psi)$
-3 Call classical planner to generate an abstract plan: $\text{plan}^{\text{abs}}=\underline{\omega_{0}},\underline{\omega_{1}},...,\underline{\omega_{m}}\leftarrow\text{Planner}(s_{0},\mathcal{O},\Psi,\Omega,g)$
+Following previous work [42, 26, 6], we leverage continuous parameter samplers to do this.
+A.2 Planner Implementation Details
+1Input: Task $\langle\mathcal{O},s_{0},g\rangle$, predicates $\Psi$, operators $\Omega$, samplers $\Sigma$.
+2 $s^{\Psi}_{0}\leftarrow\textsc{Abstract}(s_{0},\Psi)$
+3 Call classical planner to generate an abstract plan: $\text{plan}^{\text{abs}}=\underline{\omega_{0}},\underline{\omega_{1}},...,\underline{\omega_{m}}\leftarrow\text{Planner}(s^{\Psi}_{0},\mathcal{O},\Psi,\Omega,g)$
 4 Extract controller sequence with discrete args filled in: $\text{skeleton}=\underline{\mathcal{C}_{0}},\underline{\mathcal{C}_{1}},...,\underline{\mathcal{C}_{m}}$
@@
 6  Sample $\theta\sim\sigma_{\omega_{i}}$, where $\theta$ represents the continuous parameters of $\mathcal{C}_{i}$ and $\sigma_{\omega_{i}}$ is the sampler associated with $\omega_{i}$
-87  Use $\theta$ to fully ground controller $\underline{\mathcal{C}_{i}}$ into an action $a_{i}$. execute action $a_{i}$ and obtain the following state $x_{i+1}\leftarrow f(x_{i},a_{i})$
+7  Use $\theta$ to fully ground controller $\underline{\mathcal{C}_{i}}$ into an action $a_{i}$.
+8 execute action $a_{i}$ and obtain the following state $s_{i+1}\leftarrow f(s_{i},a_{i})$
 Algorithm 2 Planning and Execution pseudocode
-Algorithm 1 shows the pseudocode for the planning and execution strategy we implement following recent work (Kumar et al.
-2024b).
-Given an initial state $x_{0}$ for a task, we simply evaluate the classifiers of the given predicates $\Psi$ to convert the state into an abstract state $s_{0}$.
+Algorithm 2 shows the pseudocode for the planning and execution strategy we implement following recent work [28].
+Given an initial state $x_{0}$ for a task, we simply evaluate the classifiers of the Initial predicates $\Psi$ to convert the state into an abstract state $s_{0}$.
 We then call a classical planner with the task object set $\mathcal{O}$ and operators $\Omega$ to compute an abstract plan that achieves the goal (if one exists).
@@
 We note that more sophisticated planning strategies are possible.
-In particular, provided a photorealistic simulation environment for the transition function $f$, we could leverage many task and motion planning (Garrett et al.
-2021), such as bilevel planning (Silver et al.
-2023; Kumar et al.
-2023; Chitnis et al.
-2022).
-A.2 Operator Learning
-We adapt the “cluster and intersect” operator learning strategy from previous work (Chitnis et al.
-2022) to handle noise in predicate values inherent to our setting.
+In particular, provided a photorealistic simulation environment for the transition function $f$, we could leverage many task and motion planning [11], such as bilevel planning [42, 26, 6].
+A.3 Operator Learning
+We adapt the “cluster and intersect” operator learning strategy from previous work [6] to handle noise in predicate values inherent to our setting.
 Specifically, we learn a set of operators $\Omega$ from our demonstrations $\mathcal{D}$ and predicates $\Psi$ in four steps.
@@
   1. 1.
-    Partitioning: Each demonstration can be expressed as a sequence of transitions $\{(x,a,x^{\prime})\}$, with $x,x^{\prime}\in\mathcal{X}$ and $a\in\mathcal{A}$.
+    Partitioning: Each demonstration can be expressed as a sequence of transitions $\{(s,a,s^{\prime})\}$, with $s,s^{\prime}\in\mathcal{S}$ and $a\in\mathcal{A}$.
 Recall that each action $a$ is a controller with particular discrete and continuous arguments specified; let $\underline{\mathcal{C}}$ denote the corresponding controller with the same discrete object argument values, but continuous parameter values left unspecified.
-First, we use $\Psi$ to Abstract all states $x,x^{\prime}$ in the demonstrations $\mathcal{D}$, creating a dataset of transitions $\{(s,a,s^{\prime})\}$ with $s,s^{\prime}\in\mathcal{S}_{\Psi}$.
-Next, we partition these transitions via the following equivalence relation: $(s_{1},a_{1},s_{1}^{\prime})\equiv(s_{2},a_{2},s_{2}^{\prime})$ if the effects and controllers unify, that is, if there exists a mapping between the objects such that $\underline{\mathcal{C}_{1}}$, $(s_{1}-s_{1}^{\prime})$, and $(s_{1}^{\prime}-s_{1})$ are equivalent to $\underline{\mathcal{C}_{2}}$, $(s_{2}-s_{2}^{\prime})$, and $(s_{2}^{\prime}-s_{2})$ respectively.
-After this step, we have effectively ‘clustered’ all transitions in $\mathcal{D}$ together: we can associate each transition $\{(s,a,s^{\prime})\}$ with a particular equivalence class.
+First, we use $\Psi$ to Abstract all states $s,s^{\prime}$ in the demonstrations $\mathcal{D}$, creating a dataset of transitions $\{(s^{\Psi},a,s^{\prime\Psi})\}$ with $s^{\Psi},s^{\prime\Psi}\in\mathcal{S}^{\Psi}$.
+Next, we partition these transitions via the following equivalence relation: $(s_{1}^{\Psi},a_{1},s_{1}^{\prime\Psi})\equiv(s_{2}^{\Psi},a_{2},s_{2}^{\prime\Psi})$ if the effects and controllers unify, that is, if there exists a mapping between the objects such that $\underline{\mathcal{C}_{1}}$, $(s_{1}^{\Psi}-s_{1}^{\prime\Psi})$, and $(s_{1}^{\prime\Psi}-s_{1}^{\Psi})$ are equivalent to $\underline{\mathcal{C}_{2}}$, $(s_{2}^{\Psi}-s_{2}^{\prime\Psi})$, and $(s_{2}^{\prime\Psi}-s_{2}^{\Psi})$ respectively.
+After this step, we have effectively ‘clustered’ all transitions in $\mathcal{D}$ together: we can associate each transition $\{(s^{\Psi},a,s^{\prime\Psi})\}$ with a particular equivalence class.
   2. 2.
-    Arguments and Effects induction: For each equivalence class created in the previous step, we create $\overline{v}$ by selecting an arbitrary transition $(s,a,s^{\prime})$ and replacing each object that appears in the controller $\underline{\mathcal{C}}$ or effects with a variable of the same type.
+    Arguments and Effects induction: For each equivalence class created in the previous step, we create $\overline{v}$ by selecting an arbitrary transition $(s^{\Psi},a,s^{\prime\Psi})$ and replacing each object that appears in the controller $\underline{\mathcal{C}}$ or effects with a variable of the same type.
 This further induces a substitution $\delta:\overline{v}\to\mathcal{O}$ for the objects $\mathcal{O}$ in this transition.
-Given this, the $E^{+}$, and $E^{-}$ can then be created by applying $\delta$ to $(s^{\prime}-s)$, and $(s-s^{\prime})$ respectively.
+Given this, the $E^{+}$, and $E^{-}$ can then be created by applying $\delta$ to $(s^{\prime\Psi}-s^{\Psi})$, and $(s^{\Psi}-s^{\prime\Psi})$ respectively.
 By construction, for all other transitions $\tau$ in the same equivalence class, there exists an injective substitution $\delta_{\tau}$ under which the controller arguments and effects are equivalent to the newly created $E^{+}$, and $E^{-}$.
@@
     Precondition learning: The only remaining component required to turn each equivalence class into an operator is the operator preconditions.
-For this, we perform an intersection over all abstract states in each equivalence class (Bonet and Geffner 2019; Curtis et al.
-2021).
-Recall that an abstract state is simply the collection of ground atoms that are ‘true’, thus taking an intersection amounts to finding the set of atoms that are always true across every initial state of every transition $(s,a,s^{\prime})$ in the equivalence class.
+For this, we perform an intersection over all abstract states in each equivalence class [4, 8].
+Recall that an abstract state is simply the collection of ground atoms that are ‘true’, thus taking an intersection amounts to finding the set of atoms that are always true across every initial state of every transition $(s^{\Psi},a,s^{\prime\Psi})$ in the equivalence class.
 However, since some of our predicate classifiers might be noisy they might not always hold.
 Thus, we take a ‘soft’ intersection: we take any atom as a precondition that is true across more than a specific percentage (set by a hyperparameter $h_{\text{pre\_frac}}$) of transitions in the equivalence class.
-More specifically, $P\leftarrow\bigcap_{\tau=(s,\cdot,\cdot)}\delta^{-1}_{\tau}(s)$ if $\frac{|\delta^{-1}_{\tau}(s)|}{|\mathcal{D}|}\geq h_{\text{pre\_frac}}$, where $\delta^{-1}_{\tau}(s)$ substitutes all occurrences of the objects in $s$ with the parameters in $\overline{v}$ following an inversion of $\delta_{\tau}$, and discards any atoms involving objects that are not in the image of $\delta_{\tau}$.
-$|\delta^{-1}_{\tau}(s)|$ denotes the number of transitions in $\mathcal{D}$ in which the lifted atom $|\delta^{-1}_{\tau}(s)|$ holds, and $|\mathcal{D}|$ denotes the total number of transitions in $\mathcal{D}$.
+More specifically, $P\leftarrow\bigcap_{\tau=(s^{\Psi},\cdot,\cdot)}\delta^{-1}_{\tau}(s^{\Psi})$ if $\frac{|\delta^{-1}_{\tau}(s^{\Psi})|}{|\mathcal{D}|}\geq h_{\text{pre\_frac}}$, where $\delta^{-1}_{\tau}(s^{\Psi})$ substitutes all occurrences of the objects in $s$ with the parameters in $\overline{v}$ following an inversion of $\delta_{\tau}$, and discards any atoms involving objects that are not in the image of $\delta_{\tau}$.
+$|\delta^{-1}_{\tau}(s^{\Psi})|$ denotes the number of transitions in $\mathcal{D}$ in which the lifted atom $|\delta^{-1}_{\tau}(s^{\Psi})|$ holds, and $|\mathcal{D}|$ denotes the total number of transitions in $\mathcal{D}$.
 In our experiments, we set $h_{\text{pre\_frac}}$ to 0.8.
@@
 In our experiments, we set $h_{\text{data\_frac}}=0.05$.
-A.3 Sampler Learning
-To enable execution, we must also learn samplers for proposing continuous controller parameters (Algorithm 1).
-Adapting prior work (Silver et al.
-2023; Kumar et al.
-2023; Chitnis et al.
-2022), we train one sampler per operator, defined as:
-$\sigma(x,o_{1},\dots,o_{k})=s_{\sigma}(x[o_{1}]\oplus\cdots\oplus x[o_{k}]),$
+A.4 Sampler Learning
+To enable execution, we must also learn samplers for proposing continuous controller parameters (Algorithm 2). Adapting prior work [42, 26, 6], we train one sampler per operator, defined as:
+$\sigma(s,o_{1},\dots,o_{k})=u_{\sigma}(s[o_{1}]\oplus\cdots\oplus s[o_{k}]),$
 where $x[o]$ is the feature vector for $o$, $\oplus$ denotes concatenation, and $s_{\sigma}$ is a learned model.
 Treating this as supervised learning, we use operator-specific datasets $\tau_{\omega}$, where each of these datasets is composed of all transitions from the equivalence class used to induce this operator (Appendix A.3).
-Each transition $(s_{i},a_{i+1},s_{i+1})\in\tau_{\omega}$ maps operator arguments $\overline{v}$ to objects via $\delta:\overline{v}\to\mathcal{O}_{\tau}$.
-Using this mapping, the input for training is $x[\delta(v_{1})]\oplus\cdots\oplus x[\delta(v_{k})]$, where $(v_{1},\dots,v_{k})=\overline{v}$, and the output is the continuous parameter vector $\theta$, which are the continuous parameters used for $a_{i+1}$.
+Each transition $(s_{i}^{\Psi},a_{i+1},s_{i+1}^{\Psi})\in\tau_{\omega}$ maps operator arguments $\overline{v}$ to objects via $\delta:\overline{v}\to\mathcal{O}_{\tau}$.
+Using this mapping, the input for training is $s[\delta(v_{1})]\oplus\cdots\oplus s[\delta(v_{k})]$, where $(v_{1},\dots,v_{k})=\overline{v}$, and the output is the continuous parameter vector $\theta$, which are the continuous parameters used for $a_{i+1}$.
 Each sampler is parameterized by two neural networks.
 The first predicts a Gaussian distribution over $\theta$, regressing its mean and covariance.
-The second network classifies $(x[o_{1}]\oplus\cdots\oplus x[o_{k}],\theta)$ as valid or invalid, enabling rejection sampling from the Gaussian.
+The second network classifies $(s[o_{1}]\oplus\cdots\oplus s[o_{k}],\theta)$ as valid or invalid, enabling rejection sampling from the Gaussian.
 Negative examples are transitions outside $\tau_{\omega}$ but using the same controller.
-A.4 More on Goals
-The agent may sometimes invent visual predicates that can be interpreted as component of the task goal.
-For example, the goal of a task may be to make a burger with a cooked patty inside it.
-Here, the goal predicate would be an indicator for if this is achieved or not, perhaps called BurgerComplete(bottom_bun, patty, top_bun), in which case Cooked(patty) is an implicit component of that goal, in that BurgerComplete’s classifier implicitly checks that the patty is cooked.
-This is in contrast with a situation where the goal of a task is to fill a cup of water, and where the cup starts out upside down so that you can’t pour into it.
-Here, Upright(cup) may not be a component of HasWater(cup), in that the classifier for HasWater may not be explicitly checking if the cup is upright.
-In the former case, one could argue that there is no usefulness to inventing Cooked if we already have the goal predicate BurgerComplete, as we should be able to extract this concept from inside the goal predicate.
-But this may not be possible depending on how the classifier for the goal predicate was implemented – we may not have access to its internals.
-And even if we did, it may not be realistically possible to generate a dataset of burgers with and without cooked patties and label them with BurgerComplete’s classifier to obtain a dataset to train a classifier for Cooked.
-Furthermore, in our experimental setup, we would not necessarily need the classifier for BurgerComplete at all: at training time, any ground atoms associated with this predicate could be manually labeled in the demonstrations (easy to do when there are few goals per task), and at test time, because we execute one single plan without execution monitoring and replanning, we only need to check if the agent has solved the task once when the agent has completed its plan, which we could also do manually (analogously, the head chef at a restaurant does a manual check on the final burgers produced by the cooking staff).
 A.5 Additional Prompting Details
-Here, we provide additional details how we prompt a VLM for both labeling and proposal of atoms.
-Atom Labeling.
-Recall from Section 4.1 that the goal of atom labeling is to obtain the truth value of a particular ground atom $\underline{\psi}(o_{0},\ldots,o_{l})$ given an image-based state at timestep $t$ in a trajectory $x_{t}^{\text{img}}$.
-We do this by prompting a VLM with a text prompt, as well as a string representation of the atom (e.g.
-Cooked(patty1)) and asking it to output the truth value of the atom.
-Since we usually want to query for the values of many atoms at once, we provide all atoms in one single query and ask a VLM to label their truth values simultaneously.
+Here, we provide additional details about how we prompt a VLM for both labeling and proposal of atoms.
+A.5.1 Atom Labeling.
+Recall from Section 4.2 that the goal of atom labeling is to obtain the truth value of a particular ground atom $\underline{\psi}(o_{0},\ldots,o_{l})$ given an image-based state at timestep $t$ in a trajectory $s_{t}^{\text{img}}$.
+We do this by prompting a VLM with a text prompt that includes a string representation of the atom (e.g.
+Cooked(patty1)), along with the relevant images in $s_{t}^{\text{img}}$, and asking it to output the truth value of the atom.
+Since we usually want to query for the values of many atoms at once, we provide all atoms in one single query and ask a VLM to label their truth values in batch.
+Having separate queries to label each ground atom would probably lead to more accurate labels, but would be much slower.
 We use a different prompt depending on whether timestep $t=0$ or $t>0$.
-For $t=0$, we simply provide the initial image-based state $x_{0}^{\text{img}}$ and a prompt (shown below) that asks the model to label all the values of atoms listed in the prompt with either “True’, “False” or “Unknown”.
+For $t=0$, we simply provide the initial image-based state $s_{0}^{\text{img}}$ and a prompt (shown below) that asks the model to label all the values of atoms listed in the prompt with either “True’, “False” or “Unknown”.
 You are a vision system for a robot.
@@
 For each output value, provide an explanation as to why you labelled this predicate as having this particular value.
-Use the format: $<$predicate$>$: $<$truth_value$>$.
-$<$explanation$>$.
+Use the format: <<predicate>>: <<truth_value>>.
+<<explanation>>.
 Predicates:
-We then specify the ground atom strings whose values we’d like to query. For instance, in the “Bigger Burger” task, these might be:
+We then specify the ground atom strings whose values we’d like to query. For instance, in one of the Burger tasks, some of these might be:
 on_table(patty1)
@@
 uncooked(patty2)
-We found two aspects of this were critical to labeling accuracy.
-Firstly, chain-of-thought prompting (Wei et al.
-2022) was extremely useful: our prompt asks the VLM to explicitly provide reasoning for its choices in addition to labeling the truth value.
-Secondly, recall that the image-based state $x^{\text{img}}$ includes segmentation of object names rendered onto the image itself (as depicted in Figure 2).
-This is a form of set-of-marks prompting (Yang et al.
-2023) and is critical to the VLM being able to identify objects being referred to in the text prompt.
-For all states $x_{t}$ in a trajectory that are not the initial state (i.e., $t>0$), we prompt the VLM with the following information: (1) the image-based state from the current and previous timesteps (i.e.
-$x_{t}^{\text{img}}$ and $x_{t-1}^{\text{img}}$), (2) the action executed to get to the current state (i.e.
+Two aspects of our prompting strategy proved essential for achieving high labeling accuracy.
+First, chain-of-thought prompting [49] was extremely useful: our prompt (for $t>0$) asks the VLM to explicitly provide reasoning for its choices in addition to labeling the truth value.
+Prompting the VLM to describe differences between two scenes in a chain-of-thought fashion led to much more accurate output than simply labeling atoms based on a single scene.
+Second, we found that a form of set-of-marks prompting [50] where we annotate objects in the image with their names to be critical to the VLM being able to correctly identify objects being referred to in the text prompt.
+We only perform this annotation in the simulation environments, as we found it to be unnecessary and additional engineering effort in the real-world environments.
+Thus, in addition to assuming we have been given a set of user-specified objects (with names, an optional type, and descriptor string), we also assume, for the simulation environments (in which all objects have a type), that we have access to an object detection system that can classify objects according to their type, and track individual object instances over time, disambiguating between multiple objects of the same type across frames.
+An example of such annotations in the Burger and Coffee environments is shown in Figure 5.
+Notice that we differentiate between ”cup0” and ”cup1”.
+In the simulation environments, an annotation for an object is just the type of the object plus a number.
+Figure 5: Example object annotations in Burger and Coffee.
+For all states $s_{t}$ in a trajectory that are not the initial state (i.e., $t>0$), we prompt the VLM with the following information: (1) the image-based state from the current and previous timesteps (i.e.
+$s_{t}^{\text{img}}$ and $s_{t-1}^{\text{img}}$), (2) the action executed to get to the current state (i.e.
 $a_{t-1})$, and (3) the response of the VLM from the previous timestep.
-We modify the prompt to the VLM to be as follows (here we show some of the predicates queried and a particular action for a certain timestep in a trajectory for a Burger task):
+Here is an example of a prompt in one of the Burger tasks, at particular timestep.
 You are a vision system for a robot.
@@
 For each predicate value, provide an explanation as to why you labeled this predicate as having this particular value.
-Use the format: $<$predicate$>$: $<$truth_value$>$.
-$<$explanation$>$.
-Your response should have three sections. Here is an outline of what your response should look like:
-[START OUTLINE]
-# Expected changes based on the executed skill
-[insert your analysis on the expected changes you will see based on the skill that was executed]
-# Visual changes observed between the images
-[insert your analysis on the visual changes observed between the images]
-# Predicate values in the after image
-[insert your bulleted list of ‘* ¡predicate¿: ¡truth value¿. ¡explanation¿‘]
-[END OUTLINE]
-Predicates: available(robot)
-busy(robot)
-cooked(patty1)
-empty(bottom_bun1)
-empty(grill)
-empty(robot)
-empty_grill(grill)
-free(robot)
-Skill executed between states: Place[robot:robot, top_bun1:top_bun, patty1:patty]
-We also provide the response of the VLM from a previous timestep, an example of which is shown below.
-Predicate values in the first scene, before the skill was executed: # Expected changes based on the executed skill
-The skill executed is ”Pick[robot:robot, top_bun1:top_bun]”. We expect the robot to pick up the top bun. The top bun should no longer be on the ground or table and should be in the robot’s possession.
-# Visual changes observed between the images
-In the before image, the top bun is on the ground. In the after image, the robot is holding the top bun, indicating that the robot has successfully picked it up.
-# Predicate values in the after image
-* available(robot): Unknown. The previous value was unknown, and the pick action does not provide information about availability.
-* busy(robot): Unknown. The previous value was unknown, and the pick action does not provide information about busyness.
-* cooked(patty1): True. The patty remains cooked, as it was cooked before the action. Previously, the value was true.
-* empty(bottom_bun1): False. The bottom bun remains unchanged. Previously, the value was false.
-* empty(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
-* empty(robot): False. The robot is now holding the top bun. Previously, the value was true.
-* empty_grill(grill): True. The grill remains empty, as the action does not involve the grill. Previously, the value was true.
-* free(robot): Unknown. The previous value was unknown, and the pick action does not provide information about freedom. …
-We found that prompting the VLM to describe differences between two scenes in a chain-of-thought fashion led to much more accurate output than simply labeling atoms based on a single scene.
+Use the format: <<predicate>>: <<truth_value>>.
+<<explanation>>.
+Your response should have three sections.
+Here is an outline of what your response should look like: [START OUTLINE] # Expected changes based on the executed skill [insert your analysis on the expected changes you will see based on the skill that was executed] # Visual changes observed between the images [insert your analysis on the visual changes observed between the images] # Predicate values in the after image [insert your bulleted list of ‘* ¡predicate¿: ¡truth value¿.
+¡explanation¿‘] [END OUTLINE] Predicates: available(robot) busy(robot) cooked(patty1) empty(bottom_bun1) empty(grill) empty(robot) empty_grill(grill) free(robot) Skill executed between states: Place[robot:robot, top_bun1:top_bun, patty1:patty] Predicate values in the first scene, before the skill was executed: # Expected changes based on the executed skill The skill executed is ”Pick[robot:robot, top_bun1:top_bun]”.
+We expect the robot to pick up the top bun.
+The top bun should no longer be on the ground or table and should be in the robot’s possession.
+# Visual changes observed between the images In the before image, the top bun is on the ground.
+In the after image, the robot is holding the top bun, indicating that the robot has successfully picked it up.
+# Predicate values in the after image * available(robot): Unknown.
+The previous value was unknown, and the pick action does not provide information about availability.
+* busy(robot): Unknown.
+The previous value was unknown, and the pick action does not provide information about busyness.
+* cooked(patty1): True.
+The patty remains cooked, as it was cooked before the action.
+Previously, the value was true.
+* empty(bottom_bun1): False.
+The bottom bun remains unchanged.
+Previously, the value was false.
+* empty(grill): True.
+The grill remains empty, as the action does not involve the grill.
+Previously, the value was true.
+* empty(robot): False.
+The robot is now holding the top bun.
+Previously, the value was true.
+* empty_grill(grill): True.
+The grill remains empty, as the action does not involve the grill.
+Previously, the value was true.
+* free(robot): Unknown.
+The previous value was unknown, and the pick action does not provide information about freedom.
+…
 There are two additional prompting techniques we found helpful and important for labeling accuracy: (1) asking the VLM to “double-check” its label output based on the explanations it provided, and (2) augmenting the image-based state to include a close-up crop of the image for the objects the robot is currently interacting with.
-For this double-checking, we provide the VLM with the following prompt:
+We perform (2) only in the Burger domain.
+For double-checking, we provide the VLM with the following prompt:
 Sometimes your reasoning about the value of a predicate at the current timestep uses an incorrect value of that predicate in the previous timestep.
@@
 Regardless of whether or not it needs correctly, your reply should be formatted exactly the same as the previous answer.
-We use double-checking only at training time to maximize labeling accuracy before hill-climbing. We augment images with additional object crops only for tasks in the Burger domain.
-Figure 9: Labeling of ground atoms with visual predicates involving the lettuce object type in Burger.
-This figure depicts the visual state before and after the chop skill in demonstration #1 for the task distribution “Combo Burger”.
-It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the lettuce object in these states.
-The VLM proposes a wide variety of predicates involving the patty: assemble, chopped, clear, combine, cut, intact, sliced, uncut, and whole.
-Of these, chopped(?l:lettuce) and cut(?l:lettuce) are the most relevant to achieving the goal while also being accurately labeled.
-Our approach automatically selects the former predicate from the pool.
-Atom Proposal
+We use double-checking only at training time to maximize labeling accuracy before optimization.
+A.5.2 Atom Proposal
 Recall from Section 4.2 that the objective of atom proposal is to generate an initial pool of visual predicates given a set of demonstrations $\mathcal{D}$.
@@
 For each predicate you generate, also generate some predicates that are synonyms and antonyms so that any predicate that is even tangentially relevant to the demonstrations is generated.
-Once again, we found that chain-of-thought prompting was critical to good proposals.
-We also found that ensuring all the image-based states have segmented objects with names rendered on them was important to get a good set of proposals.
 We then provide a text representation of the action sequence executed in the demonstration trajectory. An example in a Burger task is:
@@
 An example response from the VLM in a Burger task is shown below.
-### Predicates for Each Action
-1. **Pick[robot:robot, patty1:patty]**
-- Before: on(patty1, ground)
-- After: holding(robot, patty1)
-- Synonyms: grasp(robot, patty1)
-- Antonyms: release(robot, patty1)
-2. **Place[robot:robot, patty1:patty, grill:grill]**
-- Before: holding(robot, patty1)
-- After: on(patty1, grill)
-- Synonyms: position(patty1, grill)
-- Antonyms: off(patty1, grill)
-3. **Cook[robot:robot, patty1:patty, grill:grill]**
-- Before: raw(patty1)
-- After: cooked(patty1)
-- Synonyms: grill(patty1)
-- Antonyms: uncooked(patty1)
-4. **Pick[robot:robot, patty1:patty]**
-- Before: on(patty1, grill)
-- After: holding(robot, patty1)
-- Synonyms: grasp(robot, patty1)
-- Antonyms: release(robot, patty1)
-5. **Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun]**
-- Before: holding(robot, patty1)
-- After: on(patty1, bottom_bun1)
-- Synonyms: position(patty1, bottom_bun1)
-- Antonyms: off(patty1, bottom_bun1)
-6. **Pick[robot:robot, top_bun1:top_bun]**
-- Before: on(top_bun1, ground)
-- After: holding(robot, top_bun1)
-- Synonyms: grasp(robot, top_bun1)
-- Antonyms: release(robot, top_bun1)
-7. **Place[robot:robot, top_bun1:top_bun, patty1:patty]**
-- Before: holding(robot, top_bun1)
-- After: on(top_bun1, patty1)
-- Synonyms: position(top_bun1, patty1)
-- Antonyms: off(top_bun1, patty1)
-### Other Important Predicates
-- **Initial State:**
-- on(patty1, ground)
-- on(bottom_bun1, ground)
-- on(top_bun1, ground)
-- empty(robot)
-- **Final State:**
-- on(top_bun1, patty1)
-- on(patty1, bottom_bun1)
-- empty(robot)
-- **Synonyms:**
-- empty(robot), free(robot)
-- on(obj1, obj2). above(obj1, obj2)
-- **Antonyms:** - empty(robot), holding(robot, obj)
-- on(obj1, obj2), off(obj1, obj2)
+Predicates for Each Action 1.
+Pick[robot:robot, patty1:patty] Before: on(patty1, cutting_board) After: holding(robot, patty1) Synonyms: grasping(robot, patty1), carrying(robot, patty1) Antonyms: releasing(robot, patty1), dropping(robot, patty1) 2.
+Place[robot:robot, patty1:patty, grill:grill] Before: holding(robot, patty1) After: on(patty1, grill) Synonyms: atop(patty1, grill), above(patty1, grill) Antonyms: off(patty1, grill), under(patty1, grill) 3.
+Cook[robot:robot, patty1:patty, grill:grill] Before: raw(patty1) After: cooked(patty1) Synonyms: grilled(patty1), prepared(patty1) Antonyms: raw(patty1), uncooked(patty1) 4.
+Pick[robot:robot, patty1:patty] Before: on(patty1, grill) After: holding(robot, patty1) Synonyms: grasping(robot, patty1), carrying(robot, patty1) Antonyms: releasing(robot, patty1), dropping(robot, patty1) 5.
+Place[robot:robot, patty1:patty, bottom_bun1:bottom_bun] Before: holding(robot, patty1) After: on(patty1, bottom_bun1) Synonyms: atop(patty1, bottom_bun1), above(patty1, bottom_bun1) Antonyms: off(patty1, bottom_bun1), under(patty1, bottom_bun1) 6.
+Pick[robot:robot, top_bun1:top_bun] Before: on(top_bun1, cutting_board) After: holding(robot, top_bun1) Synonyms: grasping(robot, top_bun1), carrying(robot, top_bun1) Antonyms: releasing(robot, top_bun1), dropping(robot, top_bun1) 7.
+Place[robot:robot, top_bun1:top_bun, patty1:patty] Before: holding(robot, top_bun1) After: on(top_bun1, patty1) Synonyms: atop(top_bun1, patty1), above(top_bun1, patty1) Antonyms: off(top_bun1, patty1), under(top_bun1, patty1) Other Important Predicates Additional Initial State Predicates: on(bottom_bun1, cutting_board) on(cheese1, cutting_board) on(lettuce1, cutting_board) near(robot, grill) near(robot, bottom_bun1) near(robot, top_bun1) Synonyms: on(x, y): atop(x, y), above(x, y) holding(x, y): grasping(x, y), carrying(x, y) cooked(x): grilled(x), prepared(x) raw(x): uncooked(x), fresh(x) near(x, y): close_to(x, y), adjacent_to(x, y) Antonyms: on(x, y): off(x, y), under(x, y) holding(x, y): releasing(x, y), dropping(x, y) cooked(x): raw(x), uncooked(x) raw(x): cooked(x), prepared(x) near(x, y): far_from(x, y), distant_from(x, y)
+We do the same set-of-marks prompting that we do for atom labeling, for atom proposal.
+Also, note that in the Coffee environment, we add ”Generate only 5-10 predicates” to the proposal prompt to help restrict the total number of predicates the VLM proposes.
+This alleviates some of the difficulty the hill-climbing procedure would normally have to deal with, when faced with a very large set of predicates (e.g.
+200+).
+Note that the VLM proposes atoms for each trajectory, so when we ask it to generate 5-10 atoms, it generates between $5n$ and $10n$ atoms in total if we had n demonstrations.
 A.6 Additional Experimental Details
-Additional Environment Details.
+A.6.1 Experimental Evaluation and Hyperparameters.
+We used the following hyperparameter values in our different domains. Where applicable, these values were determined via a validation set. Here $N_{\text{demo}}$ is the size of the training dataset.
+Task $J_{\text{thresh}}$ $h_{\text{pre\_frac}}$ $h_{\text{data\_frac}}$ $N_{\text{demo}}$
+Kitchen 100 0.8 0.05 3
+Bigger Burger 2000 0.8 0.05 12
+More Burger Stacks 2000 0.8 0.05 12
+Combo Burger 2000 0.8 0.05 12
+Coffee 100 0.8 0.05 5
+Cleanup 100 1.0 0.4 6
+Juice 0 1.0 0.4 10
+Table 2: Domain parameters used in experiments.
+A.6.2 Additional Environment Details.
 Here, we describe in detail the initial predicates, training demonstrations, and test tasks for each task in each of our experimental environments.
 Note that the controllers listed below have discrete object parameters (indicated by the ‘?’) as well as continuous parameters where applicable $\theta$ (shown within []).
+Figure 6: Labeling of ground atoms with visual predicates involving the patty object type in Burger.
+This figure depicts the visual state before and after the cook skill in demonstration #0 for the task distribution “Combo Burger”.
+It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the patty object in these states.
+The VLM proposes a wide variety of predicates involving the patty: clear, cooked, cool, grilling, raw, and uncook.
+Of these, cooked(?p:patty) is most relevant to achieving the goal while also being accurately labeled.
+Our approach automatically selects this predicate from the pool.
+See Figure 7 in the appendix for an analogous situation with the chop skill.
+Figure 7: Labeling of ground atoms with visual predicates involving the lettuce object type in Burger.
+This figure depicts the visual state before and after the chop skill in demonstration #1 for the task distribution “Combo Burger”.
+It also shows the labels assigned by the VLM to ground atoms with visual predicates involving the lettuce object in these states.
+The VLM proposes a wide variety of predicates involving the patty: assemble, chopped, clear, combine, cut, intact, sliced, uncut, and whole.
+Of these, chopped(?l:lettuce) and cut(?l:lettuce) are the most relevant to achieving the goal while also being accurately labeled.
+Our approach automatically selects the former predicate from the pool.
   • $\bullet$
@@
     • $\bullet$
-      Given Predicates:
+      Initial predicates:
       • $\bullet$
-        KettleBoiling(?k: kettle, ?kn: knob, ?b burner): is true only if ?kn is turned on, and the kettle ?k is ontop of ?b, and if ?kn and ?b are ’linked’ (i.e., ?kn is the knob that causes burner ?b to glow red with heat).
+        KettleBoiling(?k:kettle, ?kn:knob, ?b:burner): is true only if ?kn is turned on, and the kettle ?k is ontop of ?b, and if ?kn and ?b are ’linked’ (i.e., ?kn is the knob that causes burner ?b to glow red with heat).
     • $\bullet$
@@
     • $\bullet$
-      Cook(?r: robot, ?p: patty, ?g: grill). Given the robot is adjacent to grill ?g and patty ?p is atop the grill, cooks the patty so it appears grilled.
+      Cook(?r:robot, ?p:patty, ?g:grill, []). Given the robot is adjacent to grill ?g and patty ?p is atop the grill, cooks the patty so it appears grilled.
     • $\bullet$
-      Chop(?r: robot, ?l: lettuce, ?c: cutting_board).
+      Chop(?r:robot, ?l:lettuce, ?c:cutting_board, []).
 Given the robot is adjacent to cutting board ?c and lettuce ?l is atop the cutting board, chops the lettuce with a knife that’s on the cutting board so it appears chopped.
@@
       • $\bullet$
-        Given predicates:
+        Initial predicates:
         • $\bullet$
@@
       • $\bullet$
-        Given predicates:
+        Initial predicates:
         • $\bullet$
@@
       • $\bullet$
-        Given predicates:
+        Initial predicates:
         • $\bullet$
@@
         • $\bullet$
-          SomewhereAboveAndPrepped(?l: lettuce, ?b: bottom_bun): turns true when ?l is somewhere above ?b and ?l is choppeed.
+          SomewhereAboveAndPrepped(?l:lettuce, ?b:bottom_bun): turns true when ?l is somewhere above ?b and ?l is chopped.
         • $\bullet$
@@
 In the other 5 test tasks, we ask the agent to make two burgers, each with chopped lettuce on a cooked patty, in addition to making another burger with a single cooked patty, and the agent starts out holding a raw patty.
-A.7 Learned Predicate and Operator Examples
-Figures 10, 11, 12, and show our learned predicates and operators in our “Kitchen”, “Bigger Burger”, “Burger More Stacks” and “Combo Burger” tasks respectively.
-Figure 10: Learned predicates and operators for the “Kitchen” task.
-Figure 11: Learned predicates and operators for the “Bigger Burger” task distribution.
-Figure 12: Learned predicates and operators for the “More Burger Stacks” task distribution.
-Figure 13: Learned predicates and operators for the “Combo Burger” task distribution.
+  • $\bullet$
+    Coffee:
+    • $\bullet$
+      Initial predicates:
+      • $\bullet$
+        CupFilled(?c:cup): turns true when cup is filled with liquid above a certain threshold.
+      • $\bullet$
+        Holding(?r:robot, ?j:jug): turns true when robot is holding the jug.
+      • $\bullet$
+        OnTable(?j:jug): turns true when jug is on the table.
+      • $\bullet$
+        HandEmpty(?r:robot): turns true when the robot is not holding anything.
+      • $\bullet$
+        RobotAboveCup(?r:robot, ?c:cup): turns true when the robot is close enough to the cup to pour liquid into the cup from a jug.
+      • $\bullet$
+        JugAboveCup(?j:jug, ?c:cup): turns true when the jug is close enough to the cup to pour liquid into the cup.
+      • $\bullet$
+        NotAboveCup(?r:robot, ?j:jug): turns true when the robot is not close enough to the cup to pour liquid into the cup from a jug.
+      • $\bullet$
+        PressingButton(?r:robot, ?m:coffee_machine): turns true when the robot is pressing the power button of the coffee machine.
+      • $\bullet$
+        NotSameCup(?c1:cup, ?c2:cup): turns true when the two cups in question are not the same object.
+      • $\bullet$
+        JugPickable(?j:jug): turns true when the handle of the jug is accessible so that the robot can pick it up, usually after the skill RotateItemUntilHandleAccessible is run if not true already.
+    • $\bullet$
+      Skills:
+      • $\bullet$
+        PickJug(?r:robot, ?j:jug, []): moves the robot gripper behind the jug’s handle in the y-direction, moves down in the z-direction, and then moves forward in the y-direction before finally grasping the jug by the handle.
+      • $\bullet$
+        PlaceJugInMachine(?r:robot, ?j:jug, ?m:coffee_machine, []): picks up the jug slightly so as to avoid friction with the table, moves the jug to a position under the coffee machine’s dispenser, and then places the jug.
+      • $\bullet$
+        TurnMachineOnAndFill(?r:robot, ?m:coffee_machine, []): moves the robot up to be level with the power button of the coffee machine in the z-direction and then moves forward in the y-direction to press it, and then moves backwards.
+Terminates when the machine has finished dispensing coffee.
+      • $\bullet$
+        PourSomeLiquid(?r:robot, ?j:jug, ?c:cup, []): moves the robot next to the cup and pours from the jug until the cup is filled.
+      • $\bullet$
+        RotateItemUntilHandleAccessible(?r:robot, ?j:jug, []): moves the robot to the jug, rotates the jug until its handle is as close to the robot as possible, and then breaks contact with the jug.
+    • $\bullet$
+      Training demonstrations: we provide 4 demonstrations where a single cup is on the table and the robot fills this single cup with coffee, and 1 demonstration where two cups are on the table and the robot fills both of these cups with coffee.
+    • $\bullet$
+      Test tasks: We task the robot with filling three cups on the table with coffee.
+  • $\bullet$
+    Cleanup:
+    • $\bullet$
+      Initial predicates:
+      • $\bullet$
+        Holding(?r:rrobot, ?m:movable): turns true when a robot ?r is holding item ?m.
+We provide this predicate directly because it is not possible to infer it from the image-based state due to occlusion.
+      • $\bullet$
+        HandEmpty(?r:robot): turns true when the robot isn’t holding any objects.
+We provide this predicate directly because it is not possible to infer it from the image-based state (since no onboard camera can consistently see the hand).
+      • $\bullet$
+        InsideContainer(?m:movable, ?t:trash_can): true when the object ?m is inside the trash can ?t.
+      • $\bullet$
+        WipedOfMarkerScribbles(?tab:table): true if the table ?tab has no obvious markings atop its surface.
+    • $\bullet$
+      Skills:
+      • $\bullet$
+        PickFromTop(?g:robot, ?o:movable, ?s:immovable, []): Moves the robot ?g to grasp the top of movable object ?o, which is supported by immovable surface ?s, and lifts it.
+      • $\bullet$
+        PlaceInside(?g:robot, ?o:movable, ?s:immovable, []): Moves the robot ?g holding object ?o to position it inside immovable container ?s and releases it.
+      • $\bullet$
+        PickFromFloor(?g:robot, ?o:movable, []): Moves the robot ?g to grasp movable object ?o directly from the floor and lifts it.
+      • $\bullet$
+        WipeAndContinueHoldingEraser(?g:robot, ?o:movable, ?t:table, []): Uses the robot ?g holding movable object ?o (e.g., an eraser) to wipe the surface of table ?t while continuing to hold it.
+      • $\bullet$
+        DumpContentsOntoFloor(?g:robot, ?tc:trash_can, []): Moves the robot ?g to tilt trash can ?tc and dump its contents onto the floor.
+    • $\bullet$
+      Training demonstrations: We collect a dataset of 6 demonstrations.
+In 4 of these, the demonstrator clears an object from a table surface, then picks up an eraser and uses it to wipe a particular table of markings.
+In the remaining 2, the demonstrator dumps an object out of a narrow bin in order to pick it up and put it inside another bin.
+    • $\bullet$
+      Testing tasks: As noted in Section 5, we test 5 different task variants.
+For the ‘New object instances’ variant, we use a different table, with a different object on its surface and task the robot with achieving WipeAndContinueHoldingEraser for the new table.
+For the ‘New visual background’ variant, we use a different room.
+For ‘More objects’, we task the robot with placing multiple objects into a particular bin.
+For ‘Novel goal 1’, we setup an initial state where an object is on the floor (not atop the table), and task the robot with placing that object, as well as the eraser use for wiping, in the bin.
+For ‘Novel goal 2’, we setup an initial state where the eraser itself is inside a narrow bin: the robot must first dump the bin to obtain the eraser (never seen at training time), and then use it to wipe the surface before then putting the eraser back in the bin.
+Please see our supplementary video for a recording of the robot performing this task.
+  • $\bullet$
+    Juice:
+    • $\bullet$
+      Initial predicates:
+      • $\bullet$
+        Holding(?r:robot, ?m:movable): turns true when a robot ?r is holding item ?m.
+We provide this predicate directly because it is not possible to infer it from the image-based state due to occlusion.
+      • $\bullet$
+        HandEmpty(?r:robot): turns true when the robot isn’t holding any objects.
+We provide this predicate directly because it is not possible to infer it from the image-based state (since no onboard camera can consistently see the hand).
+      • $\bullet$
+        JuiceInCup(?j:movable, ?c:container): true when the juice object ?j is detected to be inside the container ?c, as determined by the VLM.
+      • $\bullet$
+        Empty(?c:container): true when the container ?c is detected to be empty, as determined by the vision-language model.
+      • $\bullet$
+        Inside(?m:movable, ?j:juicer): true when the movable object ?m is detected to be inside the juicer ?j, as determined by the VLM.
+    • $\bullet$
+      Skills:
+      • $\bullet$
+        PickContainer(?g:robot, ?c:container, []): Moves the robot ?g to grasp and pick up the container ?c.
+      • $\bullet$
+        PickMovable(?g:robot, ?m:movable, []): Moves the robot ?g to grasp and pick up the movable object ?m.
+      • $\bullet$
+        PlaceOnLeft(?g:robot, ?c:container, ?j:juicer, []): Moves the robot ?g holding container ?c to place it inside the waste valve region of juicer ?j.
+      • $\bullet$
+        PlaceOnRight(?g:robot, ?c:container, ?j:juicer, []): Moves the robot ?g holding container ?c to place it inside the juice valve region of juicer ?j.
+      • $\bullet$
+        PlaceInside(?g:robot, ?m:movable, ?j:juicer, []): Moves the robot ?g holding movable object ?m to place it inside juicer ?j.
+      • $\bullet$
+        DumpFromOneIntoOther(?g:robot, ?src:container, ?dest:container, []): Moves the robot ?g to dump the contents of source container ?src into destination container ?dest.
+      • $\bullet$
+        CloseLid(?g:robot, ?j:juicer, []): Moves the robot ?g to close the lid of juicer ?j.
+      • $\bullet$
+        RunMachine(?g:robot, ?j:juicer, ?c1:container, ?c2:container, []): Moves the robot ?g to turn on and run the juicer ?j, using containers ?c1 and ?c2 to collect outputs.
+    • $\bullet$
+      Training demonstrations: We collect a dataset of 10 demonstrations.
+Several demonstrations showcase putting a fruit (either an orange or an apple) into the juicer while cups are present in both the juice and waste valve regions.
+A few demonstrations show one cup absent, and that being placed under a corresponding valve before the juicer is run.
+In some of the demonstrations, one of the cups is full, and the demonstrator dumps the cup into an empty bowl or other container to empty it.
+A few demonstrations show dumping separately, and a few others simply show placing a fruit inside the juicer.
+    • $\bullet$
+      Testing tasks: As noted in Section 5, we test 5 different task variants.
+For the ‘New object instances’ variant, we use a different table, with a different fruit to be juiced.
+For the ‘New visual background’ variant, we use a different room.
+For ‘More objects’, we task the robot with placing multiple objects into the juicer.
+For ‘Novel goal 1’, we setup an initial state where there is no cup under the waste valve of the juicer, and no fruit inside the juicer.
+To solve this task, the robot must place a cup in the waste valve region, and then put the fruit in the juicer and run it.
+For ‘Novel goal 2’, we setup an initial state where there are two cups in the correct regions (waste and juice valve), but one of them is full.
+Thus, the robot must first dump contents from that cup (never seen at training time) and place it back before running the juicer to make juice.
+A.7 Real-World Experiment Details
+A.7.1 Real-Robot System Implementation
+We build on the planning and perception systems implemented by Kumar et al.
+[28].
+In particular, we reuse their perception pipeline (that leverages a combination of DETIC [53] and SAM [24]) to construct our real-world object-centric state $S^{\text{obj}}$.
+We also reuse their localization and mapping pipeline in order to implement movement, which is used implicitly in several of the skills we define for our real-world environments.
+This pipeline requires the environment be outfitted with April tags (https://april.eecs.umich.edu/software/apriltag), as can be seen in our supplementary videos (the tags are thus only used for the movement skills, and nothing else).
+Finally, we reuse several of their samplers (e.g.
+for navigation, and for grasping).
+Similar to Kumar et al. [28], all skills are implemented programmatically via low-level calls to the Boston Dynamics Spot SDK (https://dev.bostondynamics.com/).
+We construct the image-based state $S^{\text{img}}$ by directly using the $6$ on-board RGB cameras on the Boston Dynamics Spot [^1] We slightly postprocess images by rotating them to offset for any camera rotation as described here: https://dev.bostondynamics.com/python/examples/get_image/readme.
+For more details, please see our open-source code release (which will be made public upon acceptance of this paper).
+A.7.2 Robot Domains Additional Details.
+We collected training data for both our real-world domains (Cleanup and Juice) by taking images from the point-of-view of a human performing the task (Figure 1).
+We include an object ‘hand:robot’ in each demonstration to indicate that the demonstrator’s hand is performing actions.
+However, a key challenge with the Boston Dynamics Spot robot is that none of the onboard cameras can see the hand itself.
+This means that any invented visual predicates that have to do with the hand will not transfer to the robot.
+Thus, during initial pool proposal (Section 4.1), we explicitly exclude any visual atoms proposed that mention the hand.
+The only predicates that involve the hand are thus HandEmpty, and Holding since these are provided as part of the initial pool.
+In both Cleanup and Juice, we found that the VLM proposed a very large number (sometimes over $1000$) atoms, leading to an extremely slow hill-climbing optimization to subselect predicates.
+We thus modified the atom proposal prompt to have the VLM restrict itself to proposing no more than $25$ atoms per demonstration.
+In the Juice domain, we found that the VLM struggled with accurately recognizing whether cups were below specific valves regardless of how the predicate was named (e.g.
+‘UnderWasteValve’, ‘ReadyToReceiveWaste’, ‘WithinWasteRegion’).
+We thus manually demarcated and labelled (via directly marking the table surface) the waste and juice regions next to the juicer.
+We found this was necessary for learning a reasonable world model, and for accurately planning and labelling states at test time.
+A.8 Examples of Learned Predicates and Learned Action Models per Environment
+Note that predicates that end in a number are a result of the VLM proposing multiple ground atoms with the same name but with different arguments.
+For example, the VLM might propose two ground atoms, ”Clear(jug0)” and ”Clear(table0)”.
+Because each predicate must have a unique set of typed object arguments, we rename these ground atoms to ”Clear0(jug0)” and ”Clear1(table0)” during the process of parsing predicates from the VLM’s response, and would end up with two predicates, ”Clear0” and ”Clear1”.
+Further note that if you see numbers skipped in the names of the action models, that is because low-data operators were pruned, as described in Appendix A.3.
+Example learned predicates and learned action models in Kitchen Learned predicates: NOT-[[0:surface].z<=[idx_0]1.59] Learned operators: STRIPS-Op0: Parameters: [?x0:surface, ?x1:gripper, ?x2:knob] Preconditions: [KnobAndBurnerLinked(?x2:knob, ?x0:surface)] Add Effects: [NOT-[[0:surface].z<=[idx_0]1.59](?x0:surface)] Delete Effects: [] Ignore Effects: [] Skill: MoveAndTurnOnKnob(?x1:gripper, ?x2:knob) STRIPS-Op1: Parameters: [?x0:surface, ?x1:gripper, ?x2:kettle, ?x3:knob] Preconditions: [KnobAndBurnerLinked(?x3:knob, ?x0:surface), NOT-[[0:surface].z<=[idx_0]1.59](?x0:surface)] Add Effects: [KettleBoiling(?x2:kettle, ?x0:surface, ?x3:knob)] Delete Effects: [] Ignore Effects: [] Skill: PushKettleOntoBurner(?x1:gripper, ?x2:kettle, ?x0:surface)
+Example learned predicates and learned action models in More Stacks Burger Learned Predicates: cooked0(?p:patty) empty_hands0(?r:robot) Learned operators: STRIPS-Op0: Parameters: [?x0:patty, ?x1:robot] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), OnGround(?x0:patty), empty_hands0(?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:patty)] Delete Effects: [Clear(?x0:patty), OnGround(?x0:patty), empty_hands0(?x1:robot)] Ignore Effects: [] Skill: Pick(?x1:robot, ?x0:patty) STRIPS-Op1: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:grill), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty)] Add Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:grill), empty_hands0(?x2:robot)] Delete Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op2: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), On(?x1:patty, ?x0:grill), empty_hands0(?x2:robot)] Add Effects: [cooked0(?x1:patty)] Delete Effects: [] Ignore Effects: [] Skill: Cook(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op3: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), On(?x1:patty, ?x0:grill), cooked0(?x1:patty), empty_hands0(?x2:robot)] Add Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Delete Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:grill), empty_hands0(?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:patty) STRIPS-Op4: Parameters: [?x0:bottom_bun, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:bottom_bun), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty), OnGround(?x0:bottom_bun), cooked0(?x1:patty)] Add Effects: [Clear(?x1:patty), SomewhereAboveAndPrepped(?x0:bottom_bun, ?x1:patty), On(?x1:patty, ?x0:bottom_bun), empty_hands0(?x2:robot)] Delete Effects: [Clear(?x0:bottom_bun), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:bottom_bun) STRIPS-Op5: Parameters: [?x0:robot, ?x1:top_bun] Preconditions: [Clear(?x0:robot), Clear(?x1:top_bun), OnGround(?x1:top_bun), empty_hands0(?x0:robot)] Add Effects: [Holding(?x0:robot, ?x1:top_bun)] Delete Effects: [Clear(?x1:top_bun), OnGround(?x1:top_bun), empty_hands0(?x0:robot)] Ignore Effects: [] Skill: Pick(?x0:robot, ?x1:top_bun) STRIPS-Op6: Parameters: [?x0:patty, ?x1:robot, ?x2:top_bun] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), Holding(?x1:robot, ?x2:top_bun), cooked0(?x0:patty)] Add Effects: [Clear(?x2:top_bun), On(?x2:top_bun, ?x0:patty), empty_hands0(?x1:robot)] Delete Effects: [Clear(?x0:patty), Holding(?x1:robot, ?x2:top_bun)] Ignore Effects: [] Skill: Place(?x1:robot, ?x2:top_bun, ?x0:patty)
+Example learned predicates and learned action models in Bigger Burger Learned Predicates: cooked0(?p:patty) [[0:robot].fingers<=[idx_0]0.5] Learned operators: STRIPS-Op0: Parameters: [?x0:patty, ?x1:robot] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), OnGround(?x0:patty), [[0:robot].fingers<=[idx_0]0.5](?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:patty)] Delete Effects: [Clear(?x0:patty), OnGround(?x0:patty), [[0:robot].fingers<=[idx_0]0.5](?x1:robot)] Ignore Effects: [] Skill: Pick(?x1:robot, ?x0:patty) STRIPS-Op1: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:grill), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty)] Add Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:grill), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Delete Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op2: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), On(?x1:patty, ?x0:grill), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Add Effects: [RightAboveAndPrepped(?x0:grill, ?x1:patty), cooked0(?x1:patty)] Delete Effects: [] Ignore Effects: [] Skill: Cook(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op3: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), RightAboveAndPrepped(?x0:grill, ?x1:patty), On(?x1:patty, ?x0:grill), [[0:robot].fingers<=[idx_0]0.5](?x2:robot), cooked0(?x1:patty)] Add Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Delete Effects: [Clear(?x1:patty), RightAboveAndPrepped(?x0:grill, ?x1:patty), On(?x1:patty, ?x0:grill), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:patty) STRIPS-Op4: Parameters: [?x0:bottom_bun, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:bottom_bun), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty), OnGround(?x0:bottom_bun), cooked0(?x1:patty)] Add Effects: [Clear(?x1:patty), SomewhereAboveAndPrepped(?x0:bottom_bun, ?x1:patty), On(?x1:patty, ?x0:bottom_bun), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Delete Effects: [Clear(?x0:bottom_bun), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:bottom_bun) STRIPS-Op5: Parameters: [?x0:robot, ?x1:top_bun] Preconditions: [Clear(?x0:robot), Clear(?x1:top_bun), OnGround(?x1:top_bun), [[0:robot].fingers<=[idx_0]0.5](?x0:robot)] Add Effects: [Holding(?x0:robot, ?x1:top_bun)] Delete Effects: [Clear(?x1:top_bun), OnGround(?x1:top_bun), [[0:robot].fingers<=[idx_0]0.5](?x0:robot)] Ignore Effects: [] Skill: Pick(?x0:robot, ?x1:top_bun) STRIPS-Op6: Parameters: [?x0:patty, ?x1:robot, ?x2:top_bun] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), Holding(?x1:robot, ?x2:top_bun), cooked0(?x0:patty)] Add Effects: [Clear(?x2:top_bun), On(?x2:top_bun, ?x0:patty), [[0:robot].fingers<=[idx_0]0.5](?x1:robot)] Delete Effects: [Clear(?x0:patty), Holding(?x1:robot, ?x2:top_bun)] Ignore Effects: [] Skill: Place(?x1:robot, ?x2:top_bun, ?x0:patty) STRIPS-Op7: Parameters: [?x0:cutting_board, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:cutting_board), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty), cooked0(?x1:patty)] Add Effects: [Clear(?x1:patty), RightAboveAndPrepped(?x0:cutting_board, ?x1:patty), On(?x1:patty, ?x0:cutting_board), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Delete Effects: [Clear(?x0:cutting_board), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:cutting_board) STRIPS-Op8: Parameters: [?x0:patty, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:patty), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty), cooked0(?x0:patty), cooked0(?x1:patty)] Add Effects: [Clear(?x1:patty), RightAboveAndPrepped(?x0:patty, ?x1:patty), On(?x1:patty, ?x0:patty), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Delete Effects: [Clear(?x0:patty), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:patty) STRIPS-Op9: Parameters: [?x0:cutting_board, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), RightAboveAndPrepped(?x0:cutting_board, ?x1:patty), On(?x1:patty, ?x0:cutting_board), [[0:robot].fingers<=[idx_0]0.5](?x2:robot), cooked0(?x1:patty)] Add Effects: [Clear(?x0:cutting_board), Holding(?x2:robot, ?x1:patty)] Delete Effects: [Clear(?x1:patty), RightAboveAndPrepped(?x0:cutting_board, ?x1:patty), On(?x1:patty, ?x0:cutting_board), [[0:robot].fingers<=[idx_0]0.5](?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:patty)
+Example learned predicates and learned action models in Combo Burger Learned Predicates: cooked0(?p:patty) chopped0(?l:lettuce) clear5(?r:robot) Learned operators: STRIPS-Op0: Parameters: [?x0:patty, ?x1:robot] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), OnGround(?x0:patty), clear5(?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:patty)] Delete Effects: [Clear(?x0:patty), OnGround(?x0:patty), clear5(?x1:robot)] Ignore Effects: [] Skill: Pick(?x1:robot, ?x0:patty) STRIPS-Op1: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:grill), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty)] Add Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:grill), clear5(?x2:robot)] Delete Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op2: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), On(?x1:patty, ?x0:grill), clear5(?x2:robot)] Add Effects: [cooked0(?x1:patty)] Delete Effects: [] Ignore Effects: [] Skill: Cook(?x2:robot, ?x1:patty, ?x0:grill) STRIPS-Op3: Parameters: [?x0:grill, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), On(?x1:patty, ?x0:grill), clear5(?x2:robot), cooked0(?x1:patty)] Add Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:patty)] Delete Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:grill), clear5(?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:patty) STRIPS-Op4: Parameters: [?x0:bottom_bun, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:bottom_bun), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty), OnGround(?x0:bottom_bun), cooked0(?x1:patty)] Add Effects: [Clear(?x1:patty), SomewhereAboveAndPrepped(?x0:bottom_bun, ?x1:patty), On(?x1:patty, ?x0:bottom_bun), clear5(?x2:robot)] Delete Effects: [Clear(?x0:bottom_bun), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:bottom_bun) STRIPS-Op5: Parameters: [?x0:robot, ?x1:top_bun] Preconditions: [Clear(?x0:robot), Clear(?x1:top_bun), OnGround(?x1:top_bun), clear5(?x0:robot)] Add Effects: [Holding(?x0:robot, ?x1:top_bun)] Delete Effects: [Clear(?x1:top_bun), OnGround(?x1:top_bun), clear5(?x0:robot)] Ignore Effects: [] Skill: Pick(?x0:robot, ?x1:top_bun) STRIPS-Op6: Parameters: [?x0:patty, ?x1:robot, ?x2:top_bun] Preconditions: [Clear(?x0:patty), Clear(?x1:robot), Holding(?x1:robot, ?x2:top_bun), cooked0(?x0:patty)] Add Effects: [Clear(?x2:top_bun), On(?x2:top_bun, ?x0:patty), clear5(?x1:robot)] Delete Effects: [Clear(?x0:patty), Holding(?x1:robot, ?x2:top_bun)] Ignore Effects: [] Skill: Place(?x1:robot, ?x2:top_bun, ?x0:patty) STRIPS-Op7: Parameters: [?x0:lettuce, ?x1:robot] Preconditions: [Clear(?x0:lettuce), Clear(?x1:robot), OnGround(?x0:lettuce), clear5(?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:lettuce)] Delete Effects: [Clear(?x0:lettuce), OnGround(?x0:lettuce), clear5(?x1:robot)] Ignore Effects: [] Skill: Pick(?x1:robot, ?x0:lettuce) STRIPS-Op8: Parameters: [?x0:cutting_board, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x0:cutting_board), Clear(?x2:robot), Holding(?x2:robot, ?x1:lettuce)] Add Effects: [Clear(?x1:lettuce), On(?x1:lettuce, ?x0:cutting_board), clear5(?x2:robot)] Delete Effects: [Clear(?x0:cutting_board), Holding(?x2:robot, ?x1:lettuce)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:lettuce, ?x0:cutting_board) STRIPS-Op9: Parameters: [?x0:cutting_board, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x1:lettuce), Clear(?x2:robot), On(?x1:lettuce, ?x0:cutting_board), clear5(?x2:robot)] Add Effects: [chopped0(?x1:lettuce)] Delete Effects: [] Ignore Effects: [] Skill: Chop(?x2:robot, ?x1:lettuce, ?x0:cutting_board) STRIPS-Op10: Parameters: [?x0:cutting_board, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x1:lettuce), Clear(?x2:robot), On(?x1:lettuce, ?x0:cutting_board), chopped0(?x1:lettuce), clear5(?x2:robot)] Add Effects: [Clear(?x0:cutting_board), Holding(?x2:robot, ?x1:lettuce)] Delete Effects: [Clear(?x1:lettuce), On(?x1:lettuce, ?x0:cutting_board), clear5(?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:lettuce) STRIPS-Op11: Parameters: [?x0:bottom_bun, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x0:bottom_bun), Clear(?x2:robot), Holding(?x2:robot, ?x1:lettuce), OnGround(?x0:bottom_bun), chopped0(?x1:lettuce)] Add Effects: [Clear(?x1:lettuce), SomewhereAboveAndPrepped(?x0:bottom_bun, ?x1:lettuce), On(?x1:lettuce, ?x0:bottom_bun), clear5(?x2:robot)] Delete Effects: [Clear(?x0:bottom_bun), Holding(?x2:robot, ?x1:lettuce)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:lettuce, ?x0:bottom_bun) STRIPS-Op12: Parameters: [?x0:lettuce, ?x1:robot, ?x2:top_bun] Preconditions: [Clear(?x0:lettuce), Clear(?x1:robot), Holding(?x1:robot, ?x2:top_bun), chopped0(?x0:lettuce)] Add Effects: [Clear(?x2:top_bun), On(?x2:top_bun, ?x0:lettuce), clear5(?x1:robot)] Delete Effects: [Clear(?x0:lettuce), Holding(?x1:robot, ?x2:top_bun)] Ignore Effects: [] Skill: Place(?x1:robot, ?x2:top_bun, ?x0:lettuce) STRIPS-Op13: Parameters: [?x0:grill, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x0:grill), Clear(?x2:robot), Holding(?x2:robot, ?x1:lettuce), chopped0(?x1:lettuce)] Add Effects: [Clear(?x1:lettuce), On(?x1:lettuce, ?x0:grill), clear5(?x2:robot)] Delete Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:lettuce)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:lettuce, ?x0:grill) STRIPS-Op14: Parameters: [?x0:cutting_board, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x0:cutting_board), Clear(?x2:robot), Holding(?x2:robot, ?x1:patty)] Add Effects: [Clear(?x1:patty), On(?x1:patty, ?x0:cutting_board), clear5(?x2:robot)] Delete Effects: [Clear(?x0:cutting_board), Holding(?x2:robot, ?x1:patty)] Ignore Effects: [] Skill: Place(?x2:robot, ?x1:patty, ?x0:cutting_board) STRIPS-Op15: Parameters: [?x0:grill, ?x1:lettuce, ?x2:robot] Preconditions: [Clear(?x1:lettuce), Clear(?x2:robot), On(?x1:lettuce, ?x0:grill), chopped0(?x1:lettuce), clear5(?x2:robot)] Add Effects: [Clear(?x0:grill), Holding(?x2:robot, ?x1:lettuce)] Delete Effects: [Clear(?x1:lettuce), On(?x1:lettuce, ?x0:grill), clear5(?x2:robot)] Ignore Effects: [] Skill: Pick(?x2:robot, ?x1:lettuce) STRIPS-Op16: Parameters: [?x0:lettuce, ?x1:patty, ?x2:robot] Preconditions: [Clear(?x1:patty), Clear(?x2:robot), Holding(?x2:robot, ?x0:lettuce), chopped0(?x0:lettuce)] Add Effects: [Clear(?x0:lettuce), SomewhereAboveAndPrepped(?x1:patty, ?x0:lettuce), On(?x0:lettuce, ?x1:patty), clear5(?x2:robot)] Delete Effects: [Clear(?x1:patty), Holding(?x2:robot, ?x0:lettuce)] Ignore Effects: [] Skill: Place(?x2:robot, ?x0:lettuce, ?x1:patty)
+Example learned predicates and learned action models in Coffee Learned predicates: CupReceivingLiquid0(?c:cup) JugFilled0(?j:jug) JugInMachine0(?j:jug, ?m:coffee_machine) Learned operators: STRIPS-Op0: Parameters: [?x0:cup, ?x1:cup, ?x2:jug, ?x3:robot] Preconditions: [HandEmpty(?x3:robot), JugPickable(?x2:jug), NotAboveCup(?x3:robot, ?x2:jug), NotSameCup(?x0:cup, ?x1:cup), NotSameCup(?x1:cup, ?x0:cup), OnTable(?x2:jug)] Add Effects: [Holding(?x3:robot, ?x2:jug), JugAboveCup(?x2:jug, ?x0:cup), JugAboveCup(?x2:jug, ?x1:cup), RobotAboveCup(?x3:robot, ?x0:cup), RobotAboveCup(?x3:robot, ?x1:cup)] Delete Effects: [HandEmpty(?x3:robot), NotAboveCup(?x3:robot, ?x2:jug), OnTable(?x2:jug)] Ignore Effects: [] Skill: PickJug(?x3:robot, ?x2:jug) STRIPS-Op1: Parameters: [?x0:coffee_machine, ?x1:cup, ?x2:cup, ?x3:jug, ?x4:robot] Preconditions: [Holding(?x4:robot, ?x3:jug), JugAboveCup(?x3:jug, ?x1:cup), JugAboveCup(?x3:jug, ?x2:cup), JugPickable(?x3:jug), NotSameCup(?x1:cup, ?x2:cup), NotSameCup(?x2:cup, ?x1:cup), RobotAboveCup(?x4:robot, ?x1:cup), RobotAboveCup(?x4:robot, ?x2:cup)] Add Effects: [HandEmpty(?x4:robot), JugInMachine0(?x3:jug, ?x0:coffee_machine), NotAboveCup(?x4:robot, ?x3:jug)] Delete Effects: [Holding(?x4:robot, ?x3:jug), JugAboveCup(?x3:jug, ?x1:cup), JugAboveCup(?x3:jug, ?x2:cup), RobotAboveCup(?x4:robot, ?x1:cup), RobotAboveCup(?x4:robot, ?x2:cup)] Ignore Effects: [] Skill: PlaceJugInMachine(?x4:robot, ?x3:jug, ?x0:coffee_machine) STRIPS-Op2: Parameters: [?x0:coffee_machine, ?x1:jug, ?x2:robot] Preconditions: [HandEmpty(?x2:robot), JugInMachine0(?x1:jug, ?x0:coffee_machine), JugPickable(?x1:jug), NotAboveCup(?x2:robot, ?x1:jug)] Add Effects: [JugFilled0(?x1:jug), PressingButton(?x2:robot, ?x0:coffee_machine)] Delete Effects: [] Ignore Effects: [] Skill: TurnMachineOnAndFill(?x2:robot, ?x0:coffee_machine) STRIPS-Op3: Parameters: [?x0:coffee_machine, ?x1:cup, ?x2:cup, ?x3:jug, ?x4:robot] Preconditions: [HandEmpty(?x4:robot), JugFilled0(?x3:jug), JugInMachine0(?x3:jug, ?x0:coffee_machine), JugPickable(?x3:jug), NotAboveCup(?x4:robot, ?x3:jug), NotSameCup(?x1:cup, ?x2:cup), NotSameCup(?x2:cup, ?x1:cup), PressingButton(?x4:robot, ?x0:coffee_machine)] Add Effects: [Holding(?x4:robot, ?x3:jug), JugAboveCup(?x3:jug, ?x1:cup), JugAboveCup(?x3:jug, ?x2:cup), RobotAboveCup(?x4:robot, ?x1:cup), RobotAboveCup(?x4:robot, ?x2:cup)] Delete Effects: [HandEmpty(?x4:robot), JugInMachine0(?x3:jug, ?x0:coffee_machine), NotAboveCup(?x4:robot, ?x3:jug), PressingButton(?x4:robot, ?x0:coffee_machine)] Ignore Effects: [] Skill: PickJug(?x4:robot, ?x3:jug) STRIPS-Op4: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [Holding(?x2:robot, ?x1:jug), JugAboveCup(?x1:jug, ?x0:cup), JugFilled0(?x1:jug), JugPickable(?x1:jug), RobotAboveCup(?x2:robot, ?x0:cup)] Add Effects: [CupFilled(?x0:cup), CupReceivingLiquid0(?x0:cup)] Delete Effects: [JugFilled0(?x1:jug), JugPickable(?x1:jug)] Ignore Effects: [] Skill: PourSomeLiquid(?x2:robot, ?x1:jug, ?x0:cup) STRIPS-Op5: Parameters: [?x0:cup, ?x1:cup, ?x2:jug, ?x3:robot] Preconditions: [CupFilled(?x1:cup), CupReceivingLiquid0(?x1:cup), Holding(?x3:robot, ?x2:jug), JugAboveCup(?x2:jug, ?x0:cup), JugAboveCup(?x2:jug, ?x1:cup), NotSameCup(?x0:cup, ?x1:cup), NotSameCup(?x1:cup, ?x0:cup), RobotAboveCup(?x3:robot, ?x0:cup), RobotAboveCup(?x3:robot, ?x1:cup)] Add Effects: [CupFilled(?x0:cup), CupReceivingLiquid0(?x0:cup)] Delete Effects: [CupReceivingLiquid0(?x1:cup)] Ignore Effects: [] Skill: PourSomeLiquid(?x3:robot, ?x2:jug, ?x0:cup) STRIPS-Op6: Parameters: [?x0:jug, ?x1:robot] Preconditions: [HandEmpty(?x1:robot), NotAboveCup(?x1:robot, ?x0:jug), OnTable(?x0:jug)] Add Effects: [JugPickable(?x0:jug)] Delete Effects: [] Ignore Effects: [] Skill: RotateItemUntilHandleAccessible(?x1:robot, ?x0:jug) STRIPS-Op7: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [HandEmpty(?x2:robot), JugPickable(?x1:jug), NotAboveCup(?x2:robot, ?x1:jug), OnTable(?x1:jug)] Add Effects: [Holding(?x2:robot, ?x1:jug), JugAboveCup(?x1:jug, ?x0:cup), RobotAboveCup(?x2:robot, ?x0:cup)] Delete Effects: [HandEmpty(?x2:robot), NotAboveCup(?x2:robot, ?x1:jug), OnTable(?x1:jug)] Ignore Effects: [] Skill: PickJug(?x2:robot, ?x1:jug) STRIPS-Op8: Parameters: [?x0:coffee_machine, ?x1:cup, ?x2:jug, ?x3:robot] Preconditions: [Holding(?x3:robot, ?x2:jug), JugAboveCup(?x2:jug, ?x1:cup), JugPickable(?x2:jug), RobotAboveCup(?x3:robot, ?x1:cup)] Add Effects: [HandEmpty(?x3:robot), JugInMachine0(?x2:jug, ?x0:coffee_machine), NotAboveCup(?x3:robot, ?x2:jug)] Delete Effects: [Holding(?x3:robot, ?x2:jug), JugAboveCup(?x2:jug, ?x1:cup), RobotAboveCup(?x3:robot, ?x1:cup)] Ignore Effects: [] Skill: PlaceJugInMachine(?x3:robot, ?x2:jug, ?x0:coffee_machine) STRIPS-Op9: Parameters: [?x0:coffee_machine, ?x1:cup, ?x2:jug, ?x3:robot] Preconditions: [HandEmpty(?x3:robot), JugFilled0(?x2:jug), JugInMachine0(?x2:jug, ?x0:coffee_machine), JugPickable(?x2:jug), NotAboveCup(?x3:robot, ?x2:jug), PressingButton(?x3:robot, ?x0:coffee_machine)] Add Effects: [Holding(?x3:robot, ?x2:jug), JugAboveCup(?x2:jug, ?x1:cup), RobotAboveCup(?x3:robot, ?x1:cup)] Delete Effects: [HandEmpty(?x3:robot), JugInMachine0(?x2:jug, ?x0:coffee_machine), NotAboveCup(?x3:robot, ?x2:jug), PressingButton(?x3:robot, ?x0:coffee_machine)] Ignore Effects: [] Skill: PickJug(?x3:robot, ?x2:jug) STRIPS-Op10: Parameters: [?x0:cup, ?x1:jug, ?x2:robot] Preconditions: [Holding(?x2:robot, ?x1:jug), JugAboveCup(?x1:jug, ?x0:cup), JugFilled0(?x1:jug), JugPickable(?x1:jug), RobotAboveCup(?x2:robot, ?x0:cup)] Add Effects: [CupFilled(?x0:cup)] Delete Effects: [JugFilled0(?x1:jug), JugPickable(?x1:jug)] Ignore Effects: [] Skill: PourSomeLiquid(?x2:robot, ?x1:jug, ?x0:cup)
+Example learned predicates and learned action models in Juice Learned predicates: IsCup0 NearJuiceValve0 InsideWasteValveRegion0 HoldingContainer0 OutsideJuiceMachine0 LidClosed0 HoldingObj0 Functional0 JuiceMachineOpen0 IsPlastic0 HandEmpty0 HasContents0 Learned operators: STRIPS-Op0: Parameters: [?x0:robot, ?x1:juicer] Preconditions: [Functional0(?x1:juicer), HandEmpty0(?x0:robot), JuiceMachineOpen0(?x1:juicer)] Add Effects: [LidClosed0(?x1:juicer)] Delete Effects: [JuiceMachineOpen0(?x1:juicer)] Ignore Effects: [] Skill: CloseLid(?x0:robot, ?x1:juicer) STRIPS-Op1: Parameters: [?x0:robot, ?x1:container] Preconditions: [HandEmpty0(?x0:robot), IsCup0(?x1:container)] Add Effects: [HoldingContainer0(?x0:robot, ?x1:container)] Delete Effects: [HandEmpty0(?x0:robot)] Ignore Effects: [] Skill: PickContainer(?x0:robot, ?x1:container) STRIPS-Op2: Parameters: [?x0:robot, ?x1:juicer, ?x2:container] Preconditions: [Empty(?x2:container), Functional0(?x1:juicer), HoldingContainer0(?x0:robot, ?x2:container), IsCup0(?x2:container), LidClosed0(?x1:juicer)] Add Effects: [HandEmpty0(?x0:robot), NearJuiceValve0(?x2:container, ?x1:juicer)] Delete Effects: [HoldingContainer0(?x0:robot, ?x2:container)] Ignore Effects: [] Skill: PlaceInsideJuiceValveRegion(?x0:robot, ?x2:container, ?x1:juicer) STRIPS-Op5: Parameters: [?x0:robot, ?x1:juicer, ?x2:container] Preconditions: [Empty(?x2:container), Functional0(?x1:juicer), HoldingContainer0(?x0:robot, ?x2:container), IsCup0(?x2:container), IsPlastic0(?x2:container)] Add Effects: [HandEmpty0(?x0:robot), InsideWasteValveRegion0(?x2:container, ?x1:juicer)] Delete Effects: [HoldingContainer0(?x0:robot, ?x2:container)] Ignore Effects: [] Skill: PlaceInsideWasteValveRegion(?x0:robot, ?x2:container, ?x1:juicer) STRIPS-Op7: Parameters: [?x0:container, ?x1:container, ?x2:robot] Preconditions: [Empty(?x0:container), HasContents0(?x1:container), HoldingContainer0(?x2:robot, ?x1:container), IsCup0(?x1:container), IsPlastic0(?x1:container)] Add Effects: [Empty(?x1:container), HasContents0(?x0:container)] Delete Effects: [Empty(?x0:container), HasContents0(?x1:container)] Ignore Effects: [] Skill: DumpFromOneIntoOther(?x2:robot, ?x1:container, ?x0:container) STRIPS-Op10: Parameters: [?x0:container, ?x1:container, ?x2:robot, ?x3:juicer, ?x4:movable_object] Preconditions: [Empty(?x0:container), Empty(?x1:container), Functional0(?x3:juicer), HandEmpty0(?x2:robot), Inside(?x4:movable_object, ?x3:juicer), InsideWasteValveRegion0(?x1:container, ?x3:juicer), IsCup0(?x0:container), IsCup0(?x1:container), IsPlastic0(?x1:container), LidClosed0(?x3:juicer), NearJuiceValve0(?x0:container, ?x3:juicer)] Add Effects: [HasContents0(?x0:container), HasContents0(?x1:container), JuiceInCup(?x4:movable_object, ?x0:container), OutsideJuiceMachine0(?x4:movable_object)] Delete Effects: [Empty(?x0:container), Empty(?x1:container), Inside(?x4:movable_object, ?x3:juicer)] Ignore Effects: [] Skill: TurnOnAndRunMachine(?x2:robot, ?x3:juicer, ?x0:container, ?x1:container) STRIPS-Op11: Parameters: [?x0:robot, ?x1:movable_object] Preconditions: [HandEmpty0(?x0:robot), OutsideJuiceMachine0(?x1:movable_object)] Add Effects: [HoldingObj0(?x0:robot, ?x1:movable_object)] Delete Effects: [HandEmpty0(?x0:robot)] Ignore Effects: [] Skill: PickMovable(?x0:robot, ?x1:movable_object) STRIPS-Op12: Parameters: [?x0:robot, ?x1:juicer, ?x2:movable_object] Preconditions: [HoldingObj0(?x0:robot, ?x2:movable_object), JuiceMachineOpen0(?x1:juicer), OutsideJuiceMachine0(?x2:movable_object)] Add Effects: [HandEmpty0(?x0:robot), Inside(?x2:movable_object, ?x1:juicer)] Delete Effects: [HoldingObj0(?x0:robot, ?x2:movable_object), OutsideJuiceMachine0(?x2:movable_object)] Ignore Effects: [] Skill: PlaceInside(?x0:robot, ?x2:movable_object, ?x1:juicer)
+Example learned predicates and learned action models in Cleanup Learned predicates: IsOnFloor0 IsEraser0 Holding0 NoObjectsOnTop1 IsOn0 HandEmpty0 Learned operators: STRIPS-Op0: Parameters: [?x0:movable_object, ?x1:table, ?x2:robot] Preconditions: [HandEmpty0(?x2:robot), IsOn0(?x0:movable_object, ?x1:table)] Add Effects: [Holding0(?x2:robot, ?x0:movable_object), NoObjectsOnTop1(?x1:table)] Delete Effects: [HandEmpty0(?x2:robot), IsOn0(?x0:movable_object, ?x1:table)] Ignore Effects: [] Skill: PickFromTop(?x2:robot, ?x0:movable_object, ?x1:table) STRIPS-Op1: Parameters: [?x0:movable_object, ?x1:robot, ?x2:trash_can] Preconditions: [Holding0(?x1:robot, ?x0:movable_object)] Add Effects: [HandEmpty0(?x1:robot), InsideContainer(?x0:movable_object, ?x2:trash_can)] Delete Effects: [Holding0(?x1:robot, ?x0:movable_object)] Ignore Effects: [] Skill: PlaceInside(?x1:robot, ?x0:movable_object, ?x2:trash_can) STRIPS-Op2: Parameters: [?x0:movable_object, ?x1:robot] Preconditions: [HandEmpty0(?x1:robot), IsOnFloor0(?x0:movable_object)] Add Effects: [Holding0(?x1:robot, ?x0:movable_object)] Delete Effects: [HandEmpty0(?x1:robot), IsOnFloor0(?x0:movable_object)] Ignore Effects: [] Skill: PickFromFloor(?x1:robot, ?x0:movable_object) STRIPS-Op3: Parameters: [?x0:table, ?x1:movable_object, ?x2:robot] Preconditions: [Holding0(?x2:robot, ?x1:movable_object), IsEraser0(?x1:movable_object), NoObjectsOnTop1(?x0:table)] Add Effects: [WipedOfMarkerScribbles(?x0:table)] Delete Effects: [] Ignore Effects: [] Skill: WipeAndContinueHoldingEraser(?x2:robot, ?x1:movable_object, ?x0:table) STRIPS-Op7: Parameters: [?x0:movable_object, ?x1:robot, ?x2:trash_can] Preconditions: [HandEmpty0(?x1:robot), InsideContainer(?x0:movable_object, ?x2:trash_can)] Add Effects: [IsOnFloor0(?x0:movable_object)] Delete Effects: [InsideContainer(?x0:movable_object, ?x2:trash_can)] Ignore Effects: [] Skill: DumpContentsOntoFloor(?x1:robot, ?x2:trash_can)